### v2

# Channel Dependence, Limited Lookback Windows, and the Simplicity of Datasets: How Biased is Time Series Forecasting?

Abstract.
In Time Series Forecasting (TSF), the lookback window (the length of historical data used for prediction) is a critical hyperparameter that is often set arbitrarily, undermining the validity of model evaluations. We argue that the lookback window must be tuned on a per-task basis to ensure fair comparisons. Our empirical results show that failing to do so can invert performance rankings, particularly when comparing univariate and multivariate methods. Experiments on standard benchmarks reposition Channel-Independent (CI) models, such as PatchTST, as state-of-the-art methods. However, we reveal this superior performance is largely an artifact of weak inter-channel correlations and simplicity of patterns within these specific datasets. Using Granger causality analysis and ODE datasets (with implicit channel correlations), we demonstrate that the true strength of multivariate Channel-Dependent (CD) models emerges on datasets with strong, inherent cross-channel dependencies, where they significantly outperform CI models. We conclude with three key recommendations for improving TSF research: (i) treat the lookback window as a critical hyperparameter to be tuned, (ii) use statistical analysis of datasets to inform the choice between CI and CD architectures, and (iii) favor CD models in applications with limited data.
conference: ; ; copyright: none
1. Introduction
Figure 1. Test-error (MSE) on y-axis for various lookback windows on the x-axis for the weather dataset and a forecast horizon of 96. CI denotes the channel-independent variant and CD denotes the channel-dependent variant of TSF architectures. The black dotted line shows lookback window of 96, where most methods compare performance.
Time Series Forecasting (TSF) is a fundamental challenge in machine learning that drives key innovations across domains such as finance, energy, healthcare, and climate science (Godahewa et al., 2021). The core task of TSF is to predict future observations by leveraging patterns within historical time-dependent data. Nevertheless, the literature suffers from two significant experimental shortcomings that compromise the fairness and reliability of model comparisons.
First, in the multivariate TSF learning setting, the historical behavior is modeled using a fixed-length sequence of past observations—known as the lookback window—to forecast a set of future values over a given prediction horizon. Despite its fundamental role, the choice of lookback window remains surprisingly underexplored in TSF literature. Most existing Long-term Time Series Forecasting (LTSF) studies adopt a fixed lookback window (mostly 96) for all models and datasets, often without justification or transparency (Liu et al., 2024; Zeng et al., 2023; Zhang and Yan, 2023; Yi et al., 2023; Chen et al., 2023, 2025; Lin et al., 2024, 2025; Huang et al., 2025).
While this practice simplifies experimental comparisons, it masks a critical reality: the optimal lookback window can vary significantly across models and datasets, and failing to tune it can lead to misleading conclusions. As demonstrated in Figure˜1, for the weather dataset, the optimal lookback window length differs across various state-of-the-art TSF models. Consequently, performance comparisons at a single, fixed lookback window may misrepresent which models are truly state-of-the-art under realistic, well-optimized conditions. Comparable studies have been conducted by Zeng et al. (2023) on the Electricity dataset and by Tong and Yuan (2025) on the Weather dataset. However, unlike previous studies, we explicitly tune all the hyperparameters for each lookback window to generate the figure ensuring reliable comparisons.
The optimal lookback window is highly dependent on both the intrinsic properties of the dataset and the specific architecture of the forecasting model. For example, when comparing within the same weather dataset (Figure˜1), Channel Independent (CI) or univariate models tend to favor longer lookback windows (e.g., 336 or 512), enabling deeper temporal modeling per variable (Nie et al., 2023; Zeng et al., 2023; Chen et al., 2023). Whereas the Channel Dependent (CD) or multivariate models jointly model all channels and often operate with shorter lookback windows (e.g., 96), leveraging channel dependence to compensate for the reduced temporal depth (Wang et al., 2024b; Liu et al., 2024). Even within CD class of methods (i.e, across different multivariate models) certain architectures benefit more from long histories than others. For example, Crossformer-CD significantly outperforms TimeMixer-CD at a lookback window of 96, yet its performance drops below TimeMixer-CD when the lookback window is extended to 336. These findings motivate our central claim:
Tuning the lookback window as a model-specific hyperparameter is essential for fair evaluation and reliable assessment in time series forecasting.
Beyond the issue of lookback tuning, a second critical limitation undermines the fairness and practical relevance of current TSF benchmarks. Most widely-used multivariate forecasting datasets such as ETTh, ETTm, Electricity, Weather, and Traffic exhibit minimal channel dependence among their variables. This lack of inter-variable relational complexity renders these benchmarks overly simplistic; even simple linear models (Zeng et al., 2023) can achieve performance comparable to sophisticated nonlinear architectures. This raises a key question:
Does channel dependence truly matter in multivariate LTSF?
To answer this, we conduct an in-depth empirical analysis, combining traditional performance metrics with statistical tools such as Granger causality (Granger, 1969). Our analysis reveals that the benefits of multivariate modeling depend strongly on two factors: (1) the amount of historical data available, and (2) the degree of causal influence among channels. When inter-channel interactions are weak or datasets are large enough for univariate models to exploit long-term patterns, the added complexity of modeling channel dependence may offer limited returns. Additionally, through experiments on complex channel-dependent chaotic Ordinary Differential Equation (ODE) datasets (Gilpin, 2021), we reveal that the standard benchmarks lack realistic cross-channel interactions essential for multivariate forecasting. This simplicity disproportionately favors CI models and reduce the ability to fairly assess models designed to exploit channel dependencies as also shown in (Nie et al., 2023; Chen et al., 2023).
Based on these experimental and statistical findings, we offer the following practical recommendations for evaluating and designing multivariate time series forecasting models:
  • Channel dependence is unnecessary for top performance on standard benchmarks; tuning the lookback window enables CI models like PatchTST to match or surpass CD ones (Section˜6.1).
  • On complex datasets with strong channel correlations (supported by granger causality Section˜6.2), CD models outperform CI models (Section˜6.3).
  • Longer lookback windows generally favor univariate models, highlighting the need for careful lookback tuning to ensure fair comparisons (Section˜6.4).
2. Related Works
LTSF models typically fall into two categories: Channel-Dependent (CD), which model cross-channel interactions, and Channel-Independent (CI), which treat each channel in isolation and focus purely on temporal dynamics. CI models such as DLinear (Zeng et al., 2023) and PatchTST (Nie et al., 2023) process each feature independently and consistently achieve strong performance on widely-used benchmarks, casting doubt on the necessity of cross-channel modeling. In contrast, MLP-based CD methods like TSMixer (Chen et al., 2023), TimeMixer (Wang et al., 2024a), and transformer-based architectures including iTransformer (Liu et al., 2024) and Crossformer (Zhang and Yan, 2023) aim to capture inter-channel dependencies explicitly. However, the predominantly univariate nature of standard benchmark datasets limits the ability of CD methods to demonstrate their advantages. To mitigate this issue, we incorporate evaluations on ODE-based datasets (Gilpin, 2021) designed to embed implicit channel correlations and enabling more representative comparisons on channel-dependent datasets.
The effectiveness of CI methods also appears to be amplified by longer lookback windows. For instance, DLinear adopts a lookback of 336 steps, whereas transformer-based baselines were defaulted to a lookback of 96. PatchTST, a successor to DLinear, employs a fixed lookback window of 512 or 336 and addresses the limitations of its predecessor by systematically tuning baseline methods over a broader range of lookback lengths. Nonetheless, the widespread adoption of a fixed 96-step lookback, popularized by iTransformer has constrained CI models and impeded equitable benchmarking. Subsequent studies (Tong and Yuan, 2025; Chen et al., 2025; Lin et al., 2024, 2025; Huang et al., 2025) frequently report improved performance of proposed methods by retaining short fixed windows for strong CI baselines like PatchTST, thereby hindering fair comparisons due to inconsistent tuning.
Several recent studies (Wang et al., 2024a; Nie et al., 2023; Zeng et al., 2023; Tong and Yuan, 2025) acknowledge the importance of lookback window selection in forecasting performance. However, lookback tuning is frequently confined to ablation studies, with many works failing to apply consistent tuning across both baselines and proposed methods in their main experiments. For instance, TimeMixer reports results using a unified set of hyperparameters (including a fixed lookback window) in the main results, while providing tuned results only in the appendix. Another work proposed IRPA method (Tong and Yuan, 2025) that explores various lookback lengths (480, 1920, 3600) for their method, yet retains a fixed lookback of 96 for all baselines, diminishing their comparative performance and inflating IRPA’s gains. In contrast, our evaluations perform systematic lookback window tuning for all models, uncovering genuinely optimal configurations.
3. Problem Statement
Multivariate time series is a sequence $({\bm{x}}^{l})_{l=1}^{T}\in(\mathbb{R}^{C})^{T}$, where each ${\bm{x}}^{l}\in\mathbb{R}^{C}$ represents observations from $C$ channels at time step $l$. This sequence can be represented as a matrix ${\bm{X}}^{\text{full}}\in\mathbb{R}^{C\times T}$. If $C=1$, this represents the univariate forecasting setting.
To train forecasting models effectively, especially when $T$ is large, we partition the long sequence into multiple shorter subsequences using a sliding window approach. Each subsequence is split into two parts: an input segment (lookback window) of length $L$ and a target segment (forecast horizon) of length $H$. Specifically, a time-series forecasting dataset is composed of input-output pairs $({\bm{X}},{\bm{Y}})\in\mathbb{R}^{C\times L}\times\mathbb{R}^{C\times H}$, where:
${\bm{X}}=({\bm{x}}^{t},{\bm{x}}^{t+1},\dots,{\bm{x}}^{t+L-1})$ is referred to as the forecasting query, and
${\bm{Y}}=({\bm{x}}^{t+L},{\bm{x}}^{t+L+1},\dots,{\bm{x}}^{t+L+H-1})$ is the corresponding forecasting answer.
The goal of the time-series forecasting problem is to train a model $\hat{{\bm{Y}}}:\mathbb{R}^{C\times L}\rightarrow\mathbb{R}^{C\times H}$ on a dataset $\mathcal{D}_{\operatorname{train}}\subseteq\mathbb{R}^{C\times L}\times\mathbb{R}^{C\times H}$, constructed from many such overlapping or non-overlapping windows, sampled from the full sequence. This dataset is assumed to be drawn i.i.d. from some underlying distribution $p$. The objective is to minimize the expected loss:
(1) $\mathbb{E}_{({\bm{X}},{\bm{Y}})\sim p}\,\ell({\bm{Y}},\hat{{\bm{Y}}}({\bm{X}}))$
where $\ell:\mathbb{R}^{C\times H}\times\mathbb{R}^{C\times H}\to\mathbb{R}$ is a suitable loss function, commonly the mean squared error (MSE).
4. Background
Table 1. Average MSE over the 4 forecasting horizons for each variant of the baseline evaluated for the standard datasets. The best variant for each model is in bold and the best overall model is in blue. Note here due to running out of memory (OOM) for some experiments related to electricity and traffic, we reported the results with "*" from TimeMixer paper from the experiment where they tune the lookback window as well as the hyperparameters (please refer to appendix E in (Wang et al., 2024a)).
Dataset PatchTST TSMixer Crossfor. DLin. iTrans. TimeMixer
CI CD CI CD CI CD CI CD CI CD
ETTh1 0.422 0.437 0.438 0.453 0.477 0.456 0.423 0.511 0.434 0.489
Weather 0.245 0.233 0.230 0.241 0.236 0.296 0.289 0.253 0.247 0.246
Electricity 0.159 0.170 0.162 0.170 0.166 0.188 0.162 0.220 0.173 0.260
ETTh2 0.365 0.382 0.378 0.390 0.741 0.691 0.507 0.363 0.371 0.378
ETTm1 0.356 0.371 0.356 0.370 0.393 0.426 0.359 0.374 0.355 0.408
ETTm2 0.258 0.259 0.257 0.273 0.433 0.485 0.289 0.257 0.275 0.342
Traffic 0.388 OOM 0.407 0.417 OOM 0.542 0.426 OOM OOM 0.388
Wins 3 0 2 0 0 0 0 2 1 0
Avg. Rank 2.14 5.43 2.71 5.00 7.29 8.14 5.43 5.86 5.29 6.57
Table 2. Average F-scores and percentage of rejected null hypotheses $H_{0}$ ($p=0.05$) across different lag values, comparing the standard datasets with the ODE benchmark. Bold values indicate the highest correlation at each lag.
Dataset Type (lag) Average F-score $H_{0}$ rejected (%, p=0.05)
ODEs (30) 335.21 73%
Standard (30) 2.46 61%
ODEs (96) 88.50 68%
Standard (96) 1.40 45%
ODEs (192) 35.29 63%
Standard (192) 1.23 27%
Table 3. Average MSE across four forecasting horizons for each model variant on the ODE Benchmark. Bold values highlight the best variant within each model; the best overall result per row is shown in blue.
Dataset PatchTST TSMixer CrossFor. DLin. iTrans. TimeMixer
CI CD CI CD CI CD CI CD CI CD
Lorenz 0.839 0.841 0.880 0.868 0.667 0.643 0.934 0.675 0.764 0.673
BlinkingRotlet 0.424 0.426 0.580 0.487 0.340 0.311 0.522 0.426 0.433 0.520
CellCycle 0.635 0.667 0.792 0.771 0.429 0.428 0.935 0.808 0.679 0.556
DoublePendulum 0.653 0.668 0.768 0.737 0.553 0.541 0.805 0.656 0.595 0.529
Hopfield 0.420 0.346 0.507 0.435 0.335 0.316 0.690 0.300 0.622 0.245
LorenzCoupled 0.881 0.866 0.950 0.900 0.700 0.666 0.963 0.857 0.788 0.832
Wins 0 0 0 0 0 4 0 0 0 2
Avg. Rank 5.17 5.67 8.83 7.50 2.50 1.50 9.83 5.17 5.50 3.30
Fixing Lookback Windows:
The problem formulation described earlier places no intrinsic requirement on the forecasting query $\bm{X}$} to have a fixed lookback length $L$. Nevertheless, most recent LTSF models fix the lookback window during training, typically to a value of $L=96$. This constraint is usually imposed for implementation convenience, computational efficiency, or alignment with benchmarking conventions, rather than for a principled reason. As we have argued, the lookback window should instead be treated as an integral property of the model, similar to any other model-specific hyperparameter.
Channel Independence:
One of the main findings of PatchTST (Nie et al., 2023) and later by TSMixer (Chen et al., 2023) is that, on widely used benchmark datasets, CI models are sufficient for accurate forecasting. This suggests that interactions between different channels are not essential, and that each channel can be modeled in isolation. Formally, a model $\hat{{\bm{Y}}}:\mathbb{R}^{C\times L}\to\mathbb{R}^{C\times H}$ is said to be channel-independent if for all ${\bm{X}}\in\mathbb{R}^{C\times L}$ and any channel input ${X}_{c}\in\mathbb{R}^{L}$,
(2) $\hat{{\bm{Y}}}({\bm{X}})_{c}=\hat{{\bm{Y}}}({{X}}_{c})$
i.e., the prediction for any given channel $c$ is generated using only its own historical data, independent of all other channels.
Granger Causality:
The Granger causality test (Granger, 1969) evaluates whether past values of $X_{c_{2}}$ improve forecasts of $Y_{c_{1}}$ beyond what is achievable using only past values of $X_{c_{1}}$. If including $X_{c_{2}}$ reduces prediction error in an autoregressive model, then $X_{c_{2}}$ provides a useful linear signal for forecasting channel $c_{1}$. To quantify this directional relationship, we compute the F-statistic (Equation˜5) for each channel pair, measuring the predictive gain from adding $X_{c_{2}}$.
The F-statistic compares the sum of squared residuals (SSR) of two models: a univariate model using only lagged values of $c_{1}$ ($SSR_{u}$, Equation˜3), and a multivariate model that additionally includes lagged values of $c_{2}$ ($SSR_{mv}$, Equation˜4). The univariate model uses coefficients $\hat{a}$ applied to $X_{c_{1}}^{t-i}$ over a lag window of length $L$, where $i=1,\dots,L$; the multivariate model extends this with coefficients $\hat{b}$ for $X_{c_{2}}^{t-i}$. Residuals are computed for $t=L+1,\dots,L+H$.
In Equation˜5, $k_{u}$ is the number of added parameters associated with channel $c_{2}$, and $k_{mv}$ is the total number of parameters in the multivariate model. The denominator degrees of freedom are $N-k_{mv}$, where $N$ is the number of observations. A large F-statistic indicates a significant improvement in explained variance when including $c_{2}$. We reject the null hypothesis $H_{0}\!:$$c_{2}$ does not Granger-cause $c_{1}$” when this improvement is statistically significant.
(3) $SSR_{u}=\sum_{t=L+1}^{L+H}\bigg(Y_{c_{1}}^{t}-\hat{a}_{0}-\sum_{i=1}^{L}\hat{a}_{i}X_{c_{1}}^{t-i}\bigg)^{2},$
(4) $SSR_{mv}=\sum_{t=L+1}^{L+H}\bigg(Y_{c_{1}}^{t}-\hat{a}_{0}-\sum_{i=1}^{L}\hat{a}_{i}X_{c_{1}}^{t-i}-\sum_{i=1}^{L}\hat{b}_{i}X_{c_{2}}^{t-i}\bigg)^{2},$
(5) $F=\frac{(SSR_{u}-SSR_{mv})/k_{u}}{SSR_{mv}/(N-k_{mv})},$
5. General Experimental Setup
5.1. Models
We evaluate a diverse set of strong baseline models, varying in architectural complexity and treatment of channel dependencies. These include MLP-based models (TSMixer (Chen et al., 2023), TimeMixer (Wang et al., 2024b)), attention-based models (PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2024), Crossformer (Zhang and Yan, 2023)), and the simple yet competitive DLinear (Zeng et al., 2023). We use only the CI version of DLinear, consistent with its design as a lightweight baseline. For iTransformer, we use only the CD version, as removing inter-channel attention would reduce it to a basic MLP, diverging from its core design. Rest of the methods are available in or allow derivation of both CD and CI forms. Additional details on the derivation of both variants are included in the appendix. For additional experimental details, please refer to our code repository [^1] https://github.com/anonsub936/How-Biased-is-Time-Series-Forecasting-.
5.2. Datasets
As noted in the introduction, we evaluate two main dataset categories to examine how increasing complexity particularly channel correlation impacts model performance.
The Standard Datasets:
The first category includes widely used benchmarks from recent state-of-the-art studies and surveys (Nie et al., 2023; Liu et al., 2024; Wang et al., 2024b; Chen et al., 2023; Zhang and Yan, 2023; Zeng et al., 2023): Weather, Electricity, Traffic, and the Electricity Transformer Temperature (ETT) datasets. These benchmarks gained popularity following their adoption in the Informer paper (Zhou et al., 2021). We refer to this group as the standard datasets. Details on data statistics are provided in the appendix.
The Chaotic ODE Benchmark:
As a second dataset collection, we consider the benchmark proposed by Gilpin (2021). This benchmark was introduced as a time-series collection that features complex time series dynamics with clear real world data properties, providing a challenging and realistic setting for channel-dependent time-series forecasting. A demonstration for one of these datasets and how it simulates a recurrent patterns across different channels can be seen in Figure˜2. This figure particularly emphasizes periods of synchronization in the angular velocities of the two pendulums during the final 200 timesteps, indicative of inter-channel correlation. Additional plots of the selected ODE system are provided in the appendix.
Figure 2. Time series visualization of the Double Pendulum ODE system. Notably, the bottom two channels (angular velocities of the two pendulums) exhibit synchronized behavior in the last 200 time steps, demonstrating time-varying inter-channel correlation.
We select a subset of datasets from this benchmark based on two distinctive criteria: (i) The datasets follow complex trajectories that challenge most time series models and (ii) they are derived directly from the underlying ODEs, ensuring a clear mathematical correlation between variables. For criterion (i), we use the Largest Lyapunov exponent (Gilpin, 2021) to quantify dataset complexity, selecting datasets with high values to ensure forecasting remains non-trivial task that cannot be solved easily based on one channel information. To illustrate criterion (ii), we show below a simplified form of the Double Pendulum system (Elbori and Abdalsmd, 2017), assuming equal masses and rod lengths. The coupling between angular velocities $\frac{d\theta_{1}(t)}{dt},\frac{d\theta_{2}(t)}{dt}$ and angles $\theta_{1}(t),\theta_{2}(t)$ at time $t$ from Equations˜6 and 7 is evident, with parameters $g$ (gravity) and $l$ (rod length) incorporated. Such systems are essential in modeling real-world physics applications, including robotics (Jadlovská and Sarnovskỳ, 2012).
Due to space constraints, we defer detailed discussion of another example, Cell Cycle, to the appendix. This dataset reflects strong variable coupling and captures real-world measurements of Cyclin-Dependent Kinase (CDK) activity, crucial to cell cycle progression (Romond et al., 1999). Our selection yields six datasets with rich temporal dynamics and strong inter-channel dependencies, referred to as the ODE datasets. Further reasoning for particularly choosing this benchmark follows through the statistical analysis in Section˜6.2.
(6) $\frac{d^{2}\theta_{1}(t)}{dt^{2}}$ $=-\frac{2g}{l}\theta_{1}(t)+\frac{g}{l}\theta_{2}(t)$
(7) $\frac{d^{2}\theta_{2}(t)}{dt^{2}}$ $=\frac{2g}{l}\theta_{1}(t)-\frac{2g}{l}\theta_{2}(t)$
5.3. Hyperparameter Tuning
To ensure fair evaluation, we apply a consistent tuning procedure across all models. Each hyperparameter is defined either by a discrete set (categorical) or a continuous range (real-valued). Performance for each hyperparameter configuration is validated on the designated validation set through the Mean Squared Error (MSE). The best model is chosen based on the lowest validation error, then it is tested on a the separate test split to get the final evaluation. We use Bayesian optimization via Optuna (Akiba et al., 2019) to efficiently explore promising configurations, running 20 trials per setup (i.e., a specific model, dataset, and forecasting horizon).
Crucially, the lookback window is treated as a tunable hyperparameter, selected from $\{96,192,336,512,720\}$. Also more hyperparameters that are specific for each model are tuned thoroughly, allowing each model to operate under optimal conditions for fair comparison. Detailed search spaces and best hyperparameters chosen for each model are provided in the appendix.
5.4. Granger Causality
We conduct Granger causality analysis following the standard procedure outlined by Granger (1969), with several preprocessing steps to ensure robust and meaningful results:
  1. (i)
    We first assess the stationarity of each dataset. If non-stationarity is detected, differencing is applied iteratively until stationarity is achieved.
  2. (ii)
    To remove redundancy, we compute the Pearson correlation coefficient for all channel pairs. Channels with correlation exceeding 0.95 are considered highly similar, and one from each such pair is excluded from further analysis.
After the preprocessing steps, we conduct the F-test described in Section˜4. Consequently, we report two important metrics: the average F-score over all valid channel pairs, and the percentage of pairs for which the null hypothesis is rejected (indicating significant Granger causality). These results are summarized in Section˜6.2. Additional implementation details, including parameter settings and the formal hypothesis definition, are provided in the appendix.
6. Results
6.1. Evaluation of Standard Datasets
We reproduce the current time-series forecasting models in our above-mentioned evaluation settings, where we conduct fair and extensive hyperparameter tuning, including tuning the lookback window. Unless stated otherwise, all results are averaged over 5 different random seeds. Table˜1 presents the results averaged across forecasting horizons 96, 192, 336, and 720 for the standard datasets. Detailed values for each individual horizon and standard-deviations are provided in the appendix.
PatchTST remains the state-of-the-art for time-series forecasting. In 3 out of 7 datasets, the standard channel-independent version of PatchTST achieves the best performance. For three other datasets, this version ranks second or third. Its performance is not among the top three only on the Weather dataset, which may be attributed to the relatively stronger inter-channel dependencies in this dataset, as indicated by Granger causality. The average rank of PatchTST across all datasets is 2.14 showing a significant gap of 0.57 compared to the second-best model, the channel-independent version of TSMixer.
Linear models, meanwhile, remain competitive with recent forecasting approaches. The performance of DLinear on the standard datasets is comparable to more recent models. Across the 6 datasets, it achieves an average rank of 5.43 out of 10. Notably, on ETTh1, it ranks second, trailing PatchTST by only a small margin.
6.2. Granger Causality for Evaluation of Channel Correlations
Based on the findings from the previous section, which showed that multivariate models do not consistently outperform univariate ones on the standard datasets, we turn to a statistical evaluation of channel correlations across a broader set of datasets. Our goal is to identify benchmarks where inter-channel dependencies exist, potentially finding datasets and evaluation setup where multivariate models are needed.
Although the full test results are presented in the appendix, we focus here on a representative comparison between the ODE datasets and the standard datasets, to highlight key patterns. We conduct Granger causality tests using three different lag values: 30, 96, and 192. These lags correspond to different lookback window sizes used in the autoregressive models underlying the tests. To further analyze channel dependence across a broad range of time series domains, we computed Granger causality values for 41 real-world datasets from the extensive Monash Time Series Forecasting Archive (Godahewa et al., 2021). These results, calculated for lag values of 7, 30, 96, and 192, are provided in the appendix. Throughout this section, we use the terms lag and lookback window interchangeably.
Varying the lag allows us to analyze how inter-channel correlations evolve with different temporal resolutions. We can see the F-scores and the percentage of pairs of channels that exhibits granger causality in Table˜2. For both metrics, the higher values indicate higher correlation in the datasets. As seen in the table, the ODE datasets exhibit significantly higher F-scores and pass rates across all lag values, strongly indicating the presence of meaningful inter-channel dependencies. On contrary, the standard datasets show considerably weaker correlations, with average F-scores dropping below 1.0 for ETTm2 dataset at the highest lag.
These findings align with our earlier empirical observations: univariate models often match or outperform multivariate models on the standard datasets due to weak inter-channel coupling. Moreover, we observe that increasing the lag generally reduces the measured correlation for both dataset types. This effect is particularly pronounced for the standard datasets, suggesting that longer lookback windows benefit more univariate models than multivariate ones, since the former cannot leverage strong inter-channel signals.
In contrast, the ODE datasets show much stronger correlations at longer lags, e.g., average F-scores of 88.5 (lag 96) and 35.29 (lag 192), compared to 1.40 and 1.23 for the standard datasets. These results support using the ODE datasets as a valuable testbed for evaluating models under strong cross-channel dynamics.
6.3. Evaluation on the ODE Benchmark
As discussed in Section˜6.1, our results reaffirm DLinear as a competitive baseline and also suggest the superiority of CI models over their CD counterparts. In this section, we evaluate how the strong correlations between channels in the ODE datasets affect model performance. Table˜3 presents the results obtained using the same evaluation protocol described in Section˜6.1.
The evaluation on the ODE datasets reveals a markedly different performance hierarchy among models. First of all, CD variants clearly outperform CI ones. The CD variants of the evaluated models consistently outperform their CI counterparts, highlighting the importance of modeling inter-channel dependencies, especially in datasets governed by coupled differential equations. As shown in Table˜4, CD models outperform CI versions in 18 out of 24 total experiments. This strongly supports the claim that model architecture should be tailored to the underlying data structure. Secondly, Linear models perform poorly on the ODE benchmark. Contrary to their performance on standard datasets, linear models such as DLinear are no longer competitive here. Averaged over all forecasting horizons, DLinear ranks worst in 5 out of the 6 ODE datasets, with an average rank of 9.83. Compared to Crossformer, DLinear suffers from performance drops ranging from 45% to 118%, with a median degradation of 58.3% and a mean of 74%.
In summary, the commonly held view that linear models are competitive with deep architectures does not generalize to datasets with strong inter-channel dynamics. Our findings underline the importance of evaluating forecasting models on datasets with diverse structural properties. In cases where inter-channel correlation is intrinsic to the data—as with ODE systems—model architectures must be explicitly designed to capture these dependencies in order to perform well.
6.4. Lookback Window Analysis
In this section, we extend our evaluation by analyzing the distribution of selected lookback window values, grouped by both datasets and models. This analysis highlights how an optimal lookback window varies significantly depending on the model type and dataset characteristics.
As shown in Figure˜3, we report the frequency of best-performing lookback window across CI models, CD models, and datasets. In the Figure 3(a), CI models exhibit substantial variation in their preferred lookback windows. For instance, DLinear consistently favors longer windows (336 or 720), while shorter windows are rarely optimal. This finding is soldified by the fact that overall, CI models choose 96 lookback window only 11% times. In contrast, Figure 3(b) shows that CD Models more often benefit from shorter windows, with 96 or 192 selected in 55% of the cases. These counts are aggregated across the four forecasting horizons (96, 192, 336, and 720).
Figure 3(c) shows that optimal lookback windows also vary widely by dataset. This aligns with expectations: for hourly datasets like ETTh1 and ETTh2, a window of 96 covers four days, often sufficient for effective forecasting. In contrast, for 15-minute interval datasets like ETTm1 and ETTm2, more timesteps are required to capture similar temporal spans. Accordingly, ETTm2 frequently favors longer windows larger than or equal to 336, while ETTh2 consistently prefers windows shorter than 336.
Overall, this analysis reinforces the importance of tuning the lookback window, not just across model types, but also in response to dataset-specific temporal characteristics. For per-horizon results, refer to the bar charts in the appendix.
(a) CI models
(b) CD models
(c) Datasets
Figure 3. Frequencies of best-performing lookback windows grouped by: (a) CI models, (b) CD models, and (c) datasets.
7. Conclusion and Recommendations
7.1. Conclusion
Model ODE-6 Standard-7
CI CD CI CD
PatchTST 4 2 6 1
TSMixer 0 6 7 0
Crossformer 0 6 4 3
TimeMixer 2 4 4 3
Totals 6 18 21 7
Table 4. Frequency of CI and CD variants outperforming each other across models and datasets.
In this paper, as outlined in the introduction, our primary goal is to establish a fair and unbiased evaluation framework for multivariate TSF models. To achieve this, we make several critical adjustments to the experimental setup including: tuning the lookback window, analyzing statistical inter-channel correlations, and conducting empirical evaluations on datasets with varying characteristics. From this comprehensive setup, we draw the following key conclusions:
  1. (1)
    Tuning the lookback window is essential for fairness in multivariate TSF. Fixing the lookback window (commonly at 96) can unfairly disadvantage univariate models. Allowing models to tune this parameter ensures a more balanced comparison and reveals meaningful performance differences. This insight directly enables the second key finding below.
  2. (2)
    Modeling channel dependencies is unnecessary for standard datasets. For all models where a CI variant is available, it consistently outperforms the CD counterpart, by a margin of 21 to 7 on standard benchmarks. This is a significant result, especially considering that models like TSMixer, Crossformer, and TimeMixer are inherently designed to mix information across channels. Our findings indicate that inter-channel mixing is not a major contributor to performance on these datasets and can, in fact, degrade it (Table˜4).
  3. (3)
    Dataset nature plays a critical role in multivariate TSF. This aspect has been largely overlooked in prior work. Our experiments (Table˜4) reveal that datasets with different underlying dynamics require different modeling strategies. For instance, ODE Datasets exhibit significantly higher inter-channel correlation, as confirmed by statistical tests. This justifies the use of multivariate models in such cases and is further supported by the superior performance of channel-dependent models on this dataset family.
7.2. Recommendations for Multivariate Time-Series Forecasting Evaluation
In this paper, we strive to come with a useful insights that can advance the field of multivariate time series forecasting and this was done through a detailed experiments and statistical analysis from which we draw the following recommendations:
  • Do not use a fixed lookback window when a vast amount of historical data is available to ensure full potential of all models considered as seen in Section˜6.1 and Section˜6.3.
  • For proper evaluation and before choosing model architecture, consider statistically testing the channel correlation of the datasets as shown in Section˜6.2, it can help a lot in making an informed decision between deploying a CD or a CI model.
  • As noted in Section˜6.4, shorter lookback window favors in general CD models, which means for applications with limited amount of historical data, an architecture capturing channel dependency is generally recommended.
Appendix A Description of All Datasets
In this section, we would like to show in detail the different datasets used in our experiments. Starting with the standard datasets which were explained a lot in literature, so we will just give a brief description of them. First ETT- datasets which representing indication of long term power deployment which are recorded for 2 different granalurities (15 mins for ’m’ and 1 hour for ’h’) for 2 different counties (hence the post-fix ’1’ and ’2’). Electricity dataset follows the consumption of electricity over an extended amount of time (3 years) recorded hourly for 321 different households. Traffic dataset is representing the flow of traffic recorded through multiple sensors in an hourly frequency spanning over a period of approximately 2 years. All of this information as well as the train/validation/test split is shown in Table˜5
Dataset/Feature Channels Granularity Train/val/test Total Timesteps
ETTh (1/2) 7 1 hour 12/4/4 (months) 17420
ETTm (1/2) 7 15 mins 12/4/4 (months) 69680
electricity 321 1 hour 7:1:2 26304
traffic 862 1 hour 7:1:2 17544
weather 21 10 mins 7:1:2 52696
Table 5. Basic statistics about the standard datasets widely used in the time series forecasting literature. This table shows the number of features, frequency of recording the data, train/validation/test split as well as the total timesteps recorded in a given dataset.
Second type of datasets used in our experiments is based on the chaotic ODE Benchmark. Here, we elaborate more on the general common features of the used ODE datasets and we give further description of the underlying processes for generating them. For the generation process, we use a common protocol for the chosen 6 attractors which is generating 20 points per time unit, splitting the data into 7:1:2 train:validation:test split which is the common practice for the standard datasets as well. Finally, we generate a very long time series 60000 timesteps which is divided into time series forecasting samples though a rolling window (same approach that is adopted on TSF papers for the standard datasets). For reproducibility, we provide the code stating the random seed that should be used to produce the same datasets and subsequently the same results we got. For understanding more the chosen processes as well as the dimensionality of each dataset, please refer to the chaotic ODE Benchmark paper (Gilpin, 2021).
Appendix B Hyperparameter Tuning
We evaluate various Transformer-based and MLP-based time series forecasting models (e.g., DLinear, TSMixer, PatchTST, Crossformer, and our proposed FaCT) under a unified set of hyperparameter search ranges as shown in Table Table˜6.
Hyperparameter Search Range
Learning Rate $10^{-7},\,10^{-2}]$
Hidden Dimension (d_modeld\_model) $\{128,256,512,1024\}$
Feedforward Dimension (d_ffd\_ff) $\{128,256,512,1024\}$
# Encoder/Mixer Layers [1, 10][1,\,10]
Dropout [0, 0.9][0,\,0.9]
Sequence Length (seq_lenseq\_len) $\{96,192,336,512,720\}$
Table 6. Common Hyperparameter Search Ranges with all parameters being integer datatypes except for learning rate and dropout which span the whole range of floats on the respective range of values
For specific architectures, additional parameters that are model specific were tuned on the same range of parameters specified on the respective papers. For all the range $[]$ or set of values $\{\}$, refer to Table˜7.
Model Hyperparameter Range
PatchTST Patch Size $\{8,16\}$
Stride $\{4,8\}$
TSMixer Hidden Size $\{32,64,256,1024\}$
Crossformer/FaCT Segement Length $[3,12]$
Baseline $\{0,1\}$
Cross Factor $[3,20]$
Table 7. Model Specific Hyperparameter Search Ranges/set of possible values with all of them having a datatype of integer in this case.
To ensure easy and full reproducibility of our experiments, we include all the best performing hyperparameter sets corresponding to all models and datasets. Please see tables from Table˜16 to Table˜28 for the full list. We hope inclusion of these lists, make reproducibility of this paper’s results straight forward.
Appendix C Implementation Details
C.1. Libraries and Hardware
All models were implemented using the PyTorch library (version 2.4) in Python (version 3.12.1). For hyperparameter optimization, we employed Optuna (version 3.6.1). Experiments were conducted on machines equipped with NVIDIA RTX 4090 and RTX 3090 GPUs.
C.2. Training Procedure
All models are trained using the Adam optimizer. We use mean squared error (MSE) as the default loss function, and each epoch iterates over mini-batches of size 32 by default. Training proceeds for a maximum of 100 epochs, or until early stopping is triggered if the validation loss fails to improve for 10 consecutive epochs. We use Optuna to conduct 20 trials per model, selecting configurations from the defined set of hyperparameters B that yield the best validation performance.
C.3. Granger Causality
Here, we list in more details the procedure to do the granger causality test which, alongside our code, should guide an easy usage of the framework for other datasets as well as easy reproduction of current results. We pick directly the first 1000 timesteps from the loaded dataset for this statistical test purposes, which we think is reasonable sample size with regard to the lags considered here. First, we check the pearson correlation between each pair of channels, and if it is too high (greater than 0,95 in our case), then one of the two channels is dropped. This is done, because channels which are very similar can cause results that are misleading when testing pair of channels that are identical (that can easily yield a very high score that will push the average scores way higher than expected). After, the initial filtering, we get dataset with diverse channels that can still have strong correlation, but one more important aspect we address next is stationarity as grangar causality works best on stationary time series data (i.e. mean and variance should not be shifting over time). Here, we apply the augmented Dickey-Fuller test on each channel of the time series to check if it is stationary. If the channel is not stationary, we apply differencing which should help bring stationarity into the time series channel. Now we have our preprocessed data which is ready for the grangar causality to be applied on. We use the packages from statsmodels package (Seabold and Perktold, 2010).
Appendix D Derivation of CI/CD version of Models
In our paper, 6 strong recent baselines are used throughout evaluation, they vary in complexity and structure. The aim of this appendix section is to ease the understanding of making CI and CD versions of the different baselines.
Simpler Linear/MLP-based Models
We start with the simpler models used in our experiments which are DLinear and TSMixer. DLinear as mentioned earlier is based on being the simplest model that applies a trend seasonality deceomposition followed by simpler linear layers applied on both the trend and the seasonal components (Zeng et al., 2023). For TSMixer, the authors had implemented a CD version which is the original version of the model and removed the channel mixing component which is called in their paper TMix-Only. We reimplemnted the model using pytorch and made sure of reproducing the paper results before using both versions in our experiments. Throughout our paper, we call both versions as TSMixer (CD) and TSMixer (CI) for the CD and CI versions respectively.
PatchTST
Now we move into transformer-based models, starting with PatchTST as one of the first transformer models to re-establish a boost of performance over the DLinear model. In their paper, the original version of the model is CI where the transformer encoder backbone is applied independently on different channels before the application of a linear projection layer. For the CD version, they briefly discuss it in the appendix of applying the same backbone but on a flattened dimension of both number of patches and channels which means the attention is now applied jointly over patches and channels. We use the original implementation of the paper but we had to add the part related to CD implementation ourselves as it was not readily available on the paper at the time of writing our paper.
iTransformer/Crossformer
Next, transformer model that proved to be a strong baseline is Crossformer which applies a Depth-Segment-Wise embedding before application of a Two Stage Attention (TSA). Finally a decoder layer is applied through cross attention between the positional embedding of output time info and encoder output. The TSA is applied on the segments dimension (temporal) as well as followed by the channel dimension which constitutes the original CD version of the model (Zhang and Yan, 2023). For a CI version, we just removed the second stage of the TSA to make the attention only applied on the segments dimension. We used authors implementation of the CD version and made the mentioned edits on the implementation to produce the CI version used in our experiments. For iTransformer (Liu et al., 2024), we don’t introduce additional CI version to the existing CD version as otherwise the model will be simplified to a linear model which is already covered in this paper.
TimeMixer
One more prominent model proposed recently which proves be competitive on the standard datasets is TimeMixer (Wang et al., 2024a). In this model, the authors provide an MLP-based approach which is applied over multiple resolutions of the time series. This model original setting is CD where the MLP-mixing is done over both temporal and channel dimensions. For a CI version, the channel mixing component can be removed yielding a multiresolution temporal mixing based model. We use the authors implementation for both the CI and CD versions of the model.
Appendix E Detailed Results
In this appendix, we add the full results over all forecasting horizons for both types of introduced datasets. In Table˜9, we can see in detail the results for the chaotic ODE datasets, confirming the results from the main paper that Crossformer is clearly the SoTA on this benchmark with it being the best model in 15 out of possible 24 settings(for the CD version). For the standard datasets, the results are detailed in Table˜8 which also confirms the state of the current datasets where no model is clearly the best even emphasized with DLinear being the best model on 5 occurrences. This also fits the picture that was formulated on the table from the main paper text that when proper tuning is carried out for lookback window as well as all related hyperparameters, PatchTST remains competitive being the best overall 6 times. To conclude how comprehensive and reproducible our experiments are, these results represents the mean of 5 random seeds {3001, 3002, 3003, 3004, 3005} MSE error on the test split of each dataset.
Dataset / Model Horizon PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
CI CD CI CD CI CD CI CD CI CD
ETTh1 96 0.375 0.381 0.374 0.404 0.396 0.400 0.372 0.413 0.391 0.426
192 0.419 0.407 0.432 0.438 0.452 0.446 0.406 0.457 0.434 0.580
336 0.459 0.460 0.447 0.471 0.459 0.480 0.435 0.500 0.444 0.500
720 0.437 0.499 0.500 0.499 0.601 0.497 0.481 0.541 0.498 0.473
Weather 96 0.151 0.150 0.152 0.155 0.150 0.153 0.374 0.156 0.158 0.199
192 0.193 0.197 0.196 0.202 0.199 0.193 0.210 0.201 0.198 0.289
336 0.278 0.247 0.247 0.255 0.261 0.495* 0.255 0.263 0.266 0.343
720 0.359 0.336 0.330 0.354 0.335 0.342 0.316 0.319 0.342 0.442
Electricity 96 0.129* 0.141 0.131 0.137 0.139 0.150 0.135 0.135 0.130 0.143
192 0.147* 0.156 0.149 0.156 0.177 0.168 0.149 0.151 0.149 0.191
336 0.163* 0.173 0.165 0.176 0.194 0.182* 0.164 0.175 0.171 0.174
720 0.197* 0.210 0.203 0.210 0.261 0.251* 0.199 0.196 0.212 0.197
ETTh2 96 0.286 0.285 0.291 0.304 0.397 0.537 0.303 0.324 0.285 0.377
192 0.352 0.381 0.376 0.392 0.713 0.794 0.397 0.396 0.360 0.459
336 0.392 0.422 0.402 0.421 0.727 0.553 0.518 0.447 0.410 0.476
720 0.431 0.438 0.444 0.441 1.126 0.880 0.811 0.441 0.431 0.603
ETTm1 96 0.293 0.306 0.292 0.294 0.309 0.364 0.299 0.314 0.299 0.366
192 0.334 0.339 0.333 0.344 0.398 0.411 0.334 0.359 0.330 0.340
336 0.373 0.400 0.374 0.389 0.399 0.463 0.368 0.405 0.363 0.468
720 0.425 0.441 0.426 0.452 0.466 0.465 0.434 0.437 0.418 0.607
ETTm2 96 0.161 0.166 0.163 0.176 0.199 0.188 0.166 0.180 0.166 0.167
192 0.221 0.220 0.234 0.235 0.270 0.305 0.236 0.230 0.220 0.242
336 0.287 0.273 0.271 0.275 0.820 0.660 0.296 0.287 0.272 0.292
720 0.363 0.375 0.360 0.414 0.442 0.787 0.458 0.369 0.361 0.384
Traffic 96 0.360* OOM 0.386 0.384 OOM 0.514* 0.395 OOM OOM 0.360*
192 0.375* OOM 0.392 0.405 OOM 0.549* 0.406 OOM OOM 0.375*
336 0.385* OOM 0.407 0.423 OOM 0.530* 0.436 OOM OOM 0.385*
720 0.43* OOM 0.443 0.457 OOM 0.573* 0.466 OOM OOM 0.43*
Table 8. In this table, the full results over different horizons are shown over all the 7 standard datasets. These results were average of 5 different random seeds {3001,3002,3003,3004,3005}. The best result for each variant per model is in bold, while the best results overall for each horizon per dataset is highlighted in blue. we reported the results with "*" from TimeMixer paper from the experiment where they tune the lookback window as well as the hyperparameters (please refer to appendix E in (Wang et al., 2024a)).
Dataset / Model Horizon PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
CI CD CI CD CI CD CI CD CI CD
Lorenz 96 0.658 0.658 0.747 0.729 0.331 0.266 0.884 0.433 0.398 0.338
192 0.825 0.832 0.862 0.847 0.623 0.631 0.922 0.712 0.662 0.884
336 0.906 0.906 0.927 0.926 0.792 0.762 0.950 0.852 0.878 0.895
720 0.965 0.970 0.983 0.971 0.920 0.914 0.978 0.954 0.962 0.929
BlinkingRotlet 96 0.162 0.156 0.322 0.228 0.064 0.045 0.374 0.112 0.116 0.076
192 0.322 0.321 0.510 0.429 0.230 0.210 0.504 0.299 0.272 0.315
336 0.511 0.540 0.727 0.587 0.491 0.466 0.576 0.481 0.599 0.500
720 0.703 0.688 0.759 0.703 0.576 0.524 0.635 0.623 0.710 0.622
CellCycle 96 0.263 0.311 0.513 0.505 0.036 0.035 0.866 0.227 0.110 0.071
192 0.580 0.624 0.791 0.740 0.244 0.275 0.932 0.515 0.475 0.370
336 0.768 0.795 0.894 0.876 0.612 0.556 0.957 0.688 0.742 0.719
720 0.931 0.939 0.970 0.963 0.825 0.844 0.984 0.894 0.867 0.892
DoublePendulum 96 0.322 0.278 0.541 0.461 0.090 0.083 0.667 0.314 0.250 0.110
192 0.551 0.594 0.713 0.753 0.418 0.466 0.762 0.534 0.440 0.380
336 0.806 0.825 0.847 0.808 0.802 0.712 0.856 0.807 0.702 0.650
720 0.933 0.975 0.973 0.926 0.902 0.904 0.933 0.952 0.901 1.018
Hopfield 96 0.156 0.073 0.268 0.185 0.059 0.049 0.472 0.046 0.054 0.036
192 0.311 0.216 0.410 0.321 0.162 0.155 0.641 0.128 0.118 0.102
336 0.486 0.439 0.571 0.494 0.399 0.404 0.759 0.280 0.567 0.199
720 0.727 0.654 0.778 0.742 0.721 0.655 0.886 0.663 0.699 0.717
LorenzCoupled 96 0.696 0.587 0.840 0.745 0.372 0.276 0.919 0.610 0.458 0.818
192 0.883 0.901 0.949 0.895 0.687 0.635 0.959 0.840 0.795 0.947
336 0.957 0.971 0.997 0.966 0.813 0.810 0.978 0.924 0.893 0.984
720 0.990 1.007 1.013 0.994 0.926 0.942 0.996 0.984 0.969 1.003
Table 9. In this table, the full results over different horizons are shown over all the 6 chaotic ODE datasets. These results were average of 5 different random seeds {3001,3002,3003,3004,3005}. The best result for each variant per model is in bold, while the best results overall for each horizon per dataset is highlighted in blue. we reported the results with "*" from TimeMixer paper from the experiment where they tune the lookback window as well as the hyperparameters (please refer to appendix E in (Wang et al., 2024a)).
Dataset / Model PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
CI CD CI CD CI CD CI CD CI CD
Lorenz 0.0033 0.0025 0.0020 0.0115 0.0002 0.0095 0.0065 0.0303 0.0267 0.0066
BlinkingRotlet 0.0063 0.0072 0.0063 0.0083 0.0023 0.0171 0.0085 0.0580 0.0146 0.0294
CellCycle 0.0030 0.0040 0.0030 0.0047 0.0000 0.0093 0.0163 0.1127 0.0126 0.0818
DoublePendulum 0.0070 0.0067 0.0070 0.0305 0.0005 0.0110 0.0088 0.0136 0.0026 0.0052
Hopfield 0.0057 0.0070 0.0057 0.0075 0.0007 0.0049 0.0129 0.0262 0.3732 0.0076
LorenzCoupled 0.0030 0.0045 0.0030 0.0075 0.0000 0.0127 0.0056 0.0565 0.0088 0.0088
ETTh1 0.0030 0.0092 0.0030 0.0015 0.0002 0.0425 0.0222 0.0156 0.0012 0.0175
Weather 0.0058 0.0010 0.0033 0.0053 0.0004 0.0061 0.0049 0.0172 0.0037 0.0039
Electricity OOM OOM 0.0003 0.0020 0.0000 0.0041 0.0124 0.0053 0.0017 0.0037
ETTh2 0.0040 0.0047 0.0040 0.0088 0.0011 0.1265 0.1307 0.0153 0.0036 0.0389
ETTm1 0.0023 0.0047 0.0023 0.0033 0.0000 0.0131 0.0367 0.0059 0.0023 0.0161
ETTm2 0.0010 0.0020 0.0010 0.0040 0.0004 0.0928 0.0364 0.0007 0.0022 0.1017
Traffic OOM OOM 0.0007 0.0025 0.0000 OOM OOM OOM OOM OOM
Table 10. In this table, all standard deviation values are averaged over the 4 used horizons for all datasets. For each horizon, the standard deviation results were acquired by running the model for 5 different seeds with the following values {3001,3002,3003,3004,3005}
. Note for some heavier models and datasets, we got Out of Memory error, specifically speaking some of the runs for electricity and traffic datasets.
Appendix F Detailed Granger Causality Results
We report results across three dataset groups (i) standard benchmark datasets, (ii) chaotic ODE datasets, and (iii) the large-scale Monash forecasting collection. For datasets with a large number of columns, we restrict the Granger causality analysis to pairwise comparisons among 20 randomly selected channels.
For the standard benchmark datasets such as electricity and traffic, Granger causality tests reveal that these datasets exhibit moderate inter-channel correlation, with higher causality (F-score) predominantly at shorter lookback windows (lag=30) (Table˜11Table˜12, and Table˜14). This suggests that, for short lags, information from other channels can slightly aid forecasting, but as the lag increases, most predictive power lies within each channel’s own history. This outcome aligns with the observation that univariate models (CI) can often perform well for these datasets at longer lookback settings, as there is diminishing marginal benefit from including other channels when window size grows.
In contrast, chaotic ODE datasets (Table˜11Table˜12, and Table˜14) consistently display higher Granger causality scores across all lag settings (with the exception of the electricity dataset), indicating robust and persistent dependencies between channels. This is particularly evident even at longer lookback windows. Such behavior is characteristic of deterministic physical systems or simulated multivariate dynamics where the evolution of one variable intrinsically influences the others. The implication is that multivariate models leveraging channel dependencies are especially suitable for these ODE datasets.
We also report results for the Monash repository’s (Godahewa et al., 2021) large-scale real-world datasets (over 40 time series from diverse domains). We select 40 datasets from a total of 58 datasets after removing datasets with missing values. Granger causality analysis Table˜13 reveals that only a minority—six of forty—show an inter-channel F-score above 10, and that F-scores tend to decline with increasing lag (except isolated cases like the Kaggle dataset). This pattern suggests that most Monash datasets lack strong lateral dependencies and instead behave like collections of independent univariate series. For instance, many resemble traffic datasets from the M1 or M3 collections: aggregations rather than integrated multivariate systems. As lags increase (e.g., at lag=196), additional channels yield minimal information gain, hence the F-score drops.
Overall, these findings highlight that Granger causality excels at disentangling the nature of multivariate time series dependencies. Standard benchmarks typically permit effective univariate treatment at larger lags, the ODE datasets consistently demand modeling of channel interactions, and most Monash datasets demonstrate little true multivariate dependence. These results underscore the value of focusing research on genuinely multivariate datasets, such as chaotic ODEs, to advance understanding and modeling of channel dependence in time series forecasting.
Dataset / Metric Average F-score (30) $H_{0}$ rejected (%, p=0.05)
Double Pendulum 39.60 100%
Lorenz 1378.54 83%
Lorenz Coupled 579.23 97%
Cell Cycle 6.78 50%
Blinking Rotlet 3.98 33%
HopField 2.86 80%
ETTm1 1.48 50%
ETTm2 1.31 30%
ETTh1 1.87 60%
ETTh2 1.94 70%
Weather 1.55 33%
Electricity* (Ch=20) 4.34 93%
Traffic* (Ch=20) 4.71 94%
Table 11. Average F-scores and percentage of $H_{0}$ rejected (p=0.05) across datasets (lag=30).
Dataset / Metric Average F-score (96) $H_{0}$ rejected (%, p=0.05)
Double Pendulum 10.21 100%
Lorenz 353.25 83%
Lorenz Coupled 160.95 66%
Cell Cycle 2.61 53%
Blinking Rotlet 2.59 33%
HopField 1.92 70%
ETTm1 1.13 33%
ETTm2 1.11 15%
ETTh1 1.12 20%
ETTh2 1.19 36%
Weather 1.12 30%
Electricity* (Ch=20) 1.92 84%
Traffic* (Ch=20) 2.22 94%
Table 12. Average F-scores and percentage of $H_{0}$ rejected (p=0.05) across datasets (lag=96).
Monash Dataset 7 30 96 192
australian electricity demand dataset 29.38 18.06 2.71 1.62
bitcoin without missing values 1.88 1.44 1.66 2.47
car parts without missing values 1.25
cif 2016 1.90
dominick 1.26
electricity hourly 32.55 4.66 1.96 1.21
electricity weekly 1.18 1.11
fred md 1.92 1.49 1.32 1.14
hospital 1.38
kaggle web traffic without missing values 3.02 3.25 12.96 45.63
kaggle web traffic weekly 1.14 1.56
kdd cup 2018 without missing values 2.48 1.53 1.18 1.08
london smart meters without missing values 1.69 1.26 1.07 1.03
m1 monthly 2.02
m1 quarterly
m1 yearly
m3 monthly 1.59
m3 other 1.46
m3 quarterly
m3 yearly
m4 daily 1.10 1.05
m4 hourly 18.06 3.48 1.96 1.20
m4 monthly 1.76
m4 quarterly 1.06
m4 weekly 1.14
m4 yearly
nn5 daily without missing values 9.45 2.10 1.17 1.02
nn5 weekly 1.17 1.03
pedestrian counts 10.85 3.14 2.65
rideshare without missing values 3.73 2.15 1.31
solar 10 minutes 8.63 3.87 1.52 1.56
solar weekly 0.78
temperature rain without missing values 2.85 1.54 1.20 1.06
tourism monthly 12.12 1.58
tourism quarterly 1.98
tourism yearly
traffic hourly 16.00 4.44 2.08 1.25
traffic weekly 1.12 1.21
vehicle trips without missing values 1.38
weather 1.22 1.17 1.12 1.03
Table 13. Average F-scores across 40 Monash datasets for lags 7, 30, 90, 192 .
Dataset / Metric Average F-score (192) $H_{0}$ rejected (%, p=0.05)
Double Pendulum 6.35 100%
Lorenz 126.38 83%
Lorenz Coupled 72.09 57%
Cell Cycle 1.58 53%
Blinking Rotlet 3.76 33%
HopField 1.55 50%
ETTm1 1.04 23%
ETTm2 0.99 0%
ETTh1 1.07 7%
ETTh2 1.06 17%
Weather 1.82 32%
Electricity* (Ch=20) 1.21 43%
Traffic* (Ch=20) 1.39 69%
Table 14. Average F-scores and percentage of rejected $H_{0}$ (p=0.05) across datasets (lag=192). Missing values indicate that the Granger Causality failed due to statistically limited number of available time points for a proper analysis.
Appendix G Full Lookback Window Analysis
We show here the detailed results of best performing lookback windows on seperate figures for each of the forecasting horizons. We can see from Figure˜4, Figure˜5, Figure˜6, and Figure˜7 the analysis of the results based on forecasting horizons 96, 192, 336, and 720 respectively. We can see that on average the same insights noticed from the discussed forecasting horizon 96, still applies to the rest of experiments. The patterns discussed with respect to ETTh1, and ETTh2 tending to use smaller lookback windows still applies clearly on horizons 192 and 336. This changes when 720 horizon is deployed which makes sense as the difficulty of forecasting such a large amount of timesteps promotes the usage of a larger lookback windows for all datasets.
Regarding the tendency of CI models to benefit more from a larger lookback window than CD models, that still holds for the different horizons as seen in the different figures. With also more models on the larger forecasting horizon, such as 720, tend to use larger lookback window more often. This pattern follows the same pattern we discussed for the performance on the ETTh datasets. The bottom line from this analysis is to show as well the complexity of forecasting different multivariate datasets across different horizons which makes the tuning of the lookback window clearly the standard approach to go for.
Figure 4. Analysis of Frequencies of Best Performing Lookback Windows Grouped by : (a) datasets, (b) CI models, and (c) CD models. All recorded for the forecasting horizon 96.
Figure 5. Analysis of Frequencies of Best Performing Lookback Windows Grouped by : (a) datasets, (b) CI models, and (c) CD models. All recorded for the forecasting horizon 192.
Figure 6. Analysis of Frequencies of Best Performing Lookback Windows Grouped by : (a) datasets, (b) CI models, and (c) CD models. All recorded for the forecasting horizon 336.
Figure 7. Analysis of Frequencies of Best Performing Lookback Windows Grouped by : (a) datasets, (b) CI models, and (c) CD models. All recorded for the forecasting horizon 720.
Appendix H ODE Datasets Sample Visualization
In this section, a visualization is shown for the rest of the 6 ODE datasets used in this research. In Figure˜8, a visualization of a 1000 timesteps from cell cycle based ODE is shown. Two different cyclin degradation activation are shown with channel 2 (middle top) activating the degradation of cyclin C1 levels (middle bottom). Same phenomena is shown clearly on the right 2 channels of the figure for cyclin C2 levels. In Figure˜9Figure˜10Figure˜11, and Figure˜12, we can see a visualization of the first 1000 time steps of the blinking rotlet, hopfield, lorenz and lorenz coupled time series based datasets. We can see clearly on these figures, how the dynamics of these different datasets generate complex yet real-world like time series data, with highs and lows from one channel triggering highs and lows in other channels.
Figure 8. Time series visualization of the cell cycle ODE-based time series sample.
Figure 9. Time series visualization of the blinking rotlet ODE-based time series sample.
Figure 10. Time series visualization of the hopfield ODE-based time series sample.
Figure 11. Time series visualization of the lorenz ODE-based time series sample.
Figure 12. Time series visualization of the Lorenz Coupled ODE-based time series sample.
Appendix I Cell Cycle ODE Discussion
We discussed already the interesting properties of Double Pendulum ODEs. In this part of the appendix, we would like to extend more the discussion on the coupling between the ODEs and how variables interact across time. Equations 1-6 presents the system of ODEs, we used to generate the cell cycle dataset (Romond et al., 1999). These 6 equations show two different oscillators based on the concentration levels of cyclin ($C_{1}$ and $C_{2}$) and how they motivate the progression of the cell cycle. The cyclin concentration reflects directly on the activated Cyclin Dependent Kinases (cdk1, cdk2) which represented in the equations by $M_{1}$ and $M_{2}$ respectively. It also affect directly enzymes $X_{1}$ and $X_{2}$. $v_{ij}$ and $v_{dj}$ are constants that are multiplied in equation 1 for $j=1$ and equation 4 for $j=2$ as a rate of cyclin synthesis. In equations 1, we can see clearly the coupling between the ODE variables $C_{1}$ and $M_{2}$. Whereas, in equation 4, we can see the coupling between $C_{2}$ and $M_{1}$. All of this shows how information from one channel can cause spikes or change of trend in other channels, showing clear channel dependence in the generated time series data. Also, this section is valuable to show the very important application of such equation on a real life biomedical process which elevates the importance of our study more. Please note that these equations were taken from (Romond et al., 1999) and for more details, we encourage the reader to go through their work as well.
(8) $\frac{dC_{1}}{dt}=\frac{v_{i1}K_{im1}}{K_{im1}+M_{2}}-\frac{v_{d1}X_{1}C_{1}}{K_{d1}+C_{1}}-k_{d1}C_{1}$
(9) $\frac{dM_{1}}{dt}=\frac{V_{1}\left(1-M_{1}\right)}{K_{1}+\left(1-M_{1}\right)}-\frac{V_{2}M_{1}}{K_{2}+M_{1}}$
(10) $\frac{dX_{1}}{dt}=\frac{V_{3}\left(1-X_{1}\right)}{K_{3}+\left(1-X_{1}\right)}-\frac{V_{4}X_{1}}{K_{4}+X_{1}}$
(11) $\frac{dC_{2}}{dt}=\frac{v_{i2}K_{im2}}{K_{im2}+M_{1}}-\frac{v_{d2}X_{2}C_{2}}{K_{d2}+C_{2}}-k_{d2}C_{2}$
(12) $\frac{dM_{2}}{dt}=\frac{U_{1}\left(1-M_{2}\right)}{H_{1}+\left(1-M_{2}\right)}-\frac{U_{2}M_{2}}{H_{2}+M_{2}}$
(13) $\frac{dX_{2}}{dt}=\frac{U_{3}\left(1-X_{2}\right)}{H_{3}+\left(1-X_{2}\right)}-\frac{U_{4}X_{2}}{H_{4}+X_{2}}$
(14) $V_{1}=\frac{C_{1}}{K_{c1}+C_{1}}V_{M_{1}},\qquad V_{3}=M_{1}\cdot V_{M_{3}}$
(15) $U_{1}=\frac{C_{2}}{K_{c2}+C_{2}}U_{M_{1}},\qquad U_{3}=M_{2}\cdot UM_{3}$
Appendix J Ablation Study on RevIN
In this section, we include an important ablation study where we test the importance of using Reversible Instance Normalization (RevIN)  (Kim et al., 2021) on 2 standard datasets and 2 datasets from the chaotic benchmark. Crossformer was originally proposed without RevIN, so we added the RevIN component to both the CI and CD versions to measure how much of an effect window size this can have on different types of datasets. Moreover we included one more strong baseline in PatchTST in the study measuring the effect of RevIN on its main CI version.
We got 2 interesting patterns that can be seen also from table  Table˜15 which are the following:
  • Introducing revin helps to improve the performance on the both variants of Crossformer on the standard datasets.
  • On the contrary on the chaotic ODEs the model without revin is performing better than its counterpart with revin.
These findings shows that in general as confirmed in the literature RevIN can help in the performance of different model on standard datasets. While this being not totally true for PatchTST, it has been true for most models in the literature  (Chen et al., 2023). Important as emphasized throughout this paper, that is not easy task to generalize to different types of datasets, and that is no different for RevIN as it deteriorates performance if introduced on Crossformer on the chaotic ODE datasets. This shows that the different nature of Chaotic ODEs with the complex patterns can not benefit from the normalization+denormalization technique introduced in RevIN.
Dataset / Model PatchTST_CI CrossFormer_CD CrossFormer_CI
with RevIN w/o RevIN with RevIN w/o RevIN with RevIN w/o RevIN
Hopfield 0.151 0.156 0.058 0.049 0.073 0.059
LorenzCoupled 0.689 0.696 0.453 0.276 0.515 0.372
ETTh2 0.298 0.287 0.322 0.537 0.318 0.397
ETTm1 0.314 0.313 0.339 0.364 0.297 0.309
Table 15. This table shows an ablation study on the reversible instance normalization technique  (Kim et al., 2021) which showed to improve performance for most model on the standard datasets. The best variant of model is in bold and best overall for each dataset is highlighted in blue. These results are based on the forecasting horizon 96.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
PatchTST CD-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
TSMixer CI-96 lr:0.0001,num_blocks:5,hidden_size:32,dropout:0.336,activation:relu,seq_len:512
TSMixer CD-96 lr:0.0010,num_blocks:1,hidden_size:64,dropout:0.033,activation:relu,seq_len:192
CrossFormer CI-96 lr: 0.0002, e_layers: 7, d_ff: 256, d_model: 512, dropout: 0.248, seg_len: 7, baseline: 0, cross_factor: 20, seq_len: 336
CrossFormer CD-96 lr: 0.0003, e_layers: 7, d_ff: 256, d_model: 512, dropout: 0.285, seg_len: 7, baseline: 0, cross_factor: 5, seq_len: 336
DLinear CI-96 lr: 0.0026, seq_len: 720
iTransformers CD-96 lr:1.60e-05, e_layers:9, d_ff:2048, d_model:1024, dropout:0.012, seq_len:192
TimeMixer CI-96 lr:0.0028, d_ff:256, d_model:128, e_layers:4, seq_len:192
TimeMixer CD-96 lr:0.0023, d_ff:256, d_model:128, e_layers:5, seq_len:192
PatchTST CI-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
PatchTST CD-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
TSMixer CI-192 lr:0.0005,num_blocks:8,hidden_size:1.02e+03,dropout:0.705,activation:relu,seq_len:512
TSMixer CD-192 lr:0.0003,num_blocks:1,hidden_size:64,dropout:0.205,activation:gelu,seq_len:512
CrossFormer CI-192 lr: 0.0003, e_layers: 7, d_ff: 256, d_model: 512, dropout: 0.312, seg_len: 7, baseline: 0, cross_factor: 18, seq_len: 336
CrossFormer CD-192 lr: 0.0003, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.266, seg_len: 7, baseline: 0, cross_factor: 7, seq_len: 192
DLinear CI-192 lr: 0.0055, seq_len: 512
iTransformers CD-192 lr:0.0006, e_layers:9, d_ff:2048, d_model:128, dropout:0.286, seq_len:720
TimeMixer CI-192 lr:0.0023, d_ff:256, d_model:128, e_layers:5, seq_len:192
TimeMixer CD-192 lr:0.0028, d_ff:256, d_model:128, e_layers:4, seq_len:192
PatchTST CI-336 lr’:1.96e-07, ’e_layers’:5, ’d_ff’: 256, ’d_model’:1024, ’dropout’:0.295,’fc_dropout’:0.622, ’patch_size’:8,’stride’:8,’seq_len’:720
PatchTST CD-336 lr’: 9.1666e-05, ’e_layers’: 6, ’d_ff’: 1024, ’d_model’: 1024, ’dropout’: 0.086, ’fc_dropout’: 0.5099, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
TSMixer CI-336 lr:0.0012,num_blocks:6,hidden_size:32,dropout:0.264,activation:relu,seq_len:512
TSMixer CD-336 lr:0.0002,num_blocks:5,hidden_size:64,dropout:0.287,activation:relu,seq_len:512
CrossFormer CI-336 lr: 0.0002, e_layers: 7, d_ff: 256, d_model: 256, dropout: 0.309, seg_len: 8, baseline: 0, cross_factor: 5, seq_len: 336
CrossFormer CD-336 lr: 0.0002, e_layers: 7, d_ff: 256, d_model: 256, dropout: 0.309, seg_len: 8, baseline: 0, cross_factor: 18, seq_len: 336
DLinear CI-336 lr: 0.0004, seq_len: 720
iTransformers CD-336 lr:0.0006, e_layers:9, d_ff:2048, d_model:128, dropout:0.286, seq_len:720
TimeMixer CI-336 lr:0.0028, d_ff:256, d_model:128, e_layers:4, seq_len:192
TimeMixer CD-336 lr:0.0010, d_ff:1024, d_model:1024, e_layers:10, seq_len:192
PatchTST CI-720 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
PatchTST CD-720 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
TSMixer CI-720 lr:0.0004,num_blocks:8,hidden_size:1.02e+03,dropout:0.754,activation:relu,seq_len:512
TSMixer CD-720 lr:0.0001,num_blocks:2,hidden_size:32,dropout:0.297,activation:relu,seq_len:512
CrossFormer CI-720 lr: 0.0002, e_layers: 7, d_ff: 512, d_model: 256, dropout: 0.317, seg_len: 7, baseline: 1, cross_factor: 17, seq_len: 96
CrossFormer CD-720 lr: 2.20e-05, e_layers: 6, d_ff: 256, d_model: 512, dropout: 0.085, seg_len: 12, baseline: 1, cross_factor: 11, seq_len: 96
DLinear CI-720 lr: 0.0042, seq_len: 512
iTransformers CD-720 lr:0.0001, e_layers:8, d_ff:2048, d_model:1024, dropout:0.356, seq_len:192
TimeMixer CI-720 lr:2.30e-05, d_ff:512, d_model:128, e_layers:1, seq_len:512
TimeMixer CD-720 lr:0.0021, d_ff:256, d_model:128, e_layers:4, seq_len:192
Table 16. Hyperparameter settings for the BlinkingRotlet dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:0.0003,e_layers:10,d_ff:256,d_model:128,dropout:0.081,fc_dropout:0.390,patch_size:8,stride:8,seq_len:336
PatchTST CD-96 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
TSMixer CI-96 lr:0.0055,num_blocks:10,hidden_size:256,dropout:0.197,activation:gelu,seq_len:512
TSMixer CD-96 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
CrossFormer CI-96 lr: 0.0002, e_layers: 7, d_ff: 256, d_model: 512, dropout: 0.248, seg_len: 7, baseline: 0, cross_factor: 5, seq_len: 336
CrossFormer CD-96 lr: 0.0002, e_layers: 7, d_ff: 256, d_model: 512, dropout: 0.248, seg_len: 7, baseline: 0, cross_factor: 5, seq_len: 336
DLinear CI-96 lr: 0.0012, seq_len: 336
iTransformers CD-96 lr:0.0001, e_layers:8, d_ff:2048, d_model:1024, dropout:0.356, seq_len:192
TimeMixer CI-96 lr:0.0027, d_ff:256, d_model:256, e_layers:8, seq_len:96
TimeMixer CD-96 lr:0.0017, d_ff:256, d_model:512, e_layers:4, seq_len:96
PatchTST CI-192 lr:0.0003,e_layers:10,d_ff:256,d_model:128,dropout:0.081,fc_dropout:0.390,patch_size:8,stride:8,seq_len:336
PatchTST CD-192 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
TSMixer CI-192 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
TSMixer CD-192 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
CrossFormer CI-192 lr: 0.0002, e_layers: 7, d_ff: 256, d_model: 512, dropout: 0.248, seg_len: 7, baseline: 0, cross_factor: 5, seq_len: 336
CrossFormer CD-192 lr: 0.0015, e_layers: 5, d_ff: 512, d_model: 128, dropout: 0.228, seg_len: 12, baseline: 1, cross_factor: 17, seq_len: 336
DLinear CI-192 lr: 0.0066, seq_len: 720
iTransformers CD-192 lr:1.30e-05, e_layers:4, d_ff:512, d_model:1024, dropout:0.440, seq_len:336
TimeMixer CI-192 lr:0.0009, d_ff:1024, d_model:1024, e_layers:10, seq_len:192
TimeMixer CD-192 lr:0.0021, d_ff:256, d_model:128, e_layers:4, seq_len:192
PatchTST CI-336 lr’: 0.00027, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 128, ’dropout’: 0.081, ’fc_dropout’: 0.390, ’patch_size’: 8, ’stride’: 8, ’seq_len’: 336
PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.53896, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
TSMixer CI-336 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
TSMixer CD-336 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
CrossFormer CI-336 lr: 0.0006, e_layers: 8, d_ff: 512, d_model: 128, dropout: 0.361, seg_len: 6, baseline: 1, cross_factor: 8, seq_len: 512
CrossFormer CD-336 lr: 0.0015, e_layers: 5, d_ff: 512, d_model: 128, dropout: 0.228, seg_len: 12, baseline: 1, cross_factor: 17, seq_len: 336
DLinear CI-336 lr: 0.0049, seq_len: 512
iTransformers CD-336 lr:0.0004, e_layers:1, d_ff:2048, d_model:1024, dropout:0.388, seq_len:192
TimeMixer CI-336 lr:0.0020, d_ff:512, d_model:256, e_layers:7, seq_len:192
TimeMixer CD-336 lr:0.0021, d_ff:512, d_model:1024, e_layers:8, seq_len:96
PatchTST CI-720 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
PatchTST CD-720 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
TSMixer CI-720 lr:0.0006,num_blocks:8,hidden_size:64,dropout:0.181,activation:relu,seq_len:192
TSMixer CD-720 lr:0.0010,num_blocks:1,hidden_size:64,dropout:0.033,activation:relu,seq_len:192
CrossFormer CI-720 lr: 0.0006, e_layers: 7, d_ff: 128, d_model: 256, dropout: 0.309, seg_len: 7, baseline: 1, cross_factor: 18, seq_len: 336
CrossFormer CD-720 lr: 0.0002, e_layers: 6, d_ff: 128, d_model: 256, dropout: 0.302, seg_len: 7, baseline: 1, cross_factor: 5, seq_len: 512
DLinear CI-720 lr: 0.0033, seq_len: 336
iTransformers CD-720 lr:0.0001, e_layers:8, d_ff:2048, d_model:1024, dropout:0.356, seq_len:192
TimeMixer CI-720 lr:0.0019, d_ff:512, d_model:512, e_layers:8, seq_len:96
TimeMixer CD-720 lr:0.0022, d_ff:1024, d_model:512, e_layers:8, seq_len:96
Table 17. Hyperparameter settings for the CellCycle dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
PatchTST CD-96 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
TSMixer CI-96 lr:0.0012,num_blocks:6,hidden_size:64,dropout:0.302,activation:relu,seq_len:192
TSMixer CD-96 lr:0.0010,num_blocks:1,hidden_size:64,dropout:0.033,activation:relu,seq_len:192
CrossFormer CI-96 lr: 0.0015, e_layers: 9, d_ff: 512, d_model: 128, dropout: 0.228, seg_len: 6, baseline: 0, cross_factor: 20, seq_len: 336
CrossFormer CD-96 lr: 0.0003, e_layers: 7, d_ff: 256, d_model: 512, dropout: 0.285, seg_len: 7, baseline: 0, cross_factor: 5, seq_len: 336
DLinear CI-96 lr: 0.0080, seq_len: 192
iTransformers CD-96 lr:0.0001, e_layers:8, d_ff:2048, d_model:1024, dropout:0.356, seq_len:192
TimeMixer CI-96 lr:0.0026, d_ff:256, d_model:256, e_layers:8, seq_len:192
TimeMixer CD-96 lr:0.0018, d_ff:1024, d_model:128, e_layers:6, seq_len:192
PatchTST CI-192 lr:1.53e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.112,fc_dropout:0.457,patch_size:16,stride:4,seq_len:192
PatchTST CD-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
TSMixer CI-192 lr:0.0012,num_blocks:4,hidden_size:64,dropout:0.323,activation:relu,seq_len:192
TSMixer CD-192 lr:0.0032,num_blocks:7,hidden_size:1.02e+03,dropout:0.006,activation:relu,seq_len:192
CrossFormer CI-192 lr: 0.0006, e_layers: 7, d_ff: 128, d_model: 256, dropout: 0.309, seg_len: 7, baseline: 1, cross_factor: 18, seq_len: 336
CrossFormer CD-192 lr: 0.0009, e_layers: 8, d_ff: 512, d_model: 128, dropout: 0.193, seg_len: 9, baseline: 0, cross_factor: 7, seq_len: 336
DLinear CI-192 lr: 0.0097, seq_len: 192
iTransformers CD-192 lr:0.0002, e_layers:8, d_ff:2048, d_model:512, dropout:0.114, seq_len:192
TimeMixer CI-192 lr:0.0027, d_ff:256, d_model:256, e_layers:8, seq_len:96
TimeMixer CD-192 lr:0.0006, d_ff:256, d_model:256, e_layers:7, seq_len:96
PatchTST CI-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.539, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.5389, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
TSMixer CI-336 lr:0.0012,num_blocks:8,hidden_size:1.02e+03,dropout:0.307,activation:relu,seq_len:192
TSMixer CD-336 lr:0.0010,num_blocks:1,hidden_size:64,dropout:0.033,activation:relu,seq_len:192
CrossFormer CI-336 lr: 0.0005, e_layers: 6, d_ff: 512, d_model: 256, dropout: 0.326, seg_len: 7, baseline: 1, cross_factor: 8, seq_len: 512
CrossFormer CD-336 lr: 0.0002, e_layers: 7, d_ff: 512, d_model: 128, dropout: 0.309, seg_len: 11, baseline: 0, cross_factor: 8, seq_len: 336
DLinear CI-336 lr: 0.0011, seq_len: 192
iTransformers CD-336 lr:0.0001, e_layers:8, d_ff:2048, d_model:1024, dropout:0.356, seq_len:192
TimeMixer CI-336 lr:0.0022, d_ff:1024, d_model:512, e_layers:8, seq_len:96
TimeMixer CD-336 lr:0.0019, d_ff:256, d_model:256, e_layers:8, seq_len:96
PatchTST CI-720 lr:0.0001,e_layers:9,d_ff:512,d_model:512,dropout:0.065,fc_dropout:0.539,patch_size:8,stride:4,seq_len:96
PatchTST CD-720 lr:0.0002,e_layers:1,d_ff:1.02e+03,d_model:512,dropout:0.471,fc_dropout:0.710,patch_size:8,stride:8,seq_len:336
TSMixer CI-720 lr:0.0005,num_blocks:7,hidden_size:1.02e+03,dropout:0.169,activation:relu,seq_len:192
TSMixer CD-720 lr:0.0010,num_blocks:1,hidden_size:64,dropout:0.033,activation:relu,seq_len:192
CrossFormer CI-720 lr: 0.0002, e_layers: 6, d_ff: 512, d_model: 128, dropout: 0.266, seg_len: 8, baseline: 1, cross_factor: 5, seq_len: 336
CrossFormer CD-720 lr: 0.0001, e_layers: 3, d_ff: 256, d_model: 512, dropout: 0.123, seg_len: 8, baseline: 0, cross_factor: 7, seq_len: 336
DLinear CI-720 lr: 0.0032, seq_len: 512
iTransformers CD-720 lr:8.00e-06, e_layers:2, d_ff:2048, d_model:1024, dropout:0.012, seq_len:192
TimeMixer CI-720 lr:0.0018, d_ff:512, d_model:512, e_layers:7, seq_len:96
TimeMixer CD-720 lr:0.0020, d_ff:256, d_model:256, e_layers:8, seq_len:96
Table 18. Hyperparameter settings for the DoublePendulum dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:0.0001,e_layers:1,d_ff:512,d_model:256,dropout:0.572,fc_dropout:0.631,patch_size:16,stride:4,seq_len:96
PatchTST CD-96 lr:0.0001,e_layers:3,d_ff:512,d_model:512,dropout:0.518,fc_dropout:0.802,patch_size:16,stride:4,seq_len:96
TSMixer CI-96 lr:0.0097,num_blocks:9,hidden_size:32,dropout:0.748,activation:gelu,seq_len:336
TSMixer CD-96 lr:0.0002,num_blocks:6,hidden_size:256,dropout:0.139,activation:gelu,seq_len:96
CrossFormer CI-96 lr: 0.0002, e_layers: 5, d_ff: 512, d_model: 256, dropout: 0.212, seg_len: 8, baseline: 0, cross_factor: 10, seq_len: 96
CrossFormer CD-96 lr: 0.0080, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.189, seg_len: 8, baseline: 0, cross_factor: 9, seq_len: 192
DLinear CI-96 lr: 0.0094, seq_len: 336
iTransformers CD-96 lr:0.0007, e_layers:3, d_ff:1024, d_model:512, dropout:0.539, seq_len:96
TimeMixer CI-96 lr:0.0004, d_ff:256, d_model:256, e_layers:6, seq_len:192
TimeMixer CD-96 lr:0.0002, d_ff:1024, d_model:512, e_layers:4, seq_len:512
PatchTST CI-192 lr:7.27e-06,e_layers:4,d_ff:1.02e+03,d_model:512,dropout:0.053,fc_dropout:0.183,patch_size:16,stride:4,seq_len:192
PatchTST CD-192 lr:0.0001,e_layers:3,d_ff:1.02e+03,d_model:128,dropout:0.705,fc_dropout:0.640,patch_size:16,stride:8,seq_len:512
TSMixer CI-192 lr:0.0005,num_blocks:1,hidden_size:64,dropout:0.173,activation:relu,seq_len:192
TSMixer CD-192 lr:0.0003,num_blocks:1,hidden_size:64,dropout:0.166,activation:relu,seq_len:192
CrossFormer CI-192 lr: 0.0080, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.189, seg_len: 8, baseline: 0, cross_factor: 9, seq_len: 192
CrossFormer CD-192 lr: 0.0080, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.189, seg_len: 8, baseline: 0, cross_factor: 9, seq_len: 192
DLinear CI-192 lr: 0.0010, seq_len: 336
iTransformers CD-192 lr:1.60e-05, e_layers:9, d_ff:2048, d_model:1024, dropout:0.012, seq_len:336
TimeMixer CI-192 lr:1.60e-05, d_ff:512, d_model:256, e_layers:3, seq_len:336
TimeMixer CD-192 lr:0.0001, d_ff:1024, d_model:256, e_layers:9, seq_len:512
PatchTST CI-336 lr’: 5.207e-05, ’e_layers’: 1, ’d_ff’: 512, ’d_model’: 256, ’dropout’: 0.572, ’fc_dropout’: 0.6308, ’patch_size’: 16, ’stride’: 4, ’seq_len’: 96
PatchTST CD-336 lr’: 2.920e-06, ’e_layers’: 6, ’d_ff’: 1024, ’d_model’: 512, ’dropout’: 0.342, ’fc_dropout’: 0.404, ’patch_size’: 16, ’stride’: 4, ’seq_len’: 192
TSMixer CI-336 lr:0.0002,num_blocks:2,hidden_size:64,dropout:0.205,activation:relu,seq_len:512
TSMixer CD-336 lr:0.0002,num_blocks:1,hidden_size:64,dropout:0.207,activation:relu,seq_len:192
CrossFormer CI-336 lr: 0.0080, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.189, seg_len: 8, baseline: 0, cross_factor: 9, seq_len: 192
CrossFormer CD-336 lr: 0.0080, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.189, seg_len: 8, baseline: 0, cross_factor: 9, seq_len: 192
DLinear CI-336 lr: 0.0011, seq_len: 336
iTransformers CD-336 lr:1.60e-05, e_layers:9, d_ff:2048, d_model:1024, dropout:0.012, seq_len:192
TimeMixer CI-336 lr:1.50e-05, d_ff:256, d_model:256, e_layers:2, seq_len:512
TimeMixer CD-336 lr:0.0009, d_ff:1024, d_model:512, e_layers:8, seq_len:512
PatchTST CI-720 lr:1.78e-06,e_layers:2,d_ff:512,d_model:512,dropout:0.438,fc_dropout:0.405,patch_size:16,stride:4,seq_len:336
PatchTST CD-720 lr:0.0024,e_layers:2,d_ff:1.02e+03,d_model:512,dropout:0.450,fc_dropout:0.374,patch_size:16,stride:4,seq_len:512
TSMixer CI-720 lr:0.0002,num_blocks:2,hidden_size:64,dropout:0.205,activation:relu,seq_len:512
TSMixer CD-720 lr:0.0003,num_blocks:1,hidden_size:64,dropout:0.169,activation:relu,seq_len:512
CrossFormer CI-720 lr: 0.0080, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.189, seg_len: 8, baseline: 0, cross_factor: 9, seq_len: 192
CrossFormer CD-720 lr: 0.0080, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.189, seg_len: 8, baseline: 0, cross_factor: 9, seq_len: 192
DLinear CI-720 lr: 0.0078, seq_len: 512
iTransformers CD-720 lr:1.30e-05, e_layers:8, d_ff:512, d_model:512, dropout:0.017, seq_len:336
TimeMixer CI-720 lr:1.00e-05, d_ff:512, d_model:512, e_layers:1, seq_len:720
TimeMixer CD-720 lr:0.0005, d_ff:512, d_model:256, e_layers:8, seq_len:512
Table 19. Hyperparameter settings for the ETTh1 dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:3.50e-05,e_layers:1,d_ff:512,d_model:256,dropout:0.571,fc_dropout:0.486,patch_size:16,stride:4,seq_len:96
PatchTST CD-96 lr:0.0001,e_layers:3,d_ff:512,d_model:512,dropout:0.688,fc_dropout:0.393,patch_size:16,stride:4,seq_len:96
TSMixer CI-96 lr:0.0005,num_blocks:1,hidden_size:64,dropout:0.166,activation:relu,seq_len:192
TSMixer CD-96 lr:0.0011,num_blocks:6,hidden_size:1.02e+03,dropout:0.746,activation:gelu,seq_len:192
CrossFormer CI-96 lr: 0.0003, e_layers: 5, d_ff: 512, d_model: 128, dropout: 0.707, seg_len: 3, baseline: 1, cross_factor: 10, seq_len: 96
CrossFormer CD-96 lr: 0.0076, e_layers: 2, d_ff: 512, d_model: 128, dropout: 0.494, seg_len: 5, baseline: 1, cross_factor: 13, seq_len: 96
DLinear CI-96 lr: 0.0010, seq_len: 336
iTransformers CD-96 lr:0.0001, e_layers:3, d_ff:1024, d_model:512, dropout:0.518, seq_len:96
TimeMixer CI-96 lr:0.0001, d_ff:256, d_model:256, e_layers:3, seq_len:192
TimeMixer CD-96 lr:0.0002, d_ff:256, d_model:256, e_layers:6, seq_len:192
PatchTST CI-192 lr:0.0003,e_layers:1,d_ff:256,d_model:128,dropout:0.553,fc_dropout:0.631,patch_size:8,stride:4,seq_len:336
PatchTST CD-192 lr:0.0002,e_layers:3,d_ff:512,d_model:512,dropout:0.688,fc_dropout:0.594,patch_size:16,stride:4,seq_len:96
TSMixer CI-192 lr:0.0097,num_blocks:5,hidden_size:32,dropout:0.114,activation:relu,seq_len:192
TSMixer CD-192 lr:0.0001,num_blocks:2,hidden_size:64,dropout:0.252,activation:relu,seq_len:192
CrossFormer CI-192 lr: 0.0076, e_layers: 2, d_ff: 512, d_model: 128, dropout: 0.494, seg_len: 5, baseline: 1, cross_factor: 13, seq_len: 96
CrossFormer CD-192 lr: 0.0076, e_layers: 2, d_ff: 512, d_model: 128, dropout: 0.494, seg_len: 5, baseline: 1, cross_factor: 13, seq_len: 96
DLinear CI-192 lr: 0.0034, seq_len: 336
iTransformers CD-192 lr:0.0001, e_layers:3, d_ff:1024, d_model:512, dropout:0.518, seq_len:96
TimeMixer CI-192 lr:0.0001, d_ff:256, d_model:256, e_layers:3, seq_len:192
TimeMixer CD-192 lr:0.0004, d_ff:256, d_model:256, e_layers:4, seq_len:96
PatchTST CI-336 lr’: 3.495e-05, ’e_layers’: 1, ’d_ff’: 512, ’d_model’: 256, ’dropout’: 0.5706, ’fc_dropout’: 0.486, ’patch_size’: 16, ’stride’: 4, ’seq_len’: 96
PatchTST CD-336 lr’: 0.0002, ’e_layers’: 3, ’d_ff’: 512, ’d_model’: 512, ’dropout’: 0.6877, ’fc_dropout’: 0.5936, ’patch_size’: 16, ’stride’: 4, ’seq_len’: 96
TSMixer CI-336 lr:0.0097,num_blocks:5,hidden_size:32,dropout:0.331,activation:relu,seq_len:192
TSMixer CD-336 lr:0.0082,num_blocks:1,hidden_size:256,dropout:0.155,activation:relu,seq_len:512
CrossFormer CI-336 lr: 0.0017, e_layers: 2, d_ff: 256, d_model: 512, dropout: 0.512, seg_len: 4, baseline: 1, cross_factor: 8, seq_len: 96
CrossFormer CD-336 lr: 0.0076, e_layers: 2, d_ff: 512, d_model: 128, dropout: 0.494, seg_len: 5, baseline: 1, cross_factor: 13, seq_len: 96
DLinear CI-336 lr: 0.0036, seq_len: 192
iTransformers CD-336 lr:0.0001, e_layers:3, d_ff:1024, d_model:512, dropout:0.518, seq_len:96
TimeMixer CI-336 lr:0.0072, d_ff:256, d_model:256, e_layers:8, seq_len:96
TimeMixer CD-336 lr:0.0011, d_ff:256, d_model:256, e_layers:4, seq_len:96
PatchTST CI-720 lr:0.0002,e_layers:3,d_ff:512,d_model:512,dropout:0.714,fc_dropout:0.650,patch_size:16,stride:4,seq_len:96
PatchTST CD-720 lr:0.0001,e_layers:3,d_ff:512,d_model:512,dropout:0.518,fc_dropout:0.802,patch_size:16,stride:4,seq_len:96
TSMixer CI-720 lr:0.0021,num_blocks:6,hidden_size:1.02e+03,dropout:0.141,activation:relu,seq_len:192
TSMixer CD-720 lr:0.0010,num_blocks:1,hidden_size:64,dropout:0.033,activation:relu,seq_len:192
CrossFormer CI-720 lr: 0.0083, e_layers: 2, d_ff: 512, d_model: 128, dropout: 0.646, seg_len: 6, baseline: 1, cross_factor: 17, seq_len: 96
CrossFormer CD-720 lr: 0.0076, e_layers: 2, d_ff: 512, d_model: 128, dropout: 0.494, seg_len: 5, baseline: 1, cross_factor: 13, seq_len: 96
DLinear CI-720 lr: 0.0093, seq_len: 96
iTransformers CD-720 lr:0.0001, e_layers:3, d_ff:1024, d_model:512, dropout:0.518, seq_len:96
TimeMixer CI-720 lr:0.0001, d_ff:512, d_model:128, e_layers:3, seq_len:192
TimeMixer CD-720 lr:0.0033, d_ff:512, d_model:128, e_layers:8, seq_len:96
Table 20. Hyperparameter settings for the ETTh2 dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
PatchTST CD-96 lr:6.07e-06,e_layers:3,d_ff:256,d_model:1.02e+03,dropout:0.006,fc_dropout:0.642,patch_size:16,stride:8,seq_len:192
TSMixer CI-96 lr:0.0097,num_blocks:2,hidden_size:1.02e+03,dropout:0.115,activation:relu,seq_len:192
TSMixer CD-96 lr:0.0084,num_blocks:2,hidden_size:32,dropout:0.264,activation:relu,seq_len:512
CrossFormer CI-96 lr: 0.0002, e_layers: 6, d_ff: 512, d_model: 128, dropout: 0.266, seg_len: 8, baseline: 1, cross_factor: 5, seq_len: 336
CrossFormer CD-96 lr: 0.0001, e_layers: 6, d_ff: 512, d_model: 256, dropout: 0.272, seg_len: 7, baseline: 1, cross_factor: 17, seq_len: 96
DLinear CI-96 lr: 0.0006, seq_len: 336
iTransformers CD-96 lr:0.0005, e_layers:5, d_ff:512, d_model:128, dropout:0.462, seq_len:192
TimeMixer CI-96 lr:0.0004, d_ff:256, d_model:128, e_layers:2, seq_len:192
TimeMixer CD-96 lr:0.0009, d_ff:512, d_model:512, e_layers:8, seq_len:192
PatchTST CI-192 lr:2.97e-06,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.081,fc_dropout:0.435,patch_size:16,stride:8,seq_len:192
PatchTST CD-192 lr:2.97e-06,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.081,fc_dropout:0.435,patch_size:16,stride:8,seq_len:192
TSMixer CI-192 lr:0.0097,num_blocks:2,hidden_size:1.02e+03,dropout:0.115,activation:relu,seq_len:192
TSMixer CD-192 lr:0.0005,num_blocks:6,hidden_size:64,dropout:0.280,activation:relu,seq_len:192
CrossFormer CI-192 lr: 0.0002, e_layers: 7, d_ff: 512, d_model: 256, dropout: 0.317, seg_len: 7, baseline: 1, cross_factor: 17, seq_len: 96
CrossFormer CD-192 lr: 0.0002, e_layers: 7, d_ff: 512, d_model: 256, dropout: 0.317, seg_len: 7, baseline: 1, cross_factor: 17, seq_len: 96
DLinear CI-192 lr: 0.0010, seq_len: 336
iTransformers CD-192 lr:1.30e-05, e_layers:4, d_ff:512, d_model:1024, dropout:0.440, seq_len:336
TimeMixer CI-192 lr:0.0002, d_ff:256, d_model:256, e_layers:4, seq_len:192
TimeMixer CD-192 lr:0.0002, d_ff:512, d_model:1024, e_layers:7, seq_len:192
PatchTST CI-336 lr’: 0.0003, ’e_layers’: 9, ’d_ff’: 256, ’d_model’: 128, ’dropout’: 0.228, ’fc_dropout’: 0.0716, ’patch_size’: 8, ’stride’: 8, ’seq_len’: 336
PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.5389, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
TSMixer CI-336 lr:0.0006,num_blocks:8,hidden_size:1.02e+03,dropout:0.166,activation:relu,seq_len:192
TSMixer CD-336 lr:0.0004,num_blocks:8,hidden_size:64,dropout:0.215,activation:relu,seq_len:192
CrossFormer CI-336 lr: 0.0003, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.265, seg_len: 8, baseline: 0, cross_factor: 9, seq_len: 336
CrossFormer CD-336 lr: 0.0080, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.189, seg_len: 8, baseline: 0, cross_factor: 9, seq_len: 192
DLinear CI-336 lr: 0.0007, seq_len: 336
iTransformers CD-336 lr:0.0004, e_layers:1, d_ff:2048, d_model:1024, dropout:0.388, seq_len:192
TimeMixer CI-336 lr:0.0002, d_ff:256, d_model:256, e_layers:4, seq_len:192
TimeMixer CD-336 lr:0.0021, d_ff:256, d_model:128, e_layers:4, seq_len:192
PatchTST CI-720 lr:2.27e-06,e_layers:4,d_ff:1.02e+03,d_model:128,dropout:0.198,fc_dropout:0.356,patch_size:16,stride:8,seq_len:512
PatchTST CD-720 lr:0.0019,e_layers:7,d_ff:1.02e+03,d_model:512,dropout:0.471,fc_dropout:0.375,patch_size:8,stride:8,seq_len:720
TSMixer CI-720 lr:0.0097,num_blocks:2,hidden_size:1.02e+03,dropout:0.115,activation:relu,seq_len:192
TSMixer CD-720 lr:0.0011,num_blocks:6,hidden_size:1.02e+03,dropout:0.339,activation:relu,seq_len:192
CrossFormer CI-720 lr: 0.0001, e_layers: 6, d_ff: 128, d_model: 256, dropout: 0.272, seg_len: 7, baseline: 1, cross_factor: 20, seq_len: 336
CrossFormer CD-720 lr: 0.0080, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.189, seg_len: 8, baseline: 0, cross_factor: 9, seq_len: 192
DLinear CI-720 lr: 0.0097, seq_len: 192
iTransformers CD-720 lr:0.0001, e_layers:8, d_ff:2048, d_model:128, dropout:0.319, seq_len:720
TimeMixer CI-720 lr:1.00e-05, d_ff:512, d_model:512, e_layers:1, seq_len:720
TimeMixer CD-720 lr:0.0001, d_ff:1024, d_model:1024, e_layers:2, seq_len:192
Table 21. Hyperparameter settings for the ETTm1 dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:0.0002,e_layers:7,d_ff:1.02e+03,d_model:1.02e+03,dropout:0.648,fc_dropout:0.175,patch_size:8,stride:4,seq_len:512
PatchTST CD-96 lr:1.60e-06,e_layers:8,d_ff:256,d_model:1.02e+03,dropout:0.307,fc_dropout:0.623,patch_size:8,stride:4,seq_len:512
TSMixer CI-96 lr:0.0097,num_blocks:5,hidden_size:32,dropout:0.289,activation:relu,seq_len:512
TSMixer CD-96 lr:0.0003,num_blocks:1,hidden_size:64,dropout:0.077,activation:relu,seq_len:192
CrossFormer CI-96 lr: 0.0003, e_layers: 6, d_ff: 512, d_model: 128, dropout: 0.298, seg_len: 7, baseline: 1, cross_factor: 5, seq_len: 336
CrossFormer CD-96 lr: 0.0003, e_layers: 4, d_ff: 512, d_model: 128, dropout: 0.382, seg_len: 11, baseline: 1, cross_factor: 11, seq_len: 512
DLinear CI-96 lr: 0.0020, seq_len: 720
iTransformers CD-96 lr:1.20e-05, e_layers:2, d_ff:1024, d_model:512, dropout:0.083, seq_len:512
TimeMixer CI-96 lr:1.10e-05, d_ff:512, d_model:256, e_layers:4, seq_len:336
TimeMixer CD-96 lr:5.00e-06, d_ff:512, d_model:512, e_layers:3, seq_len:720
PatchTST CI-192 lr:7.02e-07,e_layers:2,d_ff:1.02e+03,d_model:512,dropout:0.292,fc_dropout:0.673,patch_size:8,stride:4,seq_len:336
PatchTST CD-192 lr:1.82e-06,e_layers:4,d_ff:1.02e+03,d_model:1.02e+03,dropout:0.471,fc_dropout:0.405,patch_size:16,stride:8,seq_len:512
TSMixer CI-192 lr:0.0010,num_blocks:1,hidden_size:64,dropout:0.033,activation:relu,seq_len:192
TSMixer CD-192 lr:0.0004,num_blocks:1,hidden_size:64,dropout:0.077,activation:relu,seq_len:192
CrossFormer CI-192 lr: 0.0001, e_layers: 3, d_ff: 512, d_model: 128, dropout: 0.180, seg_len: 9, baseline: 1, cross_factor: 14, seq_len: 720
CrossFormer CD-192 lr: 0.0001, e_layers: 3, d_ff: 512, d_model: 128, dropout: 0.180, seg_len: 9, baseline: 1, cross_factor: 14, seq_len: 720
DLinear CI-192 lr: 0.0002, seq_len: 512
iTransformers CD-192 lr:1.40e-05, e_layers:1, d_ff:1024, d_model:512, dropout:0.083, seq_len:512
TimeMixer CI-192 lr:1.00e-05, d_ff:512, d_model:512, e_layers:1, seq_len:720
TimeMixer CD-192 lr:0.0003, d_ff:512, d_model:128, e_layers:4, seq_len:720
PatchTST CI-336 lr’: 0.0002, ’e_layers’: 7, ’d_ff’: 1024, ’d_model’: 1024, ’dropout’: 0.6482, ’fc_dropout’: 0.1748, ’patch_size’: 8, ’stride’: 4, ’seq_len’: 512
PatchTST CD-336 lr’: 1.602e-06, ’e_layers’: 8, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.307, ’fc_dropout’: 0.6229, ’patch_size’: 8, ’stride’: 4, ’seq_len’: 512
TSMixer CI-336 lr:0.0097,num_blocks:5,hidden_size:32,dropout:0.328,activation:relu,seq_len:512
TSMixer CD-336 lr:0.0087,num_blocks:6,hidden_size:64,dropout:0.759,activation:gelu,seq_len:512
CrossFormer CI-336 lr: 0.0026, e_layers: 1, d_ff: 512, d_model: 128, dropout: 0.897, seg_len: 6, baseline: 1, cross_factor: 5, seq_len: 336
CrossFormer CD-336 lr: 0.0001, e_layers: 6, d_ff: 512, d_model: 256, dropout: 0.679, seg_len: 10, baseline: 1, cross_factor: 17, seq_len: 720
DLinear CI-336 lr: 0.0038, seq_len: 720
iTransformers CD-336 lr:8.00e-06, e_layers:6, d_ff:2048, d_model:128, dropout:0.184, seq_len:512
TimeMixer CI-336 lr:1.00e-05, d_ff:512, d_model:512, e_layers:1, seq_len:720
TimeMixer CD-336 lr:1.00e-05, d_ff:512, d_model:128, e_layers:1, seq_len:720
PatchTST CI-720 lr:2.27e-06,e_layers:4,d_ff:1.02e+03,d_model:128,dropout:0.198,fc_dropout:0.356,patch_size:16,stride:8,seq_len:512
PatchTST CD-720 lr:0.0032,e_layers:3,d_ff:1.02e+03,d_model:128,dropout:0.318,fc_dropout:0.402,patch_size:16,stride:8,seq_len:512
TSMixer CI-720 lr:0.0097,num_blocks:5,hidden_size:32,dropout:0.289,activation:relu,seq_len:512
TSMixer CD-720 lr:0.0086,num_blocks:9,hidden_size:32,dropout:0.759,activation:relu,seq_len:512
CrossFormer CI-720 lr: 2.75e-05, e_layers: 8, d_ff: 512, d_model: 256, dropout: 0.895, seg_len: 10, baseline: 1, cross_factor: 12, seq_len: 720
CrossFormer CD-720 lr: 0.0028, e_layers: 8, d_ff: 512, d_model: 256, dropout: 0.664, seg_len: 10, baseline: 0, cross_factor: 16, seq_len: 512
DLinear CI-720 lr: 0.0022, seq_len: 336
iTransformers CD-720 lr:9.00e-06, e_layers:2, d_ff:2048, d_model:128, dropout:0.280, seq_len:512
TimeMixer CI-720 lr:1.40e-05, d_ff:512, d_model:256, e_layers:4, seq_len:336
TimeMixer CD-720 lr:1.00e-05, d_ff:512, d_model:128, e_layers:1, seq_len:720
Table 22. Hyperparameter settings for the ETTm2 dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
PatchTST CD-96 lr:0.0003,e_layers:10,d_ff:256,d_model:128,dropout:0.081,fc_dropout:0.390,patch_size:8,stride:8,seq_len:336
TSMixer CI-96 lr:0.0002,num_blocks:6,hidden_size:256,dropout:0.334,activation:relu,seq_len:512
TSMixer CD-96 lr:0.0097,num_blocks:5,hidden_size:32,dropout:0.311,activation:relu,seq_len:512
CrossFormer CI-96 lr: 0.0003, e_layers: 7, d_ff: 256, d_model: 512, dropout: 0.336, seg_len: 6, baseline: 0, cross_factor: 17, seq_len: 512
CrossFormer CD-96 lr: 0.0001, e_layers: 3, d_ff: 256, d_model: 512, dropout: 0.123, seg_len: 8, baseline: 0, cross_factor: 7, seq_len: 336
DLinear CI-96 lr: 0.0078, seq_len: 720
iTransformers CD-96 lr:0.0002, e_layers:8, d_ff:2048, d_model:512, dropout:0.114, seq_len:192
TimeMixer CI-96 lr:0.0015, d_ff:256, d_model:256, e_layers:7, seq_len:192
TimeMixer CD-96 lr:0.0023, d_ff:256, d_model:128, e_layers:5, seq_len:192
PatchTST CI-192 lr:0.0003,e_layers:10,d_ff:256,d_model:128,dropout:0.081,fc_dropout:0.390,patch_size:8,stride:8,seq_len:336
PatchTST CD-192 lr:0.0003,e_layers:10,d_ff:256,d_model:128,dropout:0.081,fc_dropout:0.390,patch_size:8,stride:8,seq_len:336
TSMixer CI-192 lr:0.0002,num_blocks:8,hidden_size:64,dropout:0.295,activation:relu,seq_len:512
TSMixer CD-192 lr:0.0010,num_blocks:1,hidden_size:64,dropout:0.033,activation:relu,seq_len:192
CrossFormer CI-192 lr: 0.0002, e_layers: 7, d_ff: 256, d_model: 512, dropout: 0.248, seg_len: 7, baseline: 0, cross_factor: 20, seq_len: 336
CrossFormer CD-192 lr: 0.0002, e_layers: 7, d_ff: 256, d_model: 512, dropout: 0.248, seg_len: 7, baseline: 0, cross_factor: 20, seq_len: 512
DLinear CI-192 lr: 0.0082, seq_len: 512
iTransformers CD-192 lr:0.0002, e_layers:8, d_ff:2048, d_model:512, dropout:0.114, seq_len:192
TimeMixer CI-192 lr:0.0023, d_ff:256, d_model:128, e_layers:5, seq_len:192
TimeMixer CD-192 lr:0.0010, d_ff:1024, d_model:1024, e_layers:10, seq_len:192
PatchTST CI-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.014, ’fc_dropout’: 0.539, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
PatchTST CD-336 lr:0.0004,num_blocks:8,hidden_size:1.02e+03,dropout:0.159,activation:relu,seq_len:512
TSMixer CI-336 lr:0.0011,num_blocks:1,hidden_size:256,dropout:0.322,activation:relu,seq_len:512
TSMixer CD-336 lr:0.0011,num_blocks:1,hidden_size:256,dropout:0.322,activation:relu,seq_len:512
CrossFormer CI-336 lr: 0.0006, e_layers: 7, d_ff: 128, d_model: 256, dropout: 0.309, seg_len: 7, baseline: 1, cross_factor: 18, seq_len: 336
CrossFormer CD-336 lr: 0.0011, e_layers: 8, d_ff: 512, d_model: 128, dropout: 0.357, seg_len: 7, baseline: 1, cross_factor: 8, seq_len: 512
DLinear CI-336 lr: 0.0002, seq_len: 512
iTransformers CD-336 lr:0.0001, e_layers:8, d_ff:2048, d_model:1024, dropout:0.356, seq_len:192
TimeMixer CI-336 lr:0.0041, d_ff:256, d_model:1024, e_layers:6, seq_len:96
TimeMixer CD-336 lr:0.0018, d_ff:1024, d_model:128, e_layers:6, seq_len:192
PatchTST CI-720 lr:0.0003,e_layers:10,d_ff:256,d_model:128,dropout:0.081,fc_dropout:0.390,patch_size:8,stride:8,seq_len:336
PatchTST CD-720 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
TSMixer CI-720 lr:0.0002,num_blocks:8,hidden_size:1.02e+03,dropout:0.205,activation:relu,seq_len:512
TSMixer CD-720 lr:0.0010,num_blocks:1,hidden_size:64,dropout:0.033,activation:relu,seq_len:192
CrossFormer CI-720 lr: 0.0001, e_layers: 3, d_ff: 256, d_model: 512, dropout: 0.123, seg_len: 8, baseline: 0, cross_factor: 7, seq_len: 336
CrossFormer CD-720 lr: 0.0002, e_layers: 6, d_ff: 512, d_model: 128, dropout: 0.266, seg_len: 8, baseline: 1, cross_factor: 5, seq_len: 336
DLinear CI-720 lr: 0.0034, seq_len: 512
iTransformers CD-720 lr:0.0001, e_layers:8, d_ff:2048, d_model:1024, dropout:0.356, seq_len:192
TimeMixer CI-720 lr:0.0021, d_ff:256, d_model:256, e_layers:8, seq_len:512
TimeMixer CD-720 lr:0.0010, d_ff:1024, d_model:1024, e_layers:10, seq_len:192
Table 23. Hyperparameter settings for the Hopfield dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
PatchTST CD-96 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
TSMixer CI-96 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
TSMixer CD-96 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
CrossFormer CI-96 lr: 0.0002, e_layers: 7, d_ff: 256, d_model: 512, dropout: 0.309, seg_len: 3, baseline: 0, cross_factor: 18, seq_len: 336
CrossFormer CD-96 lr: 0.0004, e_layers: 6, d_ff: 512, d_model: 256, dropout: 0.228, seg_len: 7, baseline: 0, cross_factor: 8, seq_len: 336
DLinear CI-96 lr: 0.0093, seq_len: 96
iTransformers CD-96 lr:0.0001, e_layers:8, d_ff:2048, d_model:1024, dropout:0.356, seq_len:192
TimeMixer CI-96 lr:0.0028, d_ff:256, d_model:128, e_layers:4, seq_len:192
TimeMixer CD-96 lr:0.0028, d_ff:256, d_model:128, e_layers:4, seq_len:192
PatchTST CI-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
PatchTST CD-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
TSMixer CI-192 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
TSMixer CD-192 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
CrossFormer CI-192 lr: 0.0004, e_layers: 8, d_ff: 256, d_model: 256, dropout: 0.323, seg_len: 10, baseline: 0, cross_factor: 10, seq_len: 336
CrossFormer CD-192 lr: 0.0002, e_layers: 6, d_ff: 128, d_model: 256, dropout: 0.302, seg_len: 7, baseline: 1, cross_factor: 5, seq_len: 336
DLinear CI-192 lr: 0.0093, seq_len: 96
iTransformers CD-192 lr:0.0011, e_layers:4, d_ff:2048, d_model:128, dropout:0.255, seq_len:720
TimeMixer CI-192 lr:0.0013, d_ff:512, d_model:1024, e_layers:8, seq_len:96
TimeMixer CD-192 lr:0.0016, d_ff:256, d_model:128, e_layers:5, seq_len:192
PatchTST CI-336 lr’: 0.0003, ’e_layers’: 9, ’d_ff’: 256, ’d_model’: 128, ’dropout’: 0.2277, ’fc_dropout’: 0.0716, ’patch_size’: 8, ’stride’: 8, ’seq_len’: 336
PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.5389, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
TSMixer CI-336 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
TSMixer CD-336 lr:0.0080,num_blocks:7,hidden_size:64,dropout:0.204,activation:relu,seq_len:192
CrossFormer CI-336 lr: 0.0003, e_layers: 7, d_ff: 256, d_model: 512, dropout: 0.285, seg_len: 7, baseline: 0, cross_factor: 5, seq_len: 336
CrossFormer CD-336 lr: 0.0003, e_layers: 5, d_ff: 512, d_model: 256, dropout: 0.228, seg_len: 6, baseline: 0, cross_factor: 17, seq_len: 336
DLinear CI-336 lr: 0.0093, seq_len: 96
iTransformers CD-336 lr:0.0011, e_layers:4, d_ff:2048, d_model:128, dropout:0.255, seq_len:720
TimeMixer CI-336 lr:0.0010, d_ff:1024, d_model:1024, e_layers:10, seq_len:192
TimeMixer CD-336 lr:0.0009, d_ff:1024, d_model:256, e_layers:8, seq_len:192
PatchTST CI-720 lr:0.0001,e_layers:3,d_ff:512,d_model:512,dropout:0.518,fc_dropout:0.802,patch_size:16,stride:4,seq_len:96
PatchTST CD-720 lr:0.0001,e_layers:3,d_ff:512,d_model:512,dropout:0.518,fc_dropout:0.802,patch_size:16,stride:4,seq_len:96
TSMixer CI-720 lr:0.0026,num_blocks:6,hidden_size:256,dropout:0.126,activation:gelu,seq_len:96
TSMixer CD-720 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
CrossFormer CI-720 lr: 0.0006, e_layers: 7, d_ff: 128, d_model: 256, dropout: 0.309, seg_len: 7, baseline: 1, cross_factor: 18, seq_len: 336
CrossFormer CD-720 lr: 0.0006, e_layers: 8, d_ff: 512, d_model: 128, dropout: 0.361, seg_len: 6, baseline: 1, cross_factor: 10, seq_len: 336
DLinear CI-720 enc_in:3, lr:0.0027, seq_len:96
iTransformers CD-720 lr:0.0002, e_layers:7, d_ff:1024, d_model:128, dropout:0.480, seq_len:720
TimeMixer CI-720 lr:0.0010, d_ff:1024, d_model:1024, e_layers:10, seq_len:192
TimeMixer CD-720 lr:0.0041, d_ff:256, d_model:1024, e_layers:6, seq_len:96
Table 24. Hyperparameter settings for the Lorenz dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
PatchTST CD-96 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
TSMixer CI-96 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
TSMixer CD-96 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
CrossFormer CI-96 lr: 0.0003, e_layers: 7, d_ff: 256, d_model: 512, dropout: 0.336, seg_len: 10, baseline: 0, cross_factor: 5, seq_len: 336
CrossFormer CD-96 lr: 0.0004, e_layers: 7, d_ff: 512, d_model: 256, dropout: 0.323, seg_len: 6, baseline: 0, cross_factor: 8, seq_len: 336
DLinear CI-96 lr: 0.0097, seq_len: 192
iTransformers CD-96 lr:1.30e-05, e_layers:4, d_ff:2048, d_model:128, dropout:0.008, seq_len:720
TimeMixer CI-96 lr:0.0014, d_ff:1024, d_model:256, e_layers:8, seq_len:96
TimeMixer CD-96 lr:0.0006, d_ff:1024, d_model:128, e_layers:5, seq_len:192
PatchTST CI-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
PatchTST CD-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
TSMixer CI-192 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
TSMixer CD-192 lr:0.0097,num_blocks:6,hidden_size:256,dropout:0.254,activation:relu,seq_len:512
CrossFormer CI-192 lr: 0.0003, e_layers: 6, d_ff: 256, d_model: 512, dropout: 0.309, seg_len: 8, baseline: 0, cross_factor: 5, seq_len: 336
CrossFormer CD-192 lr: 0.0003, e_layers: 6, d_ff: 512, d_model: 128, dropout: 0.298, seg_len: 7, baseline: 1, cross_factor: 5, seq_len: 336
DLinear CI-192 lr: 0.0080, seq_len: 192
iTransformers CD-192 lr:1.30e-05, e_layers:4, d_ff:2048, d_model:128, dropout:0.008, seq_len:720
TimeMixer CI-192 lr:0.0011, d_ff:256, d_model:256, e_layers:7, seq_len:96
TimeMixer CD-192 lr:0.0024, d_ff:256, d_model:512, e_layers:6, seq_len:96
PatchTST CI-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.539, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.5389, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
TSMixer CI-336 lr:0.0041,num_blocks:10,hidden_size:256,dropout:0.415,activation:gelu,seq_len:96
TSMixer CD-336 lr:0.0010,num_blocks:1,hidden_size:64,dropout:0.033,activation:relu,seq_len:192
CrossFormer CI-336 lr: 0.0001, e_layers: 8, d_ff: 256, d_model: 256, dropout: 0.112, seg_len: 10, baseline: 0, cross_factor: 7, seq_len: 336
CrossFormer CD-336 lr: 0.0002, e_layers: 6, d_ff: 512, d_model: 128, dropout: 0.266, seg_len: 8, baseline: 1, cross_factor: 5, seq_len: 336
DLinear CI-336 lr: 0.0093, seq_len: 96
iTransformers CD-336 lr:0.0003, e_layers:1, d_ff:512, d_model:512, dropout:0.270, seq_len:720
TimeMixer CI-336 lr:0.0015, d_ff:512, d_model:256, e_layers:8, seq_len:96
TimeMixer CD-336 lr:0.0011, d_ff:256, d_model:256, e_layers:7, seq_len:96
PatchTST CI-720 lr:0.0001,e_layers:2,d_ff:1.02e+03,d_model:1.02e+03,dropout:0.463,fc_dropout:0.708,patch_size:16,stride:4,seq_len:192
PatchTST CD-720 lr:0.0018,e_layers:5,d_ff:1.02e+03,d_model:128,dropout:0.097,fc_dropout:0.623,patch_size:16,stride:8,seq_len:512
TSMixer CI-720 lr:0.0005,num_blocks:6,hidden_size:64,dropout:0.280,activation:relu,seq_len:192
TSMixer CD-720 lr:0.0097,num_blocks:6,hidden_size:256,dropout:0.254,activation:relu,seq_len:512
CrossFormer CI-720 lr: 0.0002, e_layers: 6, d_ff: 128, d_model: 256, dropout: 0.302, seg_len: 7, baseline: 1, cross_factor: 5, seq_len: 336
CrossFormer CD-720 lr: 0.0001, e_layers: 3, d_ff: 256, d_model: 512, dropout: 0.123, seg_len: 8, baseline: 0, cross_factor: 7, seq_len: 336
DLinear CI-720 lr: 0.0010, seq_len: 96
iTransformers CD-720 lr:0.0003, e_layers:5, d_ff:512, d_model:128, dropout:0.305, seq_len:720
TimeMixer CI-720 lr:0.0005, d_ff:256, d_model:256, e_layers:8, seq_len:96
TimeMixer CD-720 lr:0.0041, d_ff:256, d_model:1024, e_layers:6, seq_len:96
Table 25. Hyperparameter settings for the LorenzCoupled dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96
PatchTST CD-96
TSMixer CI-96 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
TSMixer CD-96 lr:0.0003,num_blocks:1,hidden_size:64,dropout:0.166,activation:relu,seq_len:512
CrossFormer CI-96 lr: 0.0024, e_layers: 1, d_ff: 128, d_model: 256, dropout: 0.305, seg_len: 6, baseline: 0, cross_factor: 7, seq_len: 192
CrossFormer CD-96 lr: 0.0032, e_layers: 1, d_ff: 128, d_model: 256, dropout: 0.294, seg_len: 7, baseline: 0, cross_factor: 8, seq_len: 192
DLinear CI-96 lr: 0.0079, seq_len: 512
iTransformers CD-96 lr:0.0032, e_layers:4, d_ff:1024, d_model:128, dropout:0.139, seq_len:720
TimeMixer CI-96 lr:0.0010, d_ff:512, d_model:256, e_layers:1, seq_len:720
TimeMixer CD-96 lr:0.0005, d_ff:512, d_model:256, e_layers:8, seq_len:512
PatchTST CI-192
PatchTST CD-192
TSMixer CI-192 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
TSMixer CD-192 lr:0.0097,num_blocks:6,hidden_size:32,dropout:0.315,activation:relu,seq_len:512
CrossFormer CI-192 OOM
CrossFormer CD-192 lr: 0.0080, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.189, seg_len: 8, baseline: 0, cross_factor: 9, seq_len: 192
DLinear CI-192 lr: 0.0058, seq_len: 512
iTransformers CD-192 lr:0.0004, e_layers:1, d_ff:2048, d_model:1024, dropout:0.080, seq_len:192
TimeMixer CI-192 lr 0.00008, d_ff 512, d_model 512, e_layers 4, seq_len 192
TimeMixer CD-192 lr:0.0015, d_ff:256, d_model:256, e_layers:7, seq_len:512
PatchTST CI-336 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
PatchTST CD-336
TSMixer CI-336
TSMixer CD-336 lr:0.0097,num_blocks:5,hidden_size:32,dropout:0.336,activation:relu,seq_len:512
CrossFormer CI-336 lr: 0.0080, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.189, seg_len: 8, baseline: 0, cross_factor: 9, seq_len: 192
CrossFormer CD-336 lr’: 0.00017626081318364395, ’e_layers’: 7, ’d_ff’: 128, ’d_model’: 512, ’dropout’: 0.3870070203430017, ’seg_len’: 10, ’baseline’: 0, ’cross_factor’: 16, ’seq_len’: 192, ’affine’: 1
DLinear CI-336 lr: 0.0098, seq_len: 512
iTransformers CD-336 lr:0.0003, e_layers:8, d_ff:2048, d_model:1024, dropout:0.277, seq_len:192
TimeMixer CI-336 lr:0.0001, d_ff:256, d_model:128, e_layers:2, seq_len:720
TimeMixer CD-336 lr:0.0021, d_ff:256, d_model:128, e_layers:4, seq_len:192
PatchTST CI-720
PatchTST CD-720
TSMixer CI-720
TSMixer CD-720 lr:0.0090,num_blocks:5,hidden_size:64,dropout:0.133,activation:relu,seq_len:192
CrossFormer CI-720 OOM
CrossFormer CD-720 OOM
DLinear CI-720 lr: 0.0098, seq_len: 512
iTransformers CD-720 lr:0.0012, e_layers:4, d_ff:2048, d_model:512, dropout:0.360, seq_len:192
TimeMixer CI-720 lr:0.0001, d_ff:256, d_model:512, e_layers:2, seq_len:720
TimeMixer CD-720 lr:0.0021, d_ff:256, d_model:128, e_layers:4, seq_len:192
Table 26. Hyperparameter settings for the electricity dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
PatchTST CD-96 lr:0.0001,e_layers:1,d_ff:1.02e+03,d_model:128,dropout:0.098,fc_dropout:0.465,patch_size:8,stride:8,seq_len:336
TSMixer CI-96 lr:0.0002,num_blocks:5,hidden_size:1.02e+03,dropout:0.205,activation:relu,seq_len:512
TSMixer CD-96 lr:0.0010,num_blocks:1,hidden_size:64,dropout:0.033,activation:relu,seq_len:192
CrossFormer CI-96 lr: 1.43e-05, e_layers: 5, d_ff: 512, d_model: 256, dropout: 0.098, seg_len: 9, baseline: 0, cross_factor: 11, seq_len: 336
CrossFormer CD-96 lr: 0.0001, e_layers: 3, d_ff: 256, d_model: 512, dropout: 0.123, seg_len: 8, baseline: 0, cross_factor: 7, seq_len: 336
DLinear CI-96 lr: 0.0020, seq_len: 720
iTransformers CD-96 lr:0.0001, e_layers:8, d_ff:2048, d_model:1024, dropout:0.356, seq_len:192
TimeMixer CI-96 lr:6.00e-06, d_ff:512, d_model:256, e_layers:1, seq_len:720
TimeMixer CD-96 lr:1.00e-05, d_ff:512, d_model:512, e_layers:1, seq_len:720
PatchTST CI-192 lr:0.0001,e_layers:4,d_ff:256,d_model:128,dropout:0.469,fc_dropout:0.405,patch_size:8,stride:8,seq_len:512
PatchTST CD-192 lr:0.0002,e_layers:1,d_ff:512,d_model:1.02e+03,dropout:0.010,fc_dropout:0.333,patch_size:16,stride:8,seq_len:192
TSMixer CI-192 lr:0.0002,num_blocks:5,hidden_size:1.02e+03,dropout:0.247,activation:relu,seq_len:512
TSMixer CD-192 lr:0.0010,num_blocks:1,hidden_size:64,dropout:0.033,activation:relu,seq_len:192
CrossFormer CI-192 lr: 0.0002, e_layers: 6, d_ff: 512, d_model: 128, dropout: 0.266, seg_len: 8, baseline: 1, cross_factor: 5, seq_len: 336
CrossFormer CD-192 lr: 0.0002, e_layers: 1, d_ff: 256, d_model: 128, dropout: 0.326, seg_len: 7, baseline: 0, cross_factor: 9, seq_len: 336
DLinear CI-192 lr: 0.0077, seq_len: 720
iTransformers CD-192 lr:1.60e-05, e_layers:9, d_ff:2048, d_model:1024, dropout:0.012, seq_len:192
TimeMixer CI-192 lr 0.00008, d_ff 512, d_model 512, e_layers 4, seq_len 192
TimeMixer CD-192 lr:0.0005, d_ff:512, d_model:128, e_layers:8, seq_len:720
PatchTST CI-336 lr:0.0007,e_layers:8,d_ff:1024,d_model:512,dropout:0.150,fc_dropout:0.021,patch_size:8,stride:8,seq_len:720
PatchTST CD-336 lr:0.0007,e_layers:8,d_ff:1024,d_model:512,dropout:0.150,fc_dropout:0.021,patch_size:8,stride:8,seq_len:720
TSMixer CI-336 lr:0.0002,num_blocks:5,hidden_size:1.02e+03,dropout:0.265,activation:relu,seq_len:512
TSMixer CD-336 lr:0.0001,num_blocks:6,hidden_size:256,dropout:0.334,activation:gelu,seq_len:512
CrossFormer CI-336 lr: 0.0002, e_layers: 6, d_ff: 512, d_model: 128, dropout: 0.266, seg_len: 8, baseline: 1, cross_factor: 5, seq_len: 336
CrossFormer CD-336 OOM
DLinear CI-336 lr: 0.0014, seq_len: 720
iTransformers CD-336 lr:1.60e-05, e_layers:9, d_ff:2048, d_model:1024, dropout:0.012, seq_len:192
TimeMixer CI-336 lr:0.0001, d_ff:256, d_model:128, e_layers:2, seq_len:720
TimeMixer CD-336 lr:1.00e-05, d_ff:512, d_model:128, e_layers:1, seq_len:720
PatchTST CI-720 lr:0.0005,e_layers:7,d_ff:256,d_model:256,dropout:0.767,fc_dropout:0.878,patch_size:16,stride:8,seq_len:720
PatchTST CD-720 lr:6.07e-06,e_layers:3,d_ff:256,d_model:1.02e+03,dropout:0.006,fc_dropout:0.642,patch_size:16,stride:8,seq_len:192
TSMixer CI-720 lr:0.0003,num_blocks:2,hidden_size:1.02e+03,dropout:0.108,activation:relu,seq_len:512
TSMixer CD-720 lr:0.0002,num_blocks:8,hidden_size:32,dropout:0.281,activation:relu,seq_len:512
CrossFormer CI-720 lr: 0.0002, e_layers: 1, d_ff: 512, d_model: 128, dropout: 0.228, seg_len: 6, baseline: 0, cross_factor: 17, seq_len: 336
CrossFormer CD-720 lr: 0.0080, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.189, seg_len: 8, baseline: 0, cross_factor: 9, seq_len: 192
DLinear CI-720 lr: 0.0030, seq_len: 720
iTransformers CD-720 lr:0.0001, e_layers:8, d_ff:1024, d_model:1024, dropout:0.356, seq_len:192
TimeMixer CI-720 lr:0.0001, d_ff:256, d_model:512, e_layers:2, seq_len:720
TimeMixer CD-720 lr:1.00e-05, d_ff:512, d_model:512, e_layers:1, seq_len:720
Table 27. Hyperparameter settings for the weather dataset. The best variant for each model is shown.
Model Setting Best Hyper. Params
PatchTST CI-96
PatchTST CD-96
TSMixer CI-96 lr:0.0097,num_blocks:9,hidden_size:32,dropout:0.271,activation:relu,seq_len:512
TSMixer CD-96 lr:0.0097,num_blocks:5,hidden_size:32,dropout:0.311,activation:relu,seq_len:512
CrossFormer CI-96 lr’: 0.000515701560401399, ’e_layers’: 7, ’d_ff’: 128, ’d_model’: 256, ’dropout’: 0.25730612136617825, ’seg_len’: 11, ’baseline’: 0, ’cross_factor’: 10, ’seq_len’: 96, ’affine’: 1
CrossFormer CD-96 lr’: 5.2122436504604685e-05, ’e_layers’: 3, ’d_ff’: 256, ’d_model’: 512, ’dropout’: 0.12303593071645767, ’seg_len’: 8, ’baseline’: 0, ’cross_factor’: 7, ’seq_len’: 336, ’affine’: 0
DLinear CI-96 lr: 0.0078, seq_len: 512
iTransformers CD-96 OOM
TimeMixer CI-96 OOM
TimeMixer CD-96 From TimeMixer Paper
PatchTST CI-192
PatchTST CD-192
TSMixer CI-192 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
TSMixer CD-192 lr:0.0097,num_blocks:5,hidden_size:32,dropout:0.311,activation:relu,seq_len:512
CrossFormer CI-192 OOM
CrossFormer CD-192 OOM
DLinear CI-192 lr: 0.0078, seq_len: 512
iTransformers CD-192 OOM
TimeMixer CI-192 OOM
TimeMixer CD-192 From TimeMixer Paper
PatchTST CI-336
PatchTST CD-336
TSMixer CI-336 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
TSMixer CD-336 lr:0.0097,num_blocks:5,hidden_size:32,dropout:0.311,activation:relu,seq_len:512
CrossFormer CI-336 OOM
CrossFormer CD-336 OOM
DLinear CI-336 lr: 0.0038, seq_len: 336
iTransformers CD-336 OOM
TimeMixer CI-336 OOM
TimeMixer CD-336 From TimeMixer Paper
PatchTST CI-720
PatchTST CD-720
TSMixer CI-720
TSMixer CD-720 lr:0.0097,num_blocks:5,hidden_size:32,dropout:0.311,activation:relu,seq_len:512
CrossFormer CI-720 OOM
CrossFormer CD-720 OOM
DLinear CI-720 lr: 0.0070, seq_len: 336
iTransformers CD-720 OOM
TimeMixer CI-720 OOM
TimeMixer CD-720 From TimeMixer Paper
Table 28. Hyperparameter settings for the Traffic dataset. The best variant for each model is shown.

### v3

@@
 Channel Dependence, Limited Lookback Windows, and the Simplicity of Datasets: How Biased is Time Series Forecasting?
-Abstract.
-In Time Series Forecasting (TSF), the lookback window (the length of historical data used for prediction) is a critical hyperparameter that is often set arbitrarily, undermining the validity of model evaluations.
+Abstract
+In Long-term Time Series Forecasting (LTSF), the lookback window is a critical hyperparameter often set arbitrarily, undermining the validity of model evaluations.
 We argue that the lookback window must be tuned on a per-task basis to ensure fair comparisons.
@@
 Using Granger causality analysis and ODE datasets (with implicit channel correlations), we demonstrate that the true strength of multivariate Channel-Dependent (CD) models emerges on datasets with strong, inherent cross-channel dependencies, where they significantly outperform CI models.
-We conclude with three key recommendations for improving TSF research: (i) treat the lookback window as a critical hyperparameter to be tuned, (ii) use statistical analysis of datasets to inform the choice between CI and CD architectures, and (iii) favor CD models in applications with limited data.
-conference: ; ; copyright: none
-1. Introduction
-Figure 1.
-Test-error (MSE) on y-axis for various lookback windows on the x-axis for the weather dataset and a forecast horizon of 96.
-CI denotes the channel-independent variant and CD denotes the channel-dependent variant of TSF architectures.
-The black dotted line shows lookback window of 96, where most methods compare performance.
-Time Series Forecasting (TSF) is a fundamental challenge in machine learning that drives key innovations across domains such as finance, energy, healthcare, and climate science (Godahewa et al., 2021).
-The core task of TSF is to predict future observations by leveraging patterns within historical time-dependent data.
+We conclude with four key recommendations for improving TSF research: (i) consider the lookback window as a key hyperparameter to tune, (ii) for standard datasets, examining CI architectures is advantageous, (iii) leverage statistical analysis of datasets to guide the choice between CI and CD architectures, and (iv) prefer CD models in scenarios with limited data.
+1 Introduction
+Figure 1: Test-error (MSE) on the y-axis for various lookback windows on the x-axis for the weather dataset and a forecast horizon of 96.
+CI denotes the channel-independent variant and CD denotes the channel-dependent variant.
+The black dotted line shows a lookback window of 96.
+Long-term Time Series Forecasting (LTSF) is a core machine learning task that aims to predict future observations by using patterns in historical time-dependent data in various domains such as finance, energy, healthcare, and climate science [6].
 Nevertheless, the literature suffers from two significant experimental shortcomings that compromise the fairness and reliability of model comparisons.
-First, in the multivariate TSF learning setting, the historical behavior is modeled using a fixed-length sequence of past observations—known as the lookback window—to forecast a set of future values over a given prediction horizon.
-Despite its fundamental role, the choice of lookback window remains surprisingly underexplored in TSF literature.
-Most existing Long-term Time Series Forecasting (LTSF) studies adopt a fixed lookback window (mostly 96) for all models and datasets, often without justification or transparency (Liu et al., 2024; Zeng et al., 2023; Zhang and Yan, 2023; Yi et al., 2023; Chen et al., 2023, 2025; Lin et al., 2024, 2025; Huang et al., 2025).
-While this practice simplifies experimental comparisons, it masks a critical reality: the optimal lookback window can vary significantly across models and datasets, and failing to tune it can lead to misleading conclusions.
-As demonstrated in Figure˜1, for the weather dataset, the optimal lookback window length differs across various state-of-the-art TSF models.
-Consequently, performance comparisons at a single, fixed lookback window may misrepresent which models are truly state-of-the-art under realistic, well-optimized conditions.
-Comparable studies have been conducted by Zeng et al.
-(2023) on the Electricity dataset and by Tong and Yuan (2025) on the Weather dataset.
-However, unlike previous studies, we explicitly tune all the hyperparameters for each lookback window to generate the figure ensuring reliable comparisons.
-The optimal lookback window is highly dependent on both the intrinsic properties of the dataset and the specific architecture of the forecasting model.
-For example, when comparing within the same weather dataset (Figure˜1), Channel Independent (CI) or univariate models tend to favor longer lookback windows (e.g., 336 or 512), enabling deeper temporal modeling per variable (Nie et al., 2023; Zeng et al., 2023; Chen et al., 2023).
-Whereas the Channel Dependent (CD) or multivariate models jointly model all channels and often operate with shorter lookback windows (e.g., 96), leveraging channel dependence to compensate for the reduced temporal depth (Wang et al., 2024b; Liu et al., 2024).
-Even within CD class of methods (i.e, across different multivariate models) certain architectures benefit more from long histories than others.
-For example, Crossformer-CD significantly outperforms TimeMixer-CD at a lookback window of 96, yet its performance drops below TimeMixer-CD when the lookback window is extended to 336.
-These findings motivate our central claim:
-Tuning the lookback window as a model-specific hyperparameter is essential for fair evaluation and reliable assessment in time series forecasting.
-Beyond the issue of lookback tuning, a second critical limitation undermines the fairness and practical relevance of current TSF benchmarks.
-Most widely-used multivariate forecasting datasets such as ETTh, ETTm, Electricity, Weather, and Traffic exhibit minimal channel dependence among their variables.
-This lack of inter-variable relational complexity renders these benchmarks overly simplistic; even simple linear models (Zeng et al., 2023) can achieve performance comparable to sophisticated nonlinear architectures.
-This raises a key question:
-Does channel dependence truly matter in multivariate LTSF?
-To answer this, we conduct an in-depth empirical analysis, combining traditional performance metrics with statistical tools such as Granger causality (Granger, 1969).
-Our analysis reveals that the benefits of multivariate modeling depend strongly on two factors: (1) the amount of historical data available, and (2) the degree of causal influence among channels.
-When inter-channel interactions are weak or datasets are large enough for univariate models to exploit long-term patterns, the added complexity of modeling channel dependence may offer limited returns.
-Additionally, through experiments on complex channel-dependent chaotic Ordinary Differential Equation (ODE) datasets (Gilpin, 2021), we reveal that the standard benchmarks lack realistic cross-channel interactions essential for multivariate forecasting.
-This simplicity disproportionately favors CI models and reduce the ability to fairly assess models designed to exploit channel dependencies as also shown in (Nie et al., 2023; Chen et al., 2023).
-Based on these experimental and statistical findings, we offer the following practical recommendations for evaluating and designing multivariate time series forecasting models:
-  • Channel dependence is unnecessary for top performance on standard benchmarks; tuning the lookback window enables CI models like PatchTST to match or surpass CD ones (Section˜6.1).
-  • On complex datasets with strong channel correlations (supported by granger causality Section˜6.2), CD models outperform CI models (Section˜6.3).
-  • Longer lookback windows generally favor univariate models, highlighting the need for careful lookback tuning to ensure fair comparisons (Section˜6.4).
-2. Related Works
-LTSF models typically fall into two categories: Channel-Dependent (CD), which model cross-channel interactions, and Channel-Independent (CI), which treat each channel in isolation and focus purely on temporal dynamics.
-CI models such as DLinear (Zeng et al., 2023) and PatchTST (Nie et al., 2023) process each feature independently and consistently achieve strong performance on widely-used benchmarks, casting doubt on the necessity of cross-channel modeling.
-In contrast, MLP-based CD methods like TSMixer (Chen et al., 2023), TimeMixer (Wang et al., 2024a), and transformer-based architectures including iTransformer (Liu et al., 2024) and Crossformer (Zhang and Yan, 2023) aim to capture inter-channel dependencies explicitly.
-However, the predominantly univariate nature of standard benchmark datasets limits the ability of CD methods to demonstrate their advantages.
-To mitigate this issue, we incorporate evaluations on ODE-based datasets (Gilpin, 2021) designed to embed implicit channel correlations and enabling more representative comparisons on channel-dependent datasets.
-The effectiveness of CI methods also appears to be amplified by longer lookback windows.
-For instance, DLinear adopts a lookback of 336 steps, whereas transformer-based baselines were defaulted to a lookback of 96.
-PatchTST, a successor to DLinear, employs a fixed lookback window of 512 or 336 and addresses the limitations of its predecessor by systematically tuning baseline methods over a broader range of lookback lengths.
-Nonetheless, the widespread adoption of a fixed 96-step lookback, popularized by iTransformer has constrained CI models and impeded equitable benchmarking.
-Subsequent studies (Tong and Yuan, 2025; Chen et al., 2025; Lin et al., 2024, 2025; Huang et al., 2025) frequently report improved performance of proposed methods by retaining short fixed windows for strong CI baselines like PatchTST, thereby hindering fair comparisons due to inconsistent tuning.
-Several recent studies (Wang et al., 2024a; Nie et al., 2023; Zeng et al., 2023; Tong and Yuan, 2025) acknowledge the importance of lookback window selection in forecasting performance.
-However, lookback tuning is frequently confined to ablation studies, with many works failing to apply consistent tuning across both baselines and proposed methods in their main experiments.
-For instance, TimeMixer reports results using a unified set of hyperparameters (including a fixed lookback window) in the main results, while providing tuned results only in the appendix.
-Another work proposed IRPA method (Tong and Yuan, 2025) that explores various lookback lengths (480, 1920, 3600) for their method, yet retains a fixed lookback of 96 for all baselines, diminishing their comparative performance and inflating IRPA’s gains.
-In contrast, our evaluations perform systematic lookback window tuning for all models, uncovering genuinely optimal configurations.
-3. Problem Statement
-Multivariate time series is a sequence $({\bm{x}}^{l})_{l=1}^{T}\in(\mathbb{R}^{C})^{T}$, where each ${\bm{x}}^{l}\in\mathbb{R}^{C}$ represents observations from $C$ channels at time step $l$.
-This sequence can be represented as a matrix ${\bm{X}}^{\text{full}}\in\mathbb{R}^{C\times T}$.
-If $C=1$, this represents the univariate forecasting setting.
+First, LTSF models utilize a fixed-length sequence of past observations, the lookback window ($L$), for prediction.
+Despite its fundamental role, The lookback window is often set arbitrarily (typically $L=96$) across benchmarks without justification [15, 25, 26, 24, 3, 2, 14, 13, 9].
+While simplifying comparisons, using a fixed lookback window masks a critical reality: the optimal lookback window varies across models and datasets (Figure˜1).
+Fixed-window comparisons may thus obscure which models truly achieve state-of-the-art performance under optimized settings.
+Unlike [25, 21], our study explicitly tunes all hyperparameters for every lookback window to ensure reliable evaluation.
+The optimal lookback window depends on both dataset properties and model architecture.
+As shown in Figure˜1, Channel-Independent (CI) models typically prefer longer windows (e.g., 336–512) to capture temporal patterns [16, 25, 3], whereas Channel-Dependent (CD) models often thrive with shorter windows (e.g., 96) by leveraging cross-channel dependencies [23, 15].
+Even within the CD class, benefits vary: Crossformer-CD outperforms TimeMixer-CD at $L=96$, but falls behind at $L=336$.
+These findings motivate our first claim: Tuning the lookback window as a model-specific hyperparameter is essential for fair evaluation and reliable assessment in time series forecasting.
+The second limitation concerns the practical relevance of current LTSF benchmarks.
+Standard multivariate datasets—such as ETT, Electricity, Weather, and Traffic—exhibit very limited dependence among channels.
+This lack of inter-variable complexity renders these benchmarks overly simplistic, raising a critical question: Does channel dependence truly matter in multivariate LTSF?
+Table 1: Frequency of CI and CD wins across models and datasets. ’ODE-6’ denotes the 6 ODE datasets with intrinsic channel correlation; ’Std.-7’ represents the 7 standard datasets.
+Model ODE-6 Std.-7
+CI CD CI CD
+PatchTST 4 2 6 1
+TSMixer 0 6 7 0
+Crossformer 0 6 4 3
+TimeMixer 2 4 4 3
+Totals 6 18 21 7
+To address this, we conduct an empirical analysis combining traditional performance metrics with Granger causality [7].
+Our results reveal that multivariate modeling benefits depend on: (1) available historical data and (2) the degree of causal influence among channels.
+When inter-channel interactions are weak, modeling channel dependence offers limited returns.
+Consequently, we utilize complex, chaotic Ordinary Differential Equation (ODE) datasets [5], which demonstrate that standard benchmarks fail to capture the cross-channel interactions crucial for multivariate forecasting.
+As shown in Table˜1, CD variants consistently outperform CI counterparts on highly coupled ODE datasets.
+Conversely, standard datasets’ simplicity disproportionately favors CI models, limiting the fair evaluation of methods designed for channel dependencies [16, 3].
+Our main contributions are as follows:
+  • Lookback Evaluation: We show that tuning the lookback window is a critical, model-specific requirement; neglecting it introduces significant baseline bias.
+  • Benchmark Simplification: We reveal that standard LTSF benchmarks lack sufficient channel dependence, unfairly handicapping CD models.
+  • Causal Analysis: Using Granger causality and ODE datasets, we prove CD models excel only when complex, verifiable inter-channel interactions are present.
+  • Guidelines: We provide actionable recommendations for choosing between CI and CD architectures based on lookback and causal complexity.
+2 Related Work
+LTSF models typically fall into two categories: Channel-Dependent (CD) and Channel-Independent (CI).
+CI models like DLinear [25] and PatchTST [16] emphasize temporal dynamics, often outperforming CD models such as TSMixer [3], iTransformer [15], and Crossformer [26].
+Recent advancements in CD modeling, including TimeFilter’s [8] graph filtration and DUET’s [18] hierarchical clustering, attempt to capture complex spatial dependencies, yet the predominantly univariate nature of common benchmarks often masks their potential.
+Recently, a surge in using Ordinary Differential Equations (ODEs) has generated rigorous, high-causality benchmarks, exemplified by Physiome [12] and chaotic ODE collections [5].
+These provide mathematically-defined channel correlations, offering a more robust multivariate testbed than standard datasets.
+A critical factor in model evaluation is the lookback window length ($L$).
+While CI methods typically benefit from $L\in\{336,512\}$, many studies [21, 2, 13, 9] fix $L=96$ for baselines.
+This practice introduces evaluation bias; for instance, IRPA [21] explores variable lookbacks for its own method but maintains $L=96$ for baselines, potentially overstating gains.
+Although some works acknowledge lookback sensitivity [22, 16], systematic tuning is often relegated to secondary ablations.
+While comprehensive benchmarks like TFB [17] advocate for fair evaluation across diverse datasets, they do not address the compounding issues of lookback bias and dataset coupling.
+In contrast, our study systematically tunes $L$ for all models and utilizes ODE-based evaluations to ensure architectural comparisons are based on truly optimal configurations.
+3 Problem Statement
 To train forecasting models effectively, especially when $T$ is large, we partition the long sequence into multiple shorter subsequences using a sliding window approach.
-Each subsequence is split into two parts: an input segment (lookback window) of length $L$ and a target segment (forecast horizon) of length $H$.
-Specifically, a time-series forecasting dataset is composed of input-output pairs $({\bm{X}},{\bm{Y}})\in\mathbb{R}^{C\times L}\times\mathbb{R}^{C\times H}$, where:
-${\bm{X}}=({\bm{x}}^{t},{\bm{x}}^{t+1},\dots,{\bm{x}}^{t+L-1})$ is referred to as the forecasting query, and
-${\bm{Y}}=({\bm{x}}^{t+L},{\bm{x}}^{t+L+1},\dots,{\bm{x}}^{t+L+H-1})$ is the corresponding forecasting answer.
-The goal of the time-series forecasting problem is to train a model $\hat{{\bm{Y}}}:\mathbb{R}^{C\times L}\rightarrow\mathbb{R}^{C\times H}$ on a dataset $\mathcal{D}_{\operatorname{train}}\subseteq\mathbb{R}^{C\times L}\times\mathbb{R}^{C\times H}$, constructed from many such overlapping or non-overlapping windows, sampled from the full sequence.
-This dataset is assumed to be drawn i.i.d.
-from some underlying distribution $p$.
-The objective is to minimize the expected loss:
-(1) $\mathbb{E}_{({\bm{X}},{\bm{Y}})\sim p}\,\ell({\bm{Y}},\hat{{\bm{Y}}}({\bm{X}}))$
-where $\ell:\mathbb{R}^{C\times H}\times\mathbb{R}^{C\times H}\to\mathbb{R}$ is a suitable loss function, commonly the mean squared error (MSE).
-4. Background
-Table 1.
-Average MSE over the 4 forecasting horizons for each variant of the baseline evaluated for the standard datasets.
-The best variant for each model is in bold and the best overall model is in blue.
-Note here due to running out of memory (OOM) for some experiments related to electricity and traffic, we reported the results with "*" from TimeMixer paper from the experiment where they tune the lookback window as well as the hyperparameters (please refer to appendix E in (Wang et al., 2024a)).
+Each subsequence is split into an input segment of length $L$ (representing the lookback window) and a target segment of length $H$ (the forecast horizon).
+A forecasting dataset thus consists of pairs $(\mathbf{X},\mathbf{Y})\in\mathbb{R}^{C\times L}\times\mathbb{R}^{C\times H}$, where $\mathbf{X}=(\mathbf{x}^{t},\dots,\mathbf{x}^{t+L-1})$ is the forecasting query and $\mathbf{Y}=(\mathbf{x}^{t+L},\dots,\mathbf{x}^{t+L+H-1})$ is the corresponding forecasting answer.
+The goal of the time-series forecasting task is to learn a model $\hat{\mathbf{Y}}:\mathbb{R}^{C\times L}\rightarrow\mathbb{R}^{C\times H}$ using a dataset $\mathcal{D}_{\text{train}}\subseteq\mathbb{R}^{C\times L}\times\mathbb{R}^{C\times H}$, constructed from many such windows sampled from the full sequence.
+The dataset is assumed to be drawn i.i.d.
+from an underlying distribution $p$.
+The objective is to minimize the expected loss, $\mathbb{E}_{(\mathbf{X},\mathbf{Y})\sim p}\,\ell(\mathbf{Y},\hat{\mathbf{Y}}(\mathbf{X}))$, where $\ell$ is typically the mean squared error (MSE).
+While the underlying time series is intrinsically dependent, this formulation treats the windowed pairs $(\mathbf{X},\mathbf{Y})$ as realizations of a stationary process for the purpose of supervised optimization.
+This allows the model $\hat{\mathbf{Y}}$ to explicitly focus on capturing the temporal and inter-channel dependencies within each local windowed segment.
+4 Background
+Fixing Lookback Windows:
+The problem formulation in Section˜3 does not strictly require the forecasting query $\mathbf{X}$ to use a fixed lookback length $L$.
+However, most recent LTSF models adopt a fixed window (commonly $L=96$) during training, primarily for implementation convenience or adherence to benchmarking standards.
+As argued previously, the lookback window should be treated as a model-specific hyperparameter, integral to the model design.
+Channel Independence:
+Recent studies [16, 3] demonstrate that on widely used benchmarks, CI models often suffice for accurate forecasting.
+This suggests that cross-channel interactions may be unnecessary for certain tasks.
+Formally, a model $\hat{\mathbf{Y}}:\mathbb{R}^{C\times L}\to\mathbb{R}^{C\times H}$ is channel-independent if there exists a univariate predictor $f:\mathbb{R}^{L}\to\mathbb{R}^{H}$ such that for each channel $c\in\{1,\dots,C\}$:
+$\hat{\mathbf{Y}}(\mathbf{X})_{c}=f(\mathbf{x}_{c,:})\quad\forall c\in\{1,\dots,C\}.$ (1)
+Granger Causality.
+Granger causality [7] evaluates whether a channel $c_{2}$ provides predictive information for $c_{1}$ beyond $c_{1}$’s own history.
+We compare the univariate sum of squared residuals ($SSR_{u}$) from a restricted model (CI assumption) and the multivariate ($SSR_{mv}$) from an unrestricted model (CD assumption) incorporating $c_{2}$.
+Using $x_{c,t}$ to denote the observation of channel $c$ at time $t$, the residuals are:
+$SSR_{u}$ $=\sum_{t=L+1}^{L+H}\bigg(x_{c_{1},t}-a_{0}-\sum_{i=1}^{L}a_{i}x_{c_{1},t-i}\bigg)^{2},$ (2)
+$SSR_{mv}$ $=\sum_{t=L+1}^{L+H}\bigg(x_{c_{1},t}-a_{0}-\sum_{i=1}^{L}a_{i}x_{c_{1},t-i}-\sum_{i=1}^{L}b_{i}x_{c_{2},t-i}\bigg)^{2}.$ (3)
+where $a_{0}$ is the intercept, and $a_{i},b_{i}$ are learned coefficients for the lagged values of $c_{1}$ and $c_{2}$ respectively.
+The $F$-statistic measures the significance of the error reduction:
+$F=\frac{(SSR_{u}-SSR_{mv})/L}{SSR_{mv}/(N-k_{mv})},$ (4)
+where $L$ is the number of added parameters (lags of $c_{2}$), $N$ the number of observations, and $k_{mv}$ the total parameters in the multivariate model.
+A high $F$-statistic indicates that $c_{2}$ Granger-causes $c_{1}$.
+5 General Experimental Setup
+5.1 Models
+We evaluate several baselines: MLP-based (TSMixer [3], TimeMixer [23]), attention-based (PatchTST [16], iTransformer [15], Crossformer [26]), and DLinear [25].
+We use DLinear-CI and iTransformer-CD to prevent architectural overlap with TSMixer; others are tested in both variants.
+Details are in our repository.
+[^1] https://github.com/IbramMedhat/How-Biased-is-Time-Series-Forecasting- and the appendix.
+5.2 Datasets
+As noted in the introduction, we evaluate two main dataset categories to examine how increasing complexity particularly channel correlation impacts model performance.
+The Standard Datasets:
+The first category includes widely used benchmarks from recent state-of-the-art studies and surveys [16, 15, 23, 3, 26, 25]: Weather, Electricity, Traffic, and the Electricity Transformer Temperature (ETT) datasets.
+These benchmarks gained popularity following their adoption in the Informer paper [27].
+The Chaotic ODE Benchmark:
+As a second dataset collection, we consider the benchmark proposed by [5].
+This benchmark was introduced as a time-series collection that features complex time series dynamics with clear real world data properties, providing a challenging and realistic setting for CD time-series forecasting.
+A demonstration for the DoublePendulum dataset and how it simulates a recurrent patterns across different channels can be seen in Figure˜2.
+This figure particularly emphasizes periods of synchronization in the angular velocities of the two pendulums during the final 200 timesteps, indicative of inter-channel correlation.
+Figure 2: Time series visualization of the Double Pendulum ODE system.
+Notably, the bottom two channels (angular velocities of the two pendulums) exhibit synchronized behavior in the last 200 time steps, demonstrating time-varying inter-channel correlation.
+We select a subset of datasets from this benchmark based on two distinctive criteria: (i) The datasets follow complex trajectories that challenge most time series models and (ii) they are derived directly from the underlying ODEs, ensuring a clear mathematical correlation between variables.
+For criterion (i), we use the Largest Lyapunov exponent [5] to quantify dataset complexity, selecting datasets with high values to ensure forecasting remains non-trivial task that cannot be solved easily based on one channel information.
+To illustrate criterion (ii), we show below a simplified form of the Double Pendulum system [4], assuming equal masses and rod lengths.
+The coupling between angular velocities $\frac{d\theta_{1}(t)}{dt},\frac{d\theta_{2}(t)}{dt}$ and angles $\theta_{1}(t),\theta_{2}(t)$ at time $t$ from Equations˜5 and 6 is evident, with parameters $g$ (gravity) and $l$ (rod length) incorporated.
+Such systems are essential in modeling real-world physics processes, including robotics [10].
+$\frac{d^{2}\theta_{1}(t)}{dt^{2}}=-\frac{2g}{l}\theta_{1}(t)+\frac{g}{l}\theta_{2}(t)$ (5)
+$\frac{d^{2}\theta_{2}(t)}{dt^{2}}=\frac{2g}{l}\theta_{1}(t)-\frac{2g}{l}\theta_{2}(t)$ (6)
+5.3 Hyperparameter Tuning
+We ensure fair evaluation via Bayesian optimization (Optuna [1]) over 20 trials per model–dataset–horizon.
+The lookback window $L$ is tuned across $\{96,\allowbreak 192,\allowbreak 336,\allowbreak 512,\allowbreak 720\}$, with the best variant selected via validation MSE.
+We prioritize MSE as the primary metric to remain consistent with LTSF benchmarks and to penalize large outliers common in complex series.
+Crucially, we address data-loading biases noted in [17] by setting drop_last=False during testing.
+This ensures no trailing points are discarded, providing a complete, unbiased evaluation across the entire distribution and confirming that lookback-CD interactions are not artifacts of data truncation.
+5.4 Granger Causality
+We perform Granger causality analysis following [7] with two preprocessing steps: (i) assessing stationarity and applying iterative differencing where necessary, and (ii) removing one channel from any pair with a Pearson correlation above 0.95 to reduce redundancy.
+Following the $F$-test (Section˜4), we report the average $F$-score and the percentage of pairs rejecting the null hypothesis.
+6 Results
+6.1 Evaluation of Standard Datasets
+We reproduce existing time-series forecasting models under our evaluation setup, performing extensive hyperparameter tuning (including lookback window) and averaging results over five random seeds.
+Reported scores in Table˜2 are averaged across horizons 96, 192, 336, and 720.
+PatchTST remains the state-of-the-art model in our fair evaluation.
+Its CI variant achieves the best performance on three of seven datasets and ranks second or third on another three.
+Only on the Weather dataset does it fall outside the top three, likely due to stronger inter-channel dependencies indicated by Granger causality.
+Overall, PatchTST attains an average rank of 2.14, a notable margin (0.57) ahead of the second-best model, CI TSMixer.
+Linear models continue to perform competitively.
+DLinear, for instance, achieves an average rank of 5.43 out of 10 across the six standard datasets and ranks second on ETTh1, slightly behind PatchTST.
+Table 2: Mean MSE across four horizons. Bold: best per model; blue: overall best. (*): sourced from [22]. OOM: Out-of-memory.
 Dataset PatchTST TSMixer Crossfor. DLin. iTrans. TimeMixer
@@
 Avg. Rank 2.14 5.43 2.71 5.00 7.29 8.14 5.43 5.86 5.29 6.57
-Table 2.
-Average F-scores and percentage of rejected null hypotheses $H_{0}$ ($p=0.05$) across different lag values, comparing the standard datasets with the ODE benchmark.
-Bold values indicate the highest correlation at each lag.
-Dataset Type (lag) Average F-score $H_{0}$ rejected (%, p=0.05)
+6.2 Granger Causality for Evaluation of Channel Correlations
+Building on previous findings that multivariate models do not consistently outperform univariate models on standard datasets, we perform a statistical analysis of channel correlations across both standard datasets and ODE datasets to determine where inter-channel dependencies exist and multivariate models might be necessary.
+We compare both benchmarks using Granger causality tests with lags (lookback windows) of 30, 96, and 192.
+F-scores and percentages of channel pairs showing causality (i.e.
+$H_{0}$ rejected) are reported in Table˜3, with higher values indicating stronger correlations.
+The ODE datasets show significantly stronger inter-channel dependencies with higher F-scores and $H_{0}$ rejection at all lags, while standard datasets have much weaker correlations, with some F-scores dropping below 1.0 at higher lags.
+This supports and explains earlier observations where univariate models often match or outperform multivariate ones on standard datasets due to weak coupling.
+Increasing lag generally reduces correlations in both dataset types, especially for standard datasets, suggesting longer lookback windows favor univariate models that do not use cross-channel signals.
+In contrast, ODE datasets maintain strong correlations even at longer lags (F-scores of 88.5 at lag 96 and 35.29 at lag 192), supporting their use as testbeds for models handling strong cross-channel dynamics.
+Table 3: Average F-scores and $H_{0}$ rejection rates ($p=0.05$). Bold indicates the highest correlation at each lag.
+Dataset Type (lag) Average F-score $H_{0}$ rejected (%)
 ODEs (30) 335.21 73%
@@
 Standard (192) 1.23 27%
-Table 3.
-Average MSE across four forecasting horizons for each model variant on the ODE Benchmark.
-Bold values highlight the best variant within each model; the best overall result per row is shown in blue.
+6.3 Evaluation on the ODE Benchmark
+As discussed in Section˜6.1, DLinear remains a competitive baseline on standard datasets, with CI models outperforming CD counterparts.
+However, on the ODE datasets, model performance differs significantly.
+CD variants consistently outperform CI models in 18 out of 24 experiments as shown in Table˜4, underscoring the importance of modeling inter-channel dependencies in datasets governed by channel-interactions.
+Interestingly, Linear models perform poorly on ODE datasets.
+DLinear ranks worst in 5 of 6 ODE datasets, with an average rank of 9.83.
+Compared to Crossformer, DLinear’s performance drops between 45% and 118%, with a median decline of 58.3%.
+These results challenge the notion that linear models compete well with deep architectures, highlighting the need to evaluate models on datasets with varying structural characteristics.
+For data with intrinsic inter-channel correlations, like ODE systems, model architectures must explicitly account for these dependencies to achieve strong performance.
+We note that while Granger causality tests provide a rigorous measure of linear statistical predictability between channels, they serve as a proxy for inter-channel dependency in this context rather than implying structural or physical causality in the strong sense.
+Table 4: Mean MSE on the ODE Benchmark. Bold: best per model; blue: overall best. Superscript references statistical details in the extended version.
 Dataset PatchTST TSMixer Crossfor. DLin. iTrans. TimeMixer
@@
 Avg. Rank 5.17 5.67 8.83 7.50 2.50 1.50 9.83 5.17 5.50 3.30
-Fixing Lookback Windows:
-The problem formulation described earlier places no intrinsic requirement on the forecasting query $\bm{X}$} to have a fixed lookback length $L$.
-Nevertheless, most recent LTSF models fix the lookback window during training, typically to a value of $L=96$.
-This constraint is usually imposed for implementation convenience, computational efficiency, or alignment with benchmarking conventions, rather than for a principled reason.
-As we have argued, the lookback window should instead be treated as an integral property of the model, similar to any other model-specific hyperparameter.
-Channel Independence:
-One of the main findings of PatchTST (Nie et al., 2023) and later by TSMixer (Chen et al., 2023) is that, on widely used benchmark datasets, CI models are sufficient for accurate forecasting.
-This suggests that interactions between different channels are not essential, and that each channel can be modeled in isolation.
-Formally, a model $\hat{{\bm{Y}}}:\mathbb{R}^{C\times L}\to\mathbb{R}^{C\times H}$ is said to be channel-independent if for all ${\bm{X}}\in\mathbb{R}^{C\times L}$ and any channel input ${X}_{c}\in\mathbb{R}^{L}$,
-(2) $\hat{{\bm{Y}}}({\bm{X}})_{c}=\hat{{\bm{Y}}}({{X}}_{c})$
-i.e., the prediction for any given channel $c$ is generated using only its own historical data, independent of all other channels.
-Granger Causality:
-The Granger causality test (Granger, 1969) evaluates whether past values of $X_{c_{2}}$ improve forecasts of $Y_{c_{1}}$ beyond what is achievable using only past values of $X_{c_{1}}$.
-If including $X_{c_{2}}$ reduces prediction error in an autoregressive model, then $X_{c_{2}}$ provides a useful linear signal for forecasting channel $c_{1}$.
-To quantify this directional relationship, we compute the F-statistic (Equation˜5) for each channel pair, measuring the predictive gain from adding $X_{c_{2}}$.
-The F-statistic compares the sum of squared residuals (SSR) of two models: a univariate model using only lagged values of $c_{1}$ ($SSR_{u}$, Equation˜3), and a multivariate model that additionally includes lagged values of $c_{2}$ ($SSR_{mv}$, Equation˜4).
-The univariate model uses coefficients $\hat{a}$ applied to $X_{c_{1}}^{t-i}$ over a lag window of length $L$, where $i=1,\dots,L$; the multivariate model extends this with coefficients $\hat{b}$ for $X_{c_{2}}^{t-i}$.
-Residuals are computed for $t=L+1,\dots,L+H$.
-In Equation˜5, $k_{u}$ is the number of added parameters associated with channel $c_{2}$, and $k_{mv}$ is the total number of parameters in the multivariate model.
-The denominator degrees of freedom are $N-k_{mv}$, where $N$ is the number of observations.
-A large F-statistic indicates a significant improvement in explained variance when including $c_{2}$.
-We reject the null hypothesis $H_{0}\!:$$c_{2}$ does not Granger-cause $c_{1}$” when this improvement is statistically significant.
-(3) $SSR_{u}=\sum_{t=L+1}^{L+H}\bigg(Y_{c_{1}}^{t}-\hat{a}_{0}-\sum_{i=1}^{L}\hat{a}_{i}X_{c_{1}}^{t-i}\bigg)^{2},$
-(4) $SSR_{mv}=\sum_{t=L+1}^{L+H}\bigg(Y_{c_{1}}^{t}-\hat{a}_{0}-\sum_{i=1}^{L}\hat{a}_{i}X_{c_{1}}^{t-i}-\sum_{i=1}^{L}\hat{b}_{i}X_{c_{2}}^{t-i}\bigg)^{2},$
-(5) $F=\frac{(SSR_{u}-SSR_{mv})/k_{u}}{SSR_{mv}/(N-k_{mv})},$
-5. General Experimental Setup
-5.1. Models
-We evaluate a diverse set of strong baseline models, varying in architectural complexity and treatment of channel dependencies.
-These include MLP-based models (TSMixer (Chen et al., 2023), TimeMixer (Wang et al., 2024b)), attention-based models (PatchTST (Nie et al., 2023), iTransformer (Liu et al., 2024), Crossformer (Zhang and Yan, 2023)), and the simple yet competitive DLinear (Zeng et al., 2023).
-We use only the CI version of DLinear, consistent with its design as a lightweight baseline.
-For iTransformer, we use only the CD version, as removing inter-channel attention would reduce it to a basic MLP, diverging from its core design.
-Rest of the methods are available in or allow derivation of both CD and CI forms.
-Additional details on the derivation of both variants are included in the appendix.
-For additional experimental details, please refer to our code repository [^1] https://github.com/anonsub936/How-Biased-is-Time-Series-Forecasting-.
-5.2. Datasets
-As noted in the introduction, we evaluate two main dataset categories to examine how increasing complexity particularly channel correlation impacts model performance.
-The Standard Datasets:
-The first category includes widely used benchmarks from recent state-of-the-art studies and surveys (Nie et al., 2023; Liu et al., 2024; Wang et al., 2024b; Chen et al., 2023; Zhang and Yan, 2023; Zeng et al., 2023): Weather, Electricity, Traffic, and the Electricity Transformer Temperature (ETT) datasets.
-These benchmarks gained popularity following their adoption in the Informer paper (Zhou et al., 2021).
-We refer to this group as the standard datasets.
-Details on data statistics are provided in the appendix.
-The Chaotic ODE Benchmark:
-As a second dataset collection, we consider the benchmark proposed by Gilpin (2021).
-This benchmark was introduced as a time-series collection that features complex time series dynamics with clear real world data properties, providing a challenging and realistic setting for channel-dependent time-series forecasting.
-A demonstration for one of these datasets and how it simulates a recurrent patterns across different channels can be seen in Figure˜2.
-This figure particularly emphasizes periods of synchronization in the angular velocities of the two pendulums during the final 200 timesteps, indicative of inter-channel correlation.
-Additional plots of the selected ODE system are provided in the appendix.
-Figure 2.
-Time series visualization of the Double Pendulum ODE system.
-Notably, the bottom two channels (angular velocities of the two pendulums) exhibit synchronized behavior in the last 200 time steps, demonstrating time-varying inter-channel correlation.
-We select a subset of datasets from this benchmark based on two distinctive criteria: (i) The datasets follow complex trajectories that challenge most time series models and (ii) they are derived directly from the underlying ODEs, ensuring a clear mathematical correlation between variables.
-For criterion (i), we use the Largest Lyapunov exponent (Gilpin, 2021) to quantify dataset complexity, selecting datasets with high values to ensure forecasting remains non-trivial task that cannot be solved easily based on one channel information.
-To illustrate criterion (ii), we show below a simplified form of the Double Pendulum system (Elbori and Abdalsmd, 2017), assuming equal masses and rod lengths.
-The coupling between angular velocities $\frac{d\theta_{1}(t)}{dt},\frac{d\theta_{2}(t)}{dt}$ and angles $\theta_{1}(t),\theta_{2}(t)$ at time $t$ from Equations˜6 and 7 is evident, with parameters $g$ (gravity) and $l$ (rod length) incorporated.
-Such systems are essential in modeling real-world physics applications, including robotics (Jadlovská and Sarnovskỳ, 2012).
-Due to space constraints, we defer detailed discussion of another example, Cell Cycle, to the appendix.
-This dataset reflects strong variable coupling and captures real-world measurements of Cyclin-Dependent Kinase (CDK) activity, crucial to cell cycle progression (Romond et al., 1999).
-Our selection yields six datasets with rich temporal dynamics and strong inter-channel dependencies, referred to as the ODE datasets.
-Further reasoning for particularly choosing this benchmark follows through the statistical analysis in Section˜6.2.
-(6) $\frac{d^{2}\theta_{1}(t)}{dt^{2}}$ $=-\frac{2g}{l}\theta_{1}(t)+\frac{g}{l}\theta_{2}(t)$
-(7) $\frac{d^{2}\theta_{2}(t)}{dt^{2}}$ $=\frac{2g}{l}\theta_{1}(t)-\frac{2g}{l}\theta_{2}(t)$
-5.3. Hyperparameter Tuning
-To ensure fair evaluation, we apply a consistent tuning procedure across all models.
-Each hyperparameter is defined either by a discrete set (categorical) or a continuous range (real-valued).
-Performance for each hyperparameter configuration is validated on the designated validation set through the Mean Squared Error (MSE).
-The best model is chosen based on the lowest validation error, then it is tested on a the separate test split to get the final evaluation.
-We use Bayesian optimization via Optuna (Akiba et al., 2019) to efficiently explore promising configurations, running 20 trials per setup (i.e., a specific model, dataset, and forecasting horizon).
-Crucially, the lookback window is treated as a tunable hyperparameter, selected from $\{96,192,336,512,720\}$.
-Also more hyperparameters that are specific for each model are tuned thoroughly, allowing each model to operate under optimal conditions for fair comparison.
-Detailed search spaces and best hyperparameters chosen for each model are provided in the appendix.
-5.4. Granger Causality
-We conduct Granger causality analysis following the standard procedure outlined by Granger (1969), with several preprocessing steps to ensure robust and meaningful results:
-  1. (i)
-    We first assess the stationarity of each dataset. If non-stationarity is detected, differencing is applied iteratively until stationarity is achieved.
-  2. (ii)
-    To remove redundancy, we compute the Pearson correlation coefficient for all channel pairs.
-Channels with correlation exceeding 0.95 are considered highly similar, and one from each such pair is excluded from further analysis.
-After the preprocessing steps, we conduct the F-test described in Section˜4.
-Consequently, we report two important metrics: the average F-score over all valid channel pairs, and the percentage of pairs for which the null hypothesis is rejected (indicating significant Granger causality).
-These results are summarized in Section˜6.2.
-Additional implementation details, including parameter settings and the formal hypothesis definition, are provided in the appendix.
-6. Results
-6.1. Evaluation of Standard Datasets
-We reproduce the current time-series forecasting models in our above-mentioned evaluation settings, where we conduct fair and extensive hyperparameter tuning, including tuning the lookback window.
-Unless stated otherwise, all results are averaged over 5 different random seeds.
-Table˜1 presents the results averaged across forecasting horizons 96, 192, 336, and 720 for the standard datasets.
-Detailed values for each individual horizon and standard-deviations are provided in the appendix.
-PatchTST remains the state-of-the-art for time-series forecasting.
-In 3 out of 7 datasets, the standard channel-independent version of PatchTST achieves the best performance.
-For three other datasets, this version ranks second or third.
-Its performance is not among the top three only on the Weather dataset, which may be attributed to the relatively stronger inter-channel dependencies in this dataset, as indicated by Granger causality.
-The average rank of PatchTST across all datasets is 2.14 showing a significant gap of 0.57 compared to the second-best model, the channel-independent version of TSMixer.
-Linear models, meanwhile, remain competitive with recent forecasting approaches.
-The performance of DLinear on the standard datasets is comparable to more recent models.
-Across the 6 datasets, it achieves an average rank of 5.43 out of 10.
-Notably, on ETTh1, it ranks second, trailing PatchTST by only a small margin.
-6.2. Granger Causality for Evaluation of Channel Correlations
-Based on the findings from the previous section, which showed that multivariate models do not consistently outperform univariate ones on the standard datasets, we turn to a statistical evaluation of channel correlations across a broader set of datasets.
-Our goal is to identify benchmarks where inter-channel dependencies exist, potentially finding datasets and evaluation setup where multivariate models are needed.
-Although the full test results are presented in the appendix, we focus here on a representative comparison between the ODE datasets and the standard datasets, to highlight key patterns.
-We conduct Granger causality tests using three different lag values: 30, 96, and 192.
-These lags correspond to different lookback window sizes used in the autoregressive models underlying the tests.
-To further analyze channel dependence across a broad range of time series domains, we computed Granger causality values for 41 real-world datasets from the extensive Monash Time Series Forecasting Archive (Godahewa et al., 2021).
-These results, calculated for lag values of 7, 30, 96, and 192, are provided in the appendix.
-Throughout this section, we use the terms lag and lookback window interchangeably.
-Varying the lag allows us to analyze how inter-channel correlations evolve with different temporal resolutions.
-We can see the F-scores and the percentage of pairs of channels that exhibits granger causality in Table˜2.
-For both metrics, the higher values indicate higher correlation in the datasets.
-As seen in the table, the ODE datasets exhibit significantly higher F-scores and pass rates across all lag values, strongly indicating the presence of meaningful inter-channel dependencies.
-On contrary, the standard datasets show considerably weaker correlations, with average F-scores dropping below 1.0 for ETTm2 dataset at the highest lag.
-These findings align with our earlier empirical observations: univariate models often match or outperform multivariate models on the standard datasets due to weak inter-channel coupling.
-Moreover, we observe that increasing the lag generally reduces the measured correlation for both dataset types.
-This effect is particularly pronounced for the standard datasets, suggesting that longer lookback windows benefit more univariate models than multivariate ones, since the former cannot leverage strong inter-channel signals.
-In contrast, the ODE datasets show much stronger correlations at longer lags, e.g., average F-scores of 88.5 (lag 96) and 35.29 (lag 192), compared to 1.40 and 1.23 for the standard datasets.
-These results support using the ODE datasets as a valuable testbed for evaluating models under strong cross-channel dynamics.
-6.3. Evaluation on the ODE Benchmark
-As discussed in Section˜6.1, our results reaffirm DLinear as a competitive baseline and also suggest the superiority of CI models over their CD counterparts.
-In this section, we evaluate how the strong correlations between channels in the ODE datasets affect model performance.
-Table˜3 presents the results obtained using the same evaluation protocol described in Section˜6.1.
-The evaluation on the ODE datasets reveals a markedly different performance hierarchy among models.
-First of all, CD variants clearly outperform CI ones.
-The CD variants of the evaluated models consistently outperform their CI counterparts, highlighting the importance of modeling inter-channel dependencies, especially in datasets governed by coupled differential equations.
-As shown in Table˜4, CD models outperform CI versions in 18 out of 24 total experiments.
-This strongly supports the claim that model architecture should be tailored to the underlying data structure.
-Secondly, Linear models perform poorly on the ODE benchmark.
-Contrary to their performance on standard datasets, linear models such as DLinear are no longer competitive here.
-Averaged over all forecasting horizons, DLinear ranks worst in 5 out of the 6 ODE datasets, with an average rank of 9.83.
-Compared to Crossformer, DLinear suffers from performance drops ranging from 45% to 118%, with a median degradation of 58.3% and a mean of 74%.
-In summary, the commonly held view that linear models are competitive with deep architectures does not generalize to datasets with strong inter-channel dynamics.
-Our findings underline the importance of evaluating forecasting models on datasets with diverse structural properties.
-In cases where inter-channel correlation is intrinsic to the data—as with ODE systems—model architectures must be explicitly designed to capture these dependencies in order to perform well.
-6.4. Lookback Window Analysis
-In this section, we extend our evaluation by analyzing the distribution of selected lookback window values, grouped by both datasets and models.
-This analysis highlights how an optimal lookback window varies significantly depending on the model type and dataset characteristics.
-As shown in Figure˜3, we report the frequency of best-performing lookback window across CI models, CD models, and datasets.
-In the Figure 3(a), CI models exhibit substantial variation in their preferred lookback windows.
-For instance, DLinear consistently favors longer windows (336 or 720), while shorter windows are rarely optimal.
-This finding is soldified by the fact that overall, CI models choose 96 lookback window only 11% times.
-In contrast, Figure 3(b) shows that CD Models more often benefit from shorter windows, with 96 or 192 selected in 55% of the cases.
-These counts are aggregated across the four forecasting horizons (96, 192, 336, and 720).
-Figure 3(c) shows that optimal lookback windows also vary widely by dataset.
-This aligns with expectations: for hourly datasets like ETTh1 and ETTh2, a window of 96 covers four days, often sufficient for effective forecasting.
-In contrast, for 15-minute interval datasets like ETTm1 and ETTm2, more timesteps are required to capture similar temporal spans.
-Accordingly, ETTm2 frequently favors longer windows larger than or equal to 336, while ETTh2 consistently prefers windows shorter than 336.
-Overall, this analysis reinforces the importance of tuning the lookback window, not just across model types, but also in response to dataset-specific temporal characteristics.
-For per-horizon results, refer to the bar charts in the appendix.
+6.4 Lookback Window Analysis
+This section analyzes the distribution of optimal lookback window values by dataset and model, showing significant variation.
+Figure˜3 shows that CI models favor longer windows (336 or 720), with 96 chosen only 11% of the time.
+In contrast, CD models often benefit from shorter windows (96 or 192) in 55% of cases.
+Theoretically, this divergence suggests a trade-off between temporal depth and inter-channel noise.
+CI models, treating each channel in isolation, require longer historical contexts to compensate for the lack of cross-variable information.
+Conversely, CD models extract relevant signals from other channels, allowing them to achieve high accuracy with shorter temporal windows.
+Furthermore, for CD models, excessively long windows may introduce redundant inter-channel noise, making it harder for the model to distinguish true dependencies from spurious correlations across variables.
+Comparing across datasets, hourly datasets (ETTh1 and ETTh2) typically prefer shorter windows around 96, covering approximately four days, while 15-minute interval datasets (ETTm1 and ETTm2) favor longer windows (greater than 336) to capture equivalent temporal spans.
+This highlights that while temporal span is a physical requirement of the data, the optimal window length is a function of the model’s capacity to process spatial (channel) versus temporal information.
 (a) CI models
@@
 (c) Datasets
-Figure 3. Frequencies of best-performing lookback windows grouped by: (a) CI models, (b) CD models, and (c) datasets.
-7. Conclusion and Recommendations
-7.1. Conclusion
-Model ODE-6 Standard-7
-CI CD CI CD
-PatchTST 4 2 6 1
-TSMixer 0 6 7 0
-Crossformer 0 6 4 3
-TimeMixer 2 4 4 3
-Totals 6 18 21 7
-Table 4. Frequency of CI and CD variants outperforming each other across models and datasets.
-In this paper, as outlined in the introduction, our primary goal is to establish a fair and unbiased evaluation framework for multivariate TSF models.
-To achieve this, we make several critical adjustments to the experimental setup including: tuning the lookback window, analyzing statistical inter-channel correlations, and conducting empirical evaluations on datasets with varying characteristics.
-From this comprehensive setup, we draw the following key conclusions:
-  1. (1)
-    Tuning the lookback window is essential for fairness in multivariate TSF.
-Fixing the lookback window (commonly at 96) can unfairly disadvantage univariate models.
-Allowing models to tune this parameter ensures a more balanced comparison and reveals meaningful performance differences.
-This insight directly enables the second key finding below.
-  2. (2)
-    Modeling channel dependencies is unnecessary for standard datasets.
-For all models where a CI variant is available, it consistently outperforms the CD counterpart, by a margin of 21 to 7 on standard benchmarks.
-This is a significant result, especially considering that models like TSMixer, Crossformer, and TimeMixer are inherently designed to mix information across channels.
-Our findings indicate that inter-channel mixing is not a major contributor to performance on these datasets and can, in fact, degrade it (Table˜4).
-  3. (3)
-    Dataset nature plays a critical role in multivariate TSF.
-This aspect has been largely overlooked in prior work.
-Our experiments (Table˜4) reveal that datasets with different underlying dynamics require different modeling strategies.
-For instance, ODE Datasets exhibit significantly higher inter-channel correlation, as confirmed by statistical tests.
-This justifies the use of multivariate models in such cases and is further supported by the superior performance of channel-dependent models on this dataset family.
-7.2. Recommendations for Multivariate Time-Series Forecasting Evaluation
-In this paper, we strive to come with a useful insights that can advance the field of multivariate time series forecasting and this was done through a detailed experiments and statistical analysis from which we draw the following recommendations:
-  • Do not use a fixed lookback window when a vast amount of historical data is available to ensure full potential of all models considered as seen in Section˜6.1 and Section˜6.3.
-  • For proper evaluation and before choosing model architecture, consider statistically testing the channel correlation of the datasets as shown in Section˜6.2, it can help a lot in making an informed decision between deploying a CD or a CI model.
-  • As noted in Section˜6.4, shorter lookback window favors in general CD models, which means for applications with limited amount of historical data, an architecture capturing channel dependency is generally recommended.
-Appendix A Description of All Datasets
-In this section, we would like to show in detail the different datasets used in our experiments.
-Starting with the standard datasets which were explained a lot in literature, so we will just give a brief description of them.
-First ETT- datasets which representing indication of long term power deployment which are recorded for 2 different granalurities (15 mins for ’m’ and 1 hour for ’h’) for 2 different counties (hence the post-fix ’1’ and ’2’).
-Electricity dataset follows the consumption of electricity over an extended amount of time (3 years) recorded hourly for 321 different households.
-Traffic dataset is representing the flow of traffic recorded through multiple sensors in an hourly frequency spanning over a period of approximately 2 years.
-All of this information as well as the train/validation/test split is shown in Table˜5
-Dataset/Feature Channels Granularity Train/val/test Total Timesteps
-ETTh (1/2) 7 1 hour 12/4/4 (months) 17420
-ETTm (1/2) 7 15 mins 12/4/4 (months) 69680
-electricity 321 1 hour 7:1:2 26304
-traffic 862 1 hour 7:1:2 17544
-weather 21 10 mins 7:1:2 52696
-Table 5.
-Basic statistics about the standard datasets widely used in the time series forecasting literature.
-This table shows the number of features, frequency of recording the data, train/validation/test split as well as the total timesteps recorded in a given dataset.
-Second type of datasets used in our experiments is based on the chaotic ODE Benchmark.
-Here, we elaborate more on the general common features of the used ODE datasets and we give further description of the underlying processes for generating them.
-For the generation process, we use a common protocol for the chosen 6 attractors which is generating 20 points per time unit, splitting the data into 7:1:2 train:validation:test split which is the common practice for the standard datasets as well.
-Finally, we generate a very long time series 60000 timesteps which is divided into time series forecasting samples though a rolling window (same approach that is adopted on TSF papers for the standard datasets).
-For reproducibility, we provide the code stating the random seed that should be used to produce the same datasets and subsequently the same results we got.
-For understanding more the chosen processes as well as the dimensionality of each dataset, please refer to the chaotic ODE Benchmark paper (Gilpin, 2021).
-Appendix B Hyperparameter Tuning
-We evaluate various Transformer-based and MLP-based time series forecasting models (e.g., DLinear, TSMixer, PatchTST, Crossformer, and our proposed FaCT) under a unified set of hyperparameter search ranges as shown in Table Table˜6.
+Figure 3: Frequencies of best-performing lookback windows grouped by: (a) CI models, (b) CD models, and (c) datasets.
+7 Conclusion and Recommendations
+Our study establishes a fair evaluation framework for multivariate time series forecasting by tuning lookback windows, analyzing inter-channel correlations, and assessing diverse datasets.
+From this, we conclude and recommend the following:
+  1. 1.
+    Lookback window tuning is essential.
+Fixing the lookback window (commonly at 96) biases results against univariate models and masks true performance differences.
+Recommendation: Always allow models to optimize lookback windows to ensure balanced and meaningful comparisons.
+  2. 2.
+    CD modeling is unnecessary for standard datasets.
+CI models generally outperform CD models on standard datasets by a margin of 21 to 7, suggesting that mixing information across channels can negatively impact performance.
+Recommendation: For standard datasets, CI models should be preferred as they typically provide better results.
+  3. 3.
+    Dataset characteristics dictate model design.
+Datasets with strong inter-channel correlations, such as ODE datasets, require CD models capturing these dependencies, supported by superior CD performance and statistical tests.
+Recommendation: Pre-evaluate channel correlations using statistical tests to select appropriate model architectures and use CD models on datasets exhibiting strong inter-channel dynamics.
+  4. 4.
+    Lookback window length interacts with channel dependency.
+CD models benefit from shorter lookback windows, while CI models may favor longer ones, reflecting differing temporal dependency needs.
+Recommendation: Tailor lookback window length jointly with model type and dataset temporal resolution to maximize forecasting accuracy.
+Limitations and Future Work:
+While our framework ensures a fair comparison, it is subject to several limitations.
+First, our evaluation relies on a predefined discrete set of lookback windows $\{96,192,336,512,720\}$; while this aligns with standard practice in LTSF literature, it remains a pragmatic subset of all possible temporal spans.
+Second, our primary reliance on MSE, while standard for penalizing large outliers, may not fully capture model performance in contexts where distributional or scale-invariant metrics are prioritized.
+Furthermore, while this study leverages chaotic ODE simulations to provide a mathematically verifiable ground truth for channel dependency, future work will explore whether these advantages transfer to highly stochastic real-world domains—such as high-frequency financial markets or complex human behavior analysis—to further validate the broader applicability of our findings.
+Finally, our observations regarding CI and CD model performance are conditioned on the current state-of-the-art model families; the emergence of hybrid architectures may shift these dynamics in future research.
+Appendix 0.A Description of All Datasets
+In this section, we provide a detailed description of the various datasets used in our experiments.
+Standard Datasets
+We first utilize standard benchmarks widely recognized in the time series forecasting literature.
+The ETT (Electricity Transformer Training) datasets represent long-term power deployment indicators, recorded at two granularities: 15-minute intervals (denoted by ’m’) and 1-hour intervals (denoted by ’h’) across two different counties (1 and 2).
+The Electricity dataset tracks hourly consumption for 321 households over three years.
+The Traffic dataset represents traffic flow recorded by 862 sensors at an hourly frequency over two years.
+Finally, the Weather dataset includes 21 meteorological features recorded every 10 minutes.
+Statistics for these datasets are summarized in Table 5.
+Table 5: Basic statistics about the standard datasets used in the time series forecasting literature.
+Dataset Channels Granularity Split (T/V/T) Total Steps
+ETTh (1/2) 7 1 hour 12/4/4 (months) 17,420
+ETTm (1/2) 7 15 mins 12/4/4 (months) 69,680
+Electricity 321 1 hour 7:1:2 26,304
+Traffic 862 1 hour 7:1:2 17,544
+Weather 21 10 mins 7:1:2 52,696
+Chaotic ODE Benchmark
+The second category of datasets is based on the Chaotic ODE Benchmark.
+We follow a common protocol for the six chosen attractors, generating 20 points per time unit and utilizing a 7:1:2 train:validation:test split.
+We generate a long time series of 60,000 timesteps, divided into forecasting samples via a rolling window.
+For reproducibility, we provide the code and random seeds required to generate these datasets.
+For further details on the dimensionality and underlying processes, we refer to the Chaotic ODE Benchmark paper [5].
+Appendix 0.B Hyperparameter Tuning
+We evaluate various Transformer-based and MLP-based models under a unified hyperparameter search space (Table 6).
+Table 6: Common Hyperparameter Search Ranges. Values are integers except for learning rate and dropout (floats).
 Hyperparameter Search Range
 Learning Rate $10^{-7},10^{-2}]$
-Hidden Dimension (d_modeld\_model) $\{128,256,512,1024\}$
-Feedforward Dimension (d_ffd\_ff) $\{128,256,512,1024\}$
-# Encoder/Mixer Layers [1, 10][1,\,10]
-Dropout [0, 0.9][0,\,0.9]
+Hidden Dimension ($d_{model}$) $\{128,256,512,1024\}$
+Feedforward Dimension ($d_{ff}$) $\{128,256,512,1024\}$
+# Encoder/Mixer Layers $[1,10]$
+Dropout $[0,0.9]$
 Sequence Length (seq_lenseq\_len) $\{96,192,336,512,720\}$
-Table 6. Common Hyperparameter Search Ranges with all parameters being integer datatypes except for learning rate and dropout which span the whole range of floats on the respective range of values
-For specific architectures, additional parameters that are model specific were tuned on the same range of parameters specified on the respective papers.
-For all the range $[]$ or set of values $\{\}$, refer to Table˜7.
+Model-specific parameters were tuned within the ranges suggested by their respective original papers (Table 7).
+Best-performing hyperparameter sets for each model-dataset pair are included in the supplementary materials to ensure full reproducibility.
+Table 7: Model Specific Hyperparameter Search Ranges (all values are integers).
 Model Hyperparameter Range
@@
 TSMixer Hidden Size $\{32,64,256,1024\}$
-Crossformer/FaCT Segement Length $[3,12]$
+Crossformer/FaCT Segment Length $[3,12]$
 Baseline $\{0,1\}$
 Cross Factor $[3,20]$
-Table 7. Model Specific Hyperparameter Search Ranges/set of possible values with all of them having a datatype of integer in this case.
-To ensure easy and full reproducibility of our experiments, we include all the best performing hyperparameter sets corresponding to all models and datasets.
-Please see tables from Table˜16 to Table˜28 for the full list.
-We hope inclusion of these lists, make reproducibility of this paper’s results straight forward.
-Appendix C Implementation Details
-C.1. Libraries and Hardware
-All models were implemented using the PyTorch library (version 2.4) in Python (version 3.12.1).
-For hyperparameter optimization, we employed Optuna (version 3.6.1).
-Experiments were conducted on machines equipped with NVIDIA RTX 4090 and RTX 3090 GPUs.
-C.2. Training Procedure
-All models are trained using the Adam optimizer.
-We use mean squared error (MSE) as the default loss function, and each epoch iterates over mini-batches of size 32 by default.
-Training proceeds for a maximum of 100 epochs, or until early stopping is triggered if the validation loss fails to improve for 10 consecutive epochs.
-We use Optuna to conduct 20 trials per model, selecting configurations from the defined set of hyperparameters B that yield the best validation performance.
-C.3. Granger Causality
-Here, we list in more details the procedure to do the granger causality test which, alongside our code, should guide an easy usage of the framework for other datasets as well as easy reproduction of current results.
-We pick directly the first 1000 timesteps from the loaded dataset for this statistical test purposes, which we think is reasonable sample size with regard to the lags considered here.
-First, we check the pearson correlation between each pair of channels, and if it is too high (greater than 0,95 in our case), then one of the two channels is dropped.
-This is done, because channels which are very similar can cause results that are misleading when testing pair of channels that are identical (that can easily yield a very high score that will push the average scores way higher than expected).
-After, the initial filtering, we get dataset with diverse channels that can still have strong correlation, but one more important aspect we address next is stationarity as grangar causality works best on stationary time series data (i.e.
-mean and variance should not be shifting over time).
-Here, we apply the augmented Dickey-Fuller test on each channel of the time series to check if it is stationary.
-If the channel is not stationary, we apply differencing which should help bring stationarity into the time series channel.
-Now we have our preprocessed data which is ready for the grangar causality to be applied on.
-We use the packages from statsmodels package (Seabold and Perktold, 2010).
-Appendix D Derivation of CI/CD version of Models
-In our paper, 6 strong recent baselines are used throughout evaluation, they vary in complexity and structure.
-The aim of this appendix section is to ease the understanding of making CI and CD versions of the different baselines.
-Simpler Linear/MLP-based Models
-We start with the simpler models used in our experiments which are DLinear and TSMixer.
-DLinear as mentioned earlier is based on being the simplest model that applies a trend seasonality deceomposition followed by simpler linear layers applied on both the trend and the seasonal components (Zeng et al., 2023).
-For TSMixer, the authors had implemented a CD version which is the original version of the model and removed the channel mixing component which is called in their paper TMix-Only.
-We reimplemnted the model using pytorch and made sure of reproducing the paper results before using both versions in our experiments.
-Throughout our paper, we call both versions as TSMixer (CD) and TSMixer (CI) for the CD and CI versions respectively.
+Appendix 0.C Implementation Details
+0.C.1 Libraries and Hardware
+All models were implemented using PyTorch (v2.4) in Python (v3.12.1).
+Hyperparameter optimization was conducted via Optuna (v3.6.1).
+Experiments were performed on machines equipped with NVIDIA RTX 4090 and RTX 3090 GPUs.
+0.C.2 Training Procedure
+Models were trained using the Adam optimizer with Mean Squared Error (MSE) loss.
+The default mini-batch size was 32.
+Training lasted for a maximum of 100 epochs, with early stopping triggered if validation loss did not improve for 10 consecutive epochs.
+0.C.3 Granger Causality
+We utilize a Granger causality framework to analyze inter-channel dependencies.
+We extract the first 1,000 timesteps for this statistical test.
+Initially, we filter out highly redundant channels (Pearson correlation >0.95>0.95).
+We then apply the Augmented Dickey-Fuller (ADF) test to ensure stationarity, applying differencing where necessary.
+The causality tests are performed using the statsmodels package [20].
+Appendix 0.D Derivation of CI/CD Model Variants
+We evaluated six recent baselines, creating both Channel-Independent (CI) and Channel-Dependent (CD) versions where applicable.
+Simpler Linear/MLP Models
+DLinear remains a CI baseline due to its decomposition approach.
+For TSMixer, we implemented a CI version by removing the channel-mixing component (referred to as TMix-Only in the original work).
+Both versions are denoted as TSMixer (CI) and TSMixer (CD).
 PatchTST
-Now we move into transformer-based models, starting with PatchTST as one of the first transformer models to re-establish a boost of performance over the DLinear model.
-In their paper, the original version of the model is CI where the transformer encoder backbone is applied independently on different channels before the application of a linear projection layer.
-For the CD version, they briefly discuss it in the appendix of applying the same backbone but on a flattened dimension of both number of patches and channels which means the attention is now applied jointly over patches and channels.
-We use the original implementation of the paper but we had to add the part related to CD implementation ourselves as it was not readily available on the paper at the time of writing our paper.
-iTransformer/Crossformer
-Next, transformer model that proved to be a strong baseline is Crossformer which applies a Depth-Segment-Wise embedding before application of a Two Stage Attention (TSA).
-Finally a decoder layer is applied through cross attention between the positional embedding of output time info and encoder output.
-The TSA is applied on the segments dimension (temporal) as well as followed by the channel dimension which constitutes the original CD version of the model (Zhang and Yan, 2023).
-For a CI version, we just removed the second stage of the TSA to make the attention only applied on the segments dimension.
-We used authors implementation of the CD version and made the mentioned edits on the implementation to produce the CI version used in our experiments.
-For iTransformer (Liu et al., 2024), we don’t introduce additional CI version to the existing CD version as otherwise the model will be simplified to a linear model which is already covered in this paper.
+The original PatchTST is CI. For our CD variant, we apply the transformer backbone to a flattened dimension of both patches and channels, enabling joint attention across both.
+Crossformer
+The standard CD version uses Two-Stage Attention (TSA) over temporal segments and channels.
+Our CI version removes the second stage of TSA to restrict attention to the temporal dimension only.
+For iTransformer, we utilize only the CD version, as a CI version would simplify the model into a basic linear projection already covered by other baselines.
 TimeMixer
-One more prominent model proposed recently which proves be competitive on the standard datasets is TimeMixer (Wang et al., 2024a).
-In this model, the authors provide an MLP-based approach which is applied over multiple resolutions of the time series.
-This model original setting is CD where the MLP-mixing is done over both temporal and channel dimensions.
-For a CI version, the channel mixing component can be removed yielding a multiresolution temporal mixing based model.
-We use the authors implementation for both the CI and CD versions of the model.
-Appendix E Detailed Results
-In this appendix, we add the full results over all forecasting horizons for both types of introduced datasets.
-In Table˜9, we can see in detail the results for the chaotic ODE datasets, confirming the results from the main paper that Crossformer is clearly the SoTA on this benchmark with it being the best model in 15 out of possible 24 settings(for the CD version).
-For the standard datasets, the results are detailed in Table˜8 which also confirms the state of the current datasets where no model is clearly the best even emphasized with DLinear being the best model on 5 occurrences.
-This also fits the picture that was formulated on the table from the main paper text that when proper tuning is carried out for lookback window as well as all related hyperparameters, PatchTST remains competitive being the best overall 6 times.
-To conclude how comprehensive and reproducible our experiments are, these results represents the mean of 5 random seeds {3001, 3002, 3003, 3004, 3005} MSE error on the test split of each dataset.
-Dataset / Model Horizon PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
-CI CD CI CD CI CD CI CD CI CD
+Originally a CD model utilizing multi-resolution MLP-mixing. We derived the CI version by removing the channel-mixing components while maintaining the multi-resolution temporal mixing structure.
+Appendix 0.E Detailed Results
+Full forecasting results across all horizons are presented in Table 9 (ODE datasets) and Table 8 (Standard datasets).
+Results represent the mean MSE over five random seeds: $\{3001,3002,3003,3004,3005\}$.
+Table 8: Full forecasting results (MSE) over standard datasets. Best result per model variant in bold; best overall result highlighted in blue.
+PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
+Dataset H CI CD CI CD CI CD CI CD CI CD
 ETTh1 96 0.375 0.381 0.374 0.404 0.396 0.400 0.372 0.413 0.391 0.426
@@
 720 0.359 0.336 0.330 0.354 0.335 0.342 0.316 0.319 0.342 0.442
-Electricity 96 0.129* 0.141 0.131 0.137 0.139 0.150 0.135 0.135 0.130 0.143
+Elect. 96 0.129* 0.141 0.131 0.137 0.139 0.150 0.135 0.135 0.130 0.143
 192 0.147* 0.156 0.149 0.156 0.177 0.168 0.149 0.151 0.149 0.191
@@
 336 0.385* OOM 0.407 0.423 OOM 0.530* 0.436 OOM OOM 0.385*
-720 0.43* OOM 0.443 0.457 OOM 0.573* 0.466 OOM OOM 0.43*
-Table 8.
-In this table, the full results over different horizons are shown over all the 7 standard datasets.
-These results were average of 5 different random seeds {3001,3002,3003,3004,3005}.
-The best result for each variant per model is in bold, while the best results overall for each horizon per dataset is highlighted in blue.
-we reported the results with "*" from TimeMixer paper from the experiment where they tune the lookback window as well as the hyperparameters (please refer to appendix E in (Wang et al., 2024a)).
-Dataset / Model Horizon PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
-CI CD CI CD CI CD CI CD CI CD
+720 0.430* OOM 0.443 0.457 OOM 0.573* 0.466 OOM OOM 0.430*
+Table 9: Full forecasting results (MSE) for Chaotic ODE datasets.
+PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
+Dataset H CI CD CI CD CI CD CI CD CI CD
 Lorenz 96 0.658 0.658 0.747 0.729 0.331 0.266 0.884 0.433 0.398 0.338
@@
 720 0.965 0.970 0.983 0.971 0.920 0.914 0.978 0.954 0.962 0.929
-BlinkingRotlet 96 0.162 0.156 0.322 0.228 0.064 0.045 0.374 0.112 0.116 0.076
+Blink.Rot. 96 0.162 0.156 0.322 0.228 0.064 0.045 0.374 0.112 0.116 0.076
 192 0.322 0.321 0.510 0.429 0.230 0.210 0.504 0.299 0.272 0.315
@@
 720 0.931 0.939 0.970 0.963 0.825 0.844 0.984 0.894 0.867 0.892
-DoublePendulum 96 0.322 0.278 0.541 0.461 0.090 0.083 0.667 0.314 0.250 0.110
+Doub.Pend. 96 0.322 0.278 0.541 0.461 0.090 0.083 0.667 0.314 0.250 0.110
 192 0.551 0.594 0.713 0.753 0.418 0.466 0.762 0.534 0.440 0.380
@@
 720 0.727 0.654 0.778 0.742 0.721 0.655 0.886 0.663 0.699 0.717
-LorenzCoupled 96 0.696 0.587 0.840 0.745 0.372 0.276 0.919 0.610 0.458 0.818
+Lor.Coupl. 96 0.696 0.587 0.840 0.745 0.372 0.276 0.919 0.610 0.458 0.818
 192 0.883 0.901 0.949 0.895 0.687 0.635 0.959 0.840 0.795 0.947
@@
 720 0.990 1.007 1.013 0.994 0.926 0.942 0.996 0.984 0.969 1.003
-Table 9.
-In this table, the full results over different horizons are shown over all the 6 chaotic ODE datasets.
-These results were average of 5 different random seeds {3001,3002,3003,3004,3005}.
-The best result for each variant per model is in bold, while the best results overall for each horizon per dataset is highlighted in blue.
-we reported the results with "*" from TimeMixer paper from the experiment where they tune the lookback window as well as the hyperparameters (please refer to appendix E in (Wang et al., 2024a)).
-Dataset / Model PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
-CI CD CI CD CI CD CI CD CI CD
+Table 10: Standard deviation values averaged over the 4 used horizons. "—" denotes Out of Memory (OOM) errors.
+PatchTST TSMixer CrossFormer DLin. iTrans. TimeMixer
+Dataset CI CD CI CD CI CD CI CD CI CD
 Lorenz 0.0033 0.0025 0.0020 0.0115 0.0002 0.0095 0.0065 0.0303 0.0267 0.0066
-BlinkingRotlet 0.0063 0.0072 0.0063 0.0083 0.0023 0.0171 0.0085 0.0580 0.0146 0.0294
+Blink.Rotlet 0.0063 0.0072 0.0063 0.0083 0.0023 0.0171 0.0085 0.0580 0.0146 0.0294
 CellCycle 0.0030 0.0040 0.0030 0.0047 0.0000 0.0093 0.0163 0.1127 0.0126 0.0818
-DoublePendulum 0.0070 0.0067 0.0070 0.0305 0.0005 0.0110 0.0088 0.0136 0.0026 0.0052
+Doub.Pend. 0.0070 0.0067 0.0070 0.0305 0.0005 0.0110 0.0088 0.0136 0.0026 0.0052
 Hopfield 0.0057 0.0070 0.0057 0.0075 0.0007 0.0049 0.0129 0.0262 0.3732 0.0076
-LorenzCoupled 0.0030 0.0045 0.0030 0.0075 0.0000 0.0127 0.0056 0.0565 0.0088 0.0088
+Lor.Coupled 0.0030 0.0045 0.0030 0.0075 0.0000 0.0127 0.0056 0.0565 0.0088 0.0088
 ETTh1 0.0030 0.0092 0.0030 0.0015 0.0002 0.0425 0.0222 0.0156 0.0012 0.0175
 Weather 0.0058 0.0010 0.0033 0.0053 0.0004 0.0061 0.0049 0.0172 0.0037 0.0039
-Electricity OOM OOM 0.0003 0.0020 0.0000 0.0041 0.0124 0.0053 0.0017 0.0037
+Electricity 0.0003 0.0020 0.0000 0.0041 0.0124 0.0053 0.0017 0.0037
 ETTh2 0.0040 0.0047 0.0040 0.0088 0.0011 0.1265 0.1307 0.0153 0.0036 0.0389
@@
 ETTm2 0.0010 0.0020 0.0010 0.0040 0.0004 0.0928 0.0364 0.0007 0.0022 0.1017
-Traffic OOM OOM 0.0007 0.0025 0.0000 OOM OOM OOM OOM OOM
-Table 10.
-In this table, all standard deviation values are averaged over the 4 used horizons for all datasets.
-For each horizon, the standard deviation results were acquired by running the model for 5 different seeds with the following values {3001,3002,3003,3004,3005}
-. Note for some heavier models and datasets, we got Out of Memory error, specifically speaking some of the runs for electricity and traffic datasets.
-Appendix F Detailed Granger Causality Results
-We report results across three dataset groups (i) standard benchmark datasets, (ii) chaotic ODE datasets, and (iii) the large-scale Monash forecasting collection.
-For datasets with a large number of columns, we restrict the Granger causality analysis to pairwise comparisons among 20 randomly selected channels.
-For the standard benchmark datasets such as electricity and traffic, Granger causality tests reveal that these datasets exhibit moderate inter-channel correlation, with higher causality (F-score) predominantly at shorter lookback windows (lag=30) (Table˜11Table˜12, and Table˜14).
-This suggests that, for short lags, information from other channels can slightly aid forecasting, but as the lag increases, most predictive power lies within each channel’s own history.
-This outcome aligns with the observation that univariate models (CI) can often perform well for these datasets at longer lookback settings, as there is diminishing marginal benefit from including other channels when window size grows.
-In contrast, chaotic ODE datasets (Table˜11Table˜12, and Table˜14) consistently display higher Granger causality scores across all lag settings (with the exception of the electricity dataset), indicating robust and persistent dependencies between channels.
-This is particularly evident even at longer lookback windows.
-Such behavior is characteristic of deterministic physical systems or simulated multivariate dynamics where the evolution of one variable intrinsically influences the others.
-The implication is that multivariate models leveraging channel dependencies are especially suitable for these ODE datasets.
-We also report results for the Monash repository’s (Godahewa et al., 2021) large-scale real-world datasets (over 40 time series from diverse domains).
-We select 40 datasets from a total of 58 datasets after removing datasets with missing values.
-Granger causality analysis Table˜13 reveals that only a minority—six of forty—show an inter-channel F-score above 10, and that F-scores tend to decline with increasing lag (except isolated cases like the Kaggle dataset).
-This pattern suggests that most Monash datasets lack strong lateral dependencies and instead behave like collections of independent univariate series.
-For instance, many resemble traffic datasets from the M1 or M3 collections: aggregations rather than integrated multivariate systems.
-As lags increase (e.g., at lag=196), additional channels yield minimal information gain, hence the F-score drops.
-Overall, these findings highlight that Granger causality excels at disentangling the nature of multivariate time series dependencies.
-Standard benchmarks typically permit effective univariate treatment at larger lags, the ODE datasets consistently demand modeling of channel interactions, and most Monash datasets demonstrate little true multivariate dependence.
-These results underscore the value of focusing research on genuinely multivariate datasets, such as chaotic ODEs, to advance understanding and modeling of channel dependence in time series forecasting.
-Dataset / Metric Average F-score (30) $H_{0}$ rejected (%, p=0.05)
-Double Pendulum 39.60 100%
-Lorenz 1378.54 83%
-Lorenz Coupled 579.23 97%
-Cell Cycle 6.78 50%
-Blinking Rotlet 3.98 33%
-HopField 2.86 80%
-ETTm1 1.48 50%
-ETTm2 1.31 30%
-ETTh1 1.87 60%
-ETTh2 1.94 70%
-Weather 1.55 33%
-Electricity* (Ch=20) 4.34 93%
-Traffic* (Ch=20) 4.71 94%
-Table 11. Average F-scores and percentage of $H_{0}$ rejected (p=0.05) across datasets (lag=30).
-Dataset / Metric Average F-score (96) $H_{0}$ rejected (%, p=0.05)
-Double Pendulum 10.21 100%
-Lorenz 353.25 83%
-Lorenz Coupled 160.95 66%
-Cell Cycle 2.61 53%
-Blinking Rotlet 2.59 33%
-HopField 1.92 70%
-ETTm1 1.13 33%
-ETTm2 1.11 15%
-ETTh1 1.12 20%
-ETTh2 1.19 36%
-Weather 1.12 30%
-Electricity* (Ch=20) 1.92 84%
-Traffic* (Ch=20) 2.22 94%
-Table 12. Average F-scores and percentage of $H_{0}$ rejected (p=0.05) across datasets (lag=96).
+Traffic 0.0007 0.0025 0.0000
+Appendix 0.F Detailed Granger Causality Analysis
+We evaluate inter-channel dependencies across three distinct dataset groups:
+  1. (i)
+    Standard benchmark datasets (e.g., ETT, Weather);
+  2. (ii)
+    Chaotic ODE datasets (e.g., Lorenz, Double Pendulum);
+  3. (iii)
+    The large-scale Monash forecasting collection [6].
+For datasets with high dimensionality, we restrict the Granger causality analysis to pairwise comparisons among 20 randomly selected channels to maintain computational efficiency.
+Standard Benchmarks
+As shown in Table 11, datasets such as Electricity and Traffic exhibit moderate inter-channel correlation.
+However, higher causality ($F$-scores) is predominantly concentrated at shorter lookback windows (lag=30).
+As the lag increases to 192, the predictive power of neighboring channels diminishes significantly.
+This quantitative decline supports the observation that Channel-Independent (CI) univariate models often perform optimally at longer lookback settings, as the marginal information gain from other channels becomes negligible over time.
+Chaotic ODE Datasets
+In contrast, chaotic ODE datasets consistently display high $F$-scores across all lag settings.
+Even at a lag of 192, systems like the Lorenz and Double Pendulum maintain robust and persistent dependencies between channels.
+This is characteristic of deterministic physical systems where the state of one variable is intrinsically and permanently linked to the evolution of others.
+These results underscore why multivariate models (CD) are particularly well-suited for physical or simulated dynamics.
+Monash Forecasting Collection
+We analyzed 40 datasets from the Monash repository after filtering for missing values.
+As detailed in Table 12, the vast majority of these real-world datasets lack strong lateral dependencies.
+Only six of the forty datasets show an inter-channel $F$-score above 10, with most scores declining sharply as lag increases.
+Many of these datasets (e.g., Hospital or Australian Electricity) behave more like collections of independent univariate series rather than integrated multivariate systems.
+Summary
+Overall, Granger causality serves as a powerful diagnostic tool for disentangling the nature of time series dependencies.
+While standard benchmarks permit effective univariate treatment at larger lags, ODE datasets demand explicit modeling of channel interactions.
+These findings suggest that research into multivariate dependencies should prioritize genuinely integrated datasets, such as chaotic ODEs, rather than simple aggregations.
+Table 11: Average $F$-scores and percentage of $F$-tests passed ($p=0.05$) across different lag settings. Asterisks (*) indicate analysis performed on 20 randomly selected channels.
+Lag = 30 Lag = 96 Lag = 192
+Dataset F-score Passed % F-score Passed % F-score Passed %
+Double Pendulum 39.60 100% 10.21 100% 6.35 100%
+Lorenz 1378.54 83% 353.25 83% 126.38 83%
+Lorenz Coupled 579.23 97% 160.95 66% 72.09 57%
+Cell Cycle 6.78 50% 2.61 53% 1.58 53%
+Blinking Rotlet 3.98 33% 2.59 33% 3.76 33%
+HopField 2.86 80% 1.92 70% 1.55 50%
+ETTm1 1.48 50% 1.13 33% 1.04 23%
+ETTm2 1.31 30% 1.11 15% 0.99 0%
+ETTh1 1.87 60% 1.12 20% 1.07 7%
+ETTh2 1.94 70% 1.19 36% 1.06 17%
+Weather 1.55 33% 1.12 30% 1.82 32%
+Electricity* 4.34 93% 1.92 84% 1.21 43%
+Traffic* 4.71 94% 2.22 94% 1.39 69%
+Table 12: Average $F$-scores across 40 Monash datasets for lags 7, 30, 96, and 192. Empty cells or "—" indicate insufficient time points for statistical validity at that lag.
 Monash Dataset 7 30 96 192
-australian electricity demand dataset 29.38 18.06 2.71 1.62
-bitcoin without missing values 1.88 1.44 1.66 2.47
-car parts without missing values 1.25
+Australian electricity demand 29.38 18.06 2.71 1.62
+Bitcoin (no missing) 1.88 1.44 1.66 2.47
+Car parts (no missing) 1.25
 CIF 2016 1.90
@@
 Hospital 1.38
-kaggle web traffic without missing values 3.02 3.25 12.96 45.63
+Kaggle web traffic (no missing) 3.02 3.25 12.96 45.63
 Kaggle web traffic weekly 1.14 1.56
-kdd cup 2018 without missing values 2.48 1.53 1.18 1.08
-london smart meters without missing values 1.69 1.26 1.07 1.03
-m1 monthly 2.02
-m1 quarterly
-m1 yearly
-m3 monthly 1.59
-m3 other 1.46
-m3 quarterly
-m3 yearly
+KDD Cup 2018 (no missing) 2.48 1.53 1.18 1.08
+London smart meters (no missing) 1.69 1.26 1.07 1.03
+M1 Monthly / Quarterly / Yearly 2.02
+M3 Monthly / Other / Quarterly / Yearly 1.59
 M4 Daily 1.10 1.05
 M4 Hourly 18.06 3.48 1.96 1.20
-m4 monthly 1.76
-m4 quarterly 1.06
-m4 weekly 1.14
-m4 yearly
-nn5 daily without missing values 9.45 2.10 1.17 1.02
+M4 Monthly / Quarterly / Weekly 1.76
+NN5 Daily (no missing) 9.45 2.10 1.17 1.02
 NN5 Weekly 1.17 1.03
 Pedestrian counts 10.85 3.14 2.65
-rideshare without missing values 3.73 2.15 1.31
+Rideshare (no missing) 3.73 2.15 1.31
 Solar 10 minutes 8.63 3.87 1.52 1.56
 Solar weekly 0.78
-temperature rain without missing values 2.85 1.54 1.20 1.06
-tourism monthly 12.12 1.58
-tourism quarterly 1.98
-tourism yearly
+Temperature rain (no missing) 2.85 1.54 1.20 1.06
+Tourism Monthly / Quarterly / Yearly 12.12 1.58
 Traffic hourly 16.00 4.44 2.08 1.25
 Traffic weekly 1.12 1.21
-vehicle trips without missing values 1.38
+Vehicle trips (no missing) 1.38
 Weather 1.22 1.17 1.12 1.03
-Table 13. Average F-scores across 40 Monash datasets for lags 7, 30, 90, 192 .
-Dataset / Metric Average F-score (192) $H_{0}$ rejected (%, p=0.05)
-Double Pendulum 6.35 100%
-Lorenz 126.38 83%
-Lorenz Coupled 72.09 57%
-Cell Cycle 1.58 53%
-Blinking Rotlet 3.76 33%
-HopField 1.55 50%
-ETTm1 1.04 23%
-ETTm2 0.99 0%
-ETTh1 1.07 7%
-ETTh2 1.06 17%
-Weather 1.82 32%
-Electricity* (Ch=20) 1.21 43%
-Traffic* (Ch=20) 1.39 69%
-Table 14.
-Average F-scores and percentage of rejected $H_{0}$ (p=0.05) across datasets (lag=192).
-Missing values indicate that the Granger Causality failed due to statistically limited number of available time points for a proper analysis.
-Appendix G Full Lookback Window Analysis
-We show here the detailed results of best performing lookback windows on seperate figures for each of the forecasting horizons.
-We can see from Figure˜4, Figure˜5, Figure˜6, and Figure˜7 the analysis of the results based on forecasting horizons 96, 192, 336, and 720 respectively.
-We can see that on average the same insights noticed from the discussed forecasting horizon 96, still applies to the rest of experiments.
-The patterns discussed with respect to ETTh1, and ETTh2 tending to use smaller lookback windows still applies clearly on horizons 192 and 336.
-This changes when 720 horizon is deployed which makes sense as the difficulty of forecasting such a large amount of timesteps promotes the usage of a larger lookback windows for all datasets.
-Regarding the tendency of CI models to benefit more from a larger lookback window than CD models, that still holds for the different horizons as seen in the different figures.
-With also more models on the larger forecasting horizon, such as 720, tend to use larger lookback window more often.
-This pattern follows the same pattern we discussed for the performance on the ETTh datasets.
-The bottom line from this analysis is to show as well the complexity of forecasting different multivariate datasets across different horizons which makes the tuning of the lookback window clearly the standard approach to go for.
-Figure 4. Analysis of Frequencies of Best Performing Lookback Windows Grouped by : (a) datasets, (b) CI models, and (c) CD models. All recorded for the forecasting horizon 96.
-Figure 5. Analysis of Frequencies of Best Performing Lookback Windows Grouped by : (a) datasets, (b) CI models, and (c) CD models. All recorded for the forecasting horizon 192.
-Figure 6. Analysis of Frequencies of Best Performing Lookback Windows Grouped by : (a) datasets, (b) CI models, and (c) CD models. All recorded for the forecasting horizon 336.
-Figure 7. Analysis of Frequencies of Best Performing Lookback Windows Grouped by : (a) datasets, (b) CI models, and (c) CD models. All recorded for the forecasting horizon 720.
-Appendix H ODE Datasets Sample Visualization
-In this section, a visualization is shown for the rest of the 6 ODE datasets used in this research.
-In Figure˜8, a visualization of a 1000 timesteps from cell cycle based ODE is shown.
-Two different cyclin degradation activation are shown with channel 2 (middle top) activating the degradation of cyclin C1 levels (middle bottom).
-Same phenomena is shown clearly on the right 2 channels of the figure for cyclin C2 levels.
-In Figure˜9Figure˜10Figure˜11, and Figure˜12, we can see a visualization of the first 1000 time steps of the blinking rotlet, hopfield, lorenz and lorenz coupled time series based datasets.
-We can see clearly on these figures, how the dynamics of these different datasets generate complex yet real-world like time series data, with highs and lows from one channel triggering highs and lows in other channels.
-Figure 8. Time series visualization of the cell cycle ODE-based time series sample.
-Figure 9. Time series visualization of the blinking rotlet ODE-based time series sample.
-Figure 10. Time series visualization of the hopfield ODE-based time series sample.
-Figure 11. Time series visualization of the lorenz ODE-based time series sample.
-Figure 12. Time series visualization of the Lorenz Coupled ODE-based time series sample.
-Appendix I Cell Cycle ODE Discussion
-We discussed already the interesting properties of Double Pendulum ODEs.
-In this part of the appendix, we would like to extend more the discussion on the coupling between the ODEs and how variables interact across time.
-Equations 1-6 presents the system of ODEs, we used to generate the cell cycle dataset (Romond et al., 1999).
-These 6 equations show two different oscillators based on the concentration levels of cyclin ($C_{1}$ and $C_{2}$) and how they motivate the progression of the cell cycle.
-The cyclin concentration reflects directly on the activated Cyclin Dependent Kinases (cdk1, cdk2) which represented in the equations by $M_{1}$ and $M_{2}$ respectively.
-It also affect directly enzymes $X_{1}$ and $X_{2}$.
-$v_{ij}$ and $v_{dj}$ are constants that are multiplied in equation 1 for $j=1$ and equation 4 for $j=2$ as a rate of cyclin synthesis.
-In equations 1, we can see clearly the coupling between the ODE variables $C_{1}$ and $M_{2}$.
-Whereas, in equation 4, we can see the coupling between $C_{2}$ and $M_{1}$.
-All of this shows how information from one channel can cause spikes or change of trend in other channels, showing clear channel dependence in the generated time series data.
-Also, this section is valuable to show the very important application of such equation on a real life biomedical process which elevates the importance of our study more.
-Please note that these equations were taken from (Romond et al., 1999) and for more details, we encourage the reader to go through their work as well.
-(8) $\frac{dC_{1}}{dt}=\frac{v_{i1}K_{im1}}{K_{im1}+M_{2}}-\frac{v_{d1}X_{1}C_{1}}{K_{d1}+C_{1}}-k_{d1}C_{1}$
-(9) $\frac{dM_{1}}{dt}=\frac{V_{1}\left(1-M_{1}\right)}{K_{1}+\left(1-M_{1}\right)}-\frac{V_{2}M_{1}}{K_{2}+M_{1}}$
-(10) $\frac{dX_{1}}{dt}=\frac{V_{3}\left(1-X_{1}\right)}{K_{3}+\left(1-X_{1}\right)}-\frac{V_{4}X_{1}}{K_{4}+X_{1}}$
-(11) $\frac{dC_{2}}{dt}=\frac{v_{i2}K_{im2}}{K_{im2}+M_{1}}-\frac{v_{d2}X_{2}C_{2}}{K_{d2}+C_{2}}-k_{d2}C_{2}$
-(12) $\frac{dM_{2}}{dt}=\frac{U_{1}\left(1-M_{2}\right)}{H_{1}+\left(1-M_{2}\right)}-\frac{U_{2}M_{2}}{H_{2}+M_{2}}$
-(13) $\frac{dX_{2}}{dt}=\frac{U_{3}\left(1-X_{2}\right)}{H_{3}+\left(1-X_{2}\right)}-\frac{U_{4}X_{2}}{H_{4}+X_{2}}$
-(14) $V_{1}=\frac{C_{1}}{K_{c1}+C_{1}}V_{M_{1}},\qquad V_{3}=M_{1}\cdot V_{M_{3}}$
-(15) $U_{1}=\frac{C_{2}}{K_{c2}+C_{2}}U_{M_{1}},\qquad U_{3}=M_{2}\cdot UM_{3}$
-Appendix J Ablation Study on RevIN
-In this section, we include an important ablation study where we test the importance of using Reversible Instance Normalization (RevIN)  (Kim et al., 2021) on 2 standard datasets and 2 datasets from the chaotic benchmark.
-Crossformer was originally proposed without RevIN, so we added the RevIN component to both the CI and CD versions to measure how much of an effect window size this can have on different types of datasets.
-Moreover we included one more strong baseline in PatchTST in the study measuring the effect of RevIN on its main CI version.
-We got 2 interesting patterns that can be seen also from table  Table˜15 which are the following:
-  • Introducing revin helps to improve the performance on the both variants of Crossformer on the standard datasets.
-  • On the contrary on the chaotic ODEs the model without revin is performing better than its counterpart with revin.
-These findings shows that in general as confirmed in the literature RevIN can help in the performance of different model on standard datasets.
-While this being not totally true for PatchTST, it has been true for most models in the literature  (Chen et al., 2023).
-Important as emphasized throughout this paper, that is not easy task to generalize to different types of datasets, and that is no different for RevIN as it deteriorates performance if introduced on Crossformer on the chaotic ODE datasets.
-This shows that the different nature of Chaotic ODEs with the complex patterns can not benefit from the normalization+denormalization technique introduced in RevIN.
-Dataset / Model PatchTST_CI CrossFormer_CD CrossFormer_CI
-with RevIN w/o RevIN with RevIN w/o RevIN with RevIN w/o RevIN
+Appendix 0.G Full Lookback Window Analysis
+This section provides a granular analysis of optimal lookback window frequencies across all evaluated forecasting horizons ($H\in\{96,192,336,720\}$).
+Figures 4 through 7 detail these results, categorized by dataset, Channel-Independent (CI) models, and Channel-Dependent (CD) models.
+On average, the insights derived from the $H=96$ horizon remain consistent across the extended experiments.
+For instance, the tendency of ETTh1 and ETTh2 to utilize smaller lookback windows persists for horizons 192 and 336.
+However, at $H=720$, we observe a systemic shift toward larger lookback windows.
+This transition is intuitive; the increased difficulty of forecasting across an extended horizon promotes the use of broader historical context for all datasets.
+Furthermore, the observation that CI models typically benefit more from larger lookback windows than CD models holds true across horizons.
+As the forecasting horizon expands to 720, a greater number of models across all categories favor larger windows.
+These patterns underscore the complexity of multivariate time series forecasting and confirm that the systematic tuning of lookback windows is essential for optimal performance across diverse horizons.
+(a) By Dataset
+(b) By CI Models
+(c) By CD Models
+Figure 4: Frequency analysis of best-performing lookback windows for forecasting horizon 96.
+(a) By Dataset
+(b) By CI Models
+(c) By CD Models
+Figure 5: Frequency analysis of best-performing lookback windows for forecasting horizon 192.
+(a) By Dataset
+(b) By CI Models
+(c) By CD Models
+Figure 6: Frequency analysis of best-performing lookback windows for forecasting horizon 336.
+(a) By Dataset
+(b) By CI Models
+(c) By CD Models
+Figure 7: Frequency analysis of best-performing lookback windows for forecasting horizon 720.
+Appendix 0.H ODE Datasets Sample Visualization
+In this section, we provide visualizations for the remaining six ODE datasets utilized in this research.
+Figure 8 illustrates 1,000 timesteps from the Cell Cycle-based ODE.
+Specifically, two distinct cyclin degradation activations are observable: channel 2 (middle-top) activates the degradation of cyclin C1 levels (middle-bottom).
+This phenomenon is replicated on the rightmost two channels for cyclin C2 levels.
+Figures 9, 10, 11, and 12 display the first 1,000 timesteps for the Blinking Rotlet, Hopfield, Lorenz, and Lorenz-Coupled datasets, respectively.
+These visualizations demonstrate how deterministic dynamics generate complex, real-world-like multivariate patterns, where temporal fluctuations in one channel directly influence or trigger corresponding shifts in associated channels.
+Figure 8: Time series visualization of the cell cycle ODE-based time series sample.
+Figure 9: Time series visualization of the blinking rotlet ODE-based time series sample.
+Figure 10: Time series visualization of the hopfield ODE-based time series sample.
+Figure 11: Time series visualization of the lorenz ODE-based time series sample.
+Figure 12: Time series visualization of the Lorenz Coupled ODE-based time series sample.
+Appendix 0.I Cell Cycle ODE Discussion
+We have already discussed the interesting properties of Double Pendulum ODEs.
+In this part of the appendix, we would like to further extend the discussion on the coupling between the ODEs and how variables interact across time.
+Equations 1–6 present the system of ODEs we used to generate the cell cycle dataset [19].
+These six equations show two different oscillators based on the concentration levels of cyclin ($C_{1}$ and $C_{2}$) and how they motivate the progression of the cell cycle.
+The cyclin concentration reflects directly on the activated Cyclin-Dependent Kinases (cdk1, cdk2), which are represented in the equations by $M_{1}$ and $M_{2}$, respectively; it also directly affects enzymes $X_{1}$ and $X_{2}$.
+The terms $v_{ij}$ and $v_{dj}$ are constants multiplied in Equation 1 (for $j=1$) and Equation 4 (for $j=2$) as a rate of cyclin synthesis.
+In Equation 1, we can see clearly the coupling between the ODE variables $C_{1}$ and $M_{2}$, whereas in Equation 4, we see the coupling between $C_{2}$ and $M_{1}$.
+All of this demonstrates how information from one channel can cause spikes or changes of trend in other channels, showing clear channel dependence in the generated time-series data.
+Furthermore, this section is valuable for demonstrating the important application of such equations to a real-life biomedical process, which further elevates the importance of our study.
+Please note that these equations were taken from [19]; for more details, we encourage the reader to review their work.
+$\frac{dC_{1}}{dt}=\frac{v_{i1}K_{im1}}{K_{im1}+M_{2}}-\frac{v_{d1}X_{1}C_{1}}{K_{d1}+C_{1}}-k_{d1}C_{1}$ (7)
+$\frac{dM_{1}}{dt}=\frac{V_{1}(1-M_{1})}{K_{1}+(1-M_{1})}-\frac{V_{2}M_{1}}{K_{2}+M_{1}}$ (8)
+$\frac{dX_{1}}{dt}=\frac{V_{3}(1-X_{1})}{K_{3}+(1-X_{1})}-\frac{V_{4}X_{1}}{K_{4}+X_{1}}$ (9)
+$\frac{dC_{2}}{dt}=\frac{v_{i2}K_{im2}}{K_{im2}+M_{1}}-\frac{v_{d2}X_{2}C_{2}}{K_{d2}+C_{2}}-k_{d2}C_{2}$ (10)
+$\frac{dM_{2}}{dt}=\frac{U_{1}(1-M_{2})}{H_{1}+(1-M_{2})}-\frac{U_{2}M_{2}}{H_{2}+M_{2}}$ (11)
+$\frac{dX_{2}}{dt}=\frac{U_{3}(1-X_{2})}{H_{3}+(1-X_{2})}-\frac{U_{4}X_{2}}{H_{4}+X_{2}}$ (12)
+$V_{1}=\frac{C_{1}}{K_{c1}+C_{1}}V_{M_{1}},\qquad V_{3}=M_{1}\cdot V_{M_{3}}$ (13)
+$U_{1}=\frac{C_{2}}{K_{c2}+C_{2}}U_{M_{1}},\qquad U_{3}=M_{2}\cdot U_{M_{3}}$ (14)
+Appendix 0.J Ablation Study on RevIN
+In this section, we present an ablation study evaluating the impact of Reversible Instance Normalization (RevIN) [11] on two standard datasets and two datasets from the chaotic benchmark.
+Since Crossformer was originally proposed without RevIN, we integrated the RevIN component into both the Channel-Independent (CI) and Channel-Dependent (CD) versions to measure its effect across different dataset types.
+Additionally, we included PatchTST as a strong baseline to measure the effect of RevIN on its primary CI version.
+We observed two distinct patterns in the results, as detailed in Table 13:
+  • The introduction of RevIN improves the performance of both Crossformer variants on the standard datasets.
+  • Conversely, on the chaotic ODE datasets, the models without RevIN outperform their counterparts that include the normalization technique.
+These findings confirm that, as established in the literature [3], RevIN generally enhances model performance on standard datasets.
+While this was not universally true for PatchTST in our tests, it holds for most models in existing studies.
+Importantly, as emphasized throughout this paper, generalizing across different types of datasets is a non-trivial task; RevIN is no exception, as it deteriorates performance when applied to Crossformer on chaotic ODE datasets.
+This suggests that the complex, non-stationary patterns inherent in chaotic ODEs may not benefit from the normalization-denormalization technique introduced by RevIN.
+Table 13: Ablation study on Reversible Instance Normalization (RevIN).
+The results demonstrate that while RevIN improves performance on standard datasets (ETTh2, ETTm1), it can be detrimental to chaotic ODE datasets (Hopfield, LorenzCoupled).
+The best variant for each model is in bold, and the best overall result for each dataset is highlighted in blue.
+Forecasting horizon $L=96$.
+Dataset / Model PatchTST (CI) CrossFormer (CD) CrossFormer (CI)
++RevIN –RevIN +RevIN –RevIN +RevIN –RevIN
 Hopfield 0.151 0.156 0.058 0.049 0.073 0.059
@@
 ETTm1 0.314 0.313 0.339 0.364 0.297 0.309
-Table 15.
-This table shows an ablation study on the reversible instance normalization technique  (Kim et al., 2021) which showed to improve performance for most model on the standard datasets.
-The best variant of model is in bold and best overall for each dataset is highlighted in blue.
-These results are based on the forecasting horizon 96.
-Model Setting Best Hyper. Params
-PatchTST CI-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
-PatchTST CD-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
+Hyperparameter settings for the BlinkingRotlet dataset across different forecasting horizons. The table is split across pages for readability.
+Model Setting Best Hyperparameters
+[…]
+Model Setting Best Hyperparameters
+[…]
+PatchTST CD-96 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
 TSMixer CI-96 lr: 0.0001, num_blocks: 5, hidden_size: 32, dropout: 0.336, activation: relu, seq_len: 512
@@
 TimeMixer CD-96 lr: 0.0023, d_ff: 256, d_model: 128, e_layers: 5, seq_len: 192
-PatchTST CI-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
-PatchTST CD-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
-TSMixer CI-192 lr:0.0005,num_blocks:8,hidden_size:1.02e+03,dropout:0.705,activation:relu,seq_len:512
+PatchTST CI-192 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
+PatchTST CD-192 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
+TSMixer CI-192 lr: 0.0005, num_blocks: 8, hidden_size: 1024, dropout: 0.705, activation: relu, seq_len: 512
 TSMixer CD-192 lr: 0.0003, num_blocks: 1, hidden_size: 64, dropout: 0.205, activation: gelu, seq_len: 512
@@
 TimeMixer CD-192 lr: 0.0028, d_ff: 256, d_model: 128, e_layers: 4, seq_len: 192
-PatchTST CI-336 lr’:1.96e-07, ’e_layers’:5, ’d_ff’: 256, ’d_model’:1024, ’dropout’:0.295,’fc_dropout’:0.622, ’patch_size’:8,’stride’:8,’seq_len’:720
-PatchTST CD-336 lr’: 9.1666e-05, ’e_layers’: 6, ’d_ff’: 1024, ’d_model’: 1024, ’dropout’: 0.086, ’fc_dropout’: 0.5099, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
+PatchTST CI-336 lr: 1.96e-07, e_layers: 5, d_ff: 256, d_model: 1024, dropout: 0.295, fc_dropout: 0.622, patch_size: 8, stride: 8, seq_len: 720
+PatchTST CD-336 lr: 9.16e-05, e_layers: 6, d_ff: 1024, d_model: 1024, dropout: 0.086, fc_dropout: 0.509, patch_size: 16, stride: 8, seq_len: 192
 TSMixer CI-336 lr: 0.0012, num_blocks: 6, hidden_size: 32, dropout: 0.264, activation: relu, seq_len: 512
@@
 TimeMixer CD-336 lr: 0.0010, d_ff: 1024, d_model: 1024, e_layers: 10, seq_len: 192
-PatchTST CI-720 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
+PatchTST CI-720 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
 PatchTST CD-720 lr: 0.0003, e_layers: 9, d_ff: 256, d_model: 128, dropout: 0.228, fc_dropout: 0.072, patch_size: 8, stride: 8, seq_len: 336
-TSMixer CI-720 lr:0.0004,num_blocks:8,hidden_size:1.02e+03,dropout:0.754,activation:relu,seq_len:512
+TSMixer CI-720 lr: 0.0004, num_blocks: 8, hidden_size: 1024, dropout: 0.754, activation: relu, seq_len: 512
 TSMixer CD-720 lr: 0.0001, num_blocks: 2, hidden_size: 32, dropout: 0.297, activation: relu, seq_len: 512
@@
 TimeMixer CD-720 lr: 0.0021, d_ff: 256, d_model: 128, e_layers: 4, seq_len: 192
-Table 16. Hyperparameter settings for the BlinkingRotlet dataset. The best variant for each model is shown.
-Model Setting Best Hyper. Params
-PatchTST CI-96 lr:0.0003,e_layers:10,d_ff:256,d_model:128,dropout:0.081,fc_dropout:0.390,patch_size:8,stride:8,seq_len:336
+Hyperparameter settings for the CellCycle dataset. The table is split across pages for readability.
+Model Setting Best Hyperparameters
+[…]
+Model Setting Best Hyperparameters
+[…]
 PatchTST CD-96 lr: 0.0003, e_layers: 9, d_ff: 256, d_model: 128, dropout: 0.228, fc_dropout: 0.072, patch_size: 8, stride: 8, seq_len: 336
@@
 TimeMixer CD-192 lr: 0.0021, d_ff: 256, d_model: 128, e_layers: 4, seq_len: 192
-PatchTST CI-336 lr’: 0.00027, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 128, ’dropout’: 0.081, ’fc_dropout’: 0.390, ’patch_size’: 8, ’stride’: 8, ’seq_len’: 336
-PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.53896, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
+PatchTST CI-336 lr: 0.00027, e_layers: 10, d_ff: 256, d_model: 128, dropout: 0.081, fc_dropout: 0.390, patch_size: 8, stride: 8, seq_len: 336
+PatchTST CD-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
 TSMixer CI-336 lr: 0.0041, num_blocks: 10, hidden_size: 256, dropout: 0.415, activation: gelu, seq_len: 96
@@
 TimeMixer CD-336 lr: 0.0021, d_ff: 512, d_model: 1024, e_layers: 8, seq_len: 96
-PatchTST CI-720 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
-PatchTST CD-720 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
+PatchTST CI-720 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
+PatchTST CD-720 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
 TSMixer CI-720 lr: 0.0006, num_blocks: 8, hidden_size: 64, dropout: 0.181, activation: relu, seq_len: 192
@@
 TimeMixer CD-720 lr: 0.0022, d_ff: 1024, d_model: 512, e_layers: 8, seq_len: 96
-Table 17. Hyperparameter settings for the CellCycle dataset. The best variant for each model is shown.
-Model Setting Best Hyper. Params
-PatchTST CI-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
+Hyperparameter settings for the DoublePendulum dataset. The table is split across pages for readability.
+Model Setting Best Hyperparameters
+[…]
+Model Setting Best Hyperparameters
+[…]
 PatchTST CD-96 lr: 0.0003, e_layers: 9, d_ff: 256, d_model: 128, dropout: 0.228, fc_dropout: 0.072, patch_size: 8, stride: 8, seq_len: 336
@@
 TimeMixer CD-96 lr: 0.0018, d_ff: 1024, d_model: 128, e_layers: 6, seq_len: 192
-PatchTST CI-192 lr:1.53e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.112,fc_dropout:0.457,patch_size:16,stride:4,seq_len:192
-PatchTST CD-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
+PatchTST CI-192 lr: 1.53e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.112, fc_dropout: 0.457, patch_size: 16, stride: 4, seq_len: 192
+PatchTST CD-192 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
 TSMixer CI-192 lr: 0.0012, num_blocks: 4, hidden_size: 64, dropout: 0.323, activation: relu, seq_len: 192
-TSMixer CD-192 lr:0.0032,num_blocks:7,hidden_size:1.02e+03,dropout:0.006,activation:relu,seq_len:192
+TSMixer CD-192 lr: 0.0032, num_blocks: 7, hidden_size: 1024, dropout: 0.006, activation: relu, seq_len: 192
 CrossFormer CI-192 lr: 0.0006, e_layers: 7, d_ff: 128, d_model: 256, dropout: 0.309, seg_len: 7, baseline: 1, cross_factor: 18, seq_len: 336
@@
 TimeMixer CD-192 lr: 0.0006, d_ff: 256, d_model: 256, e_layers: 7, seq_len: 96
-PatchTST CI-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.539, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
-PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.5389, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
-TSMixer CI-336 lr:0.0012,num_blocks:8,hidden_size:1.02e+03,dropout:0.307,activation:relu,seq_len:192
+PatchTST CI-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
+PatchTST CD-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
+TSMixer CI-336 lr: 0.0012, num_blocks: 8, hidden_size: 1024, dropout: 0.307, activation: relu, seq_len: 192
 TSMixer CD-336 lr: 0.0010, num_blocks: 1, hidden_size: 64, dropout: 0.033, activation: relu, seq_len: 192
@@
 PatchTST CI-720 lr: 0.0001, e_layers: 9, d_ff: 512, d_model: 512, dropout: 0.065, fc_dropout: 0.539, patch_size: 8, stride: 4, seq_len: 96
-PatchTST CD-720 lr:0.0002,e_layers:1,d_ff:1.02e+03,d_model:512,dropout:0.471,fc_dropout:0.710,patch_size:8,stride:8,seq_len:336
-TSMixer CI-720 lr:0.0005,num_blocks:7,hidden_size:1.02e+03,dropout:0.169,activation:relu,seq_len:192
+PatchTST CD-720 lr: 0.0002, e_layers: 1, d_ff: 1024, d_model: 512, dropout: 0.471, fc_dropout: 0.710, patch_size: 8, stride: 8, seq_len: 336
+TSMixer CI-720 lr: 0.0005, num_blocks: 7, hidden_size: 1024, dropout: 0.169, activation: relu, seq_len: 192
 TSMixer CD-720 lr: 0.0010, num_blocks: 1, hidden_size: 64, dropout: 0.033, activation: relu, seq_len: 192
@@
 TimeMixer CD-720 lr: 0.0020, d_ff: 256, d_model: 256, e_layers: 8, seq_len: 96
-Table 18. Hyperparameter settings for the DoublePendulum dataset. The best variant for each model is shown.
-Model Setting Best Hyper. Params
-PatchTST CI-96 lr:0.0001,e_layers:1,d_ff:512,d_model:256,dropout:0.572,fc_dropout:0.631,patch_size:16,stride:4,seq_len:96
+Hyperparameter settings for the ETTh1 dataset across different forecasting horizons.
+Model Setting Best Hyperparameters
+[…]
+Model Setting Best Hyperparameters
+[…]
 PatchTST CD-96 lr: 0.0001, e_layers: 3, d_ff: 512, d_model: 512, dropout: 0.518, fc_dropout: 0.802, patch_size: 16, stride: 4, seq_len: 96
@@
 TimeMixer CD-96 lr: 0.0002, d_ff: 1024, d_model: 512, e_layers: 4, seq_len: 512
-PatchTST CI-192 lr:7.27e-06,e_layers:4,d_ff:1.02e+03,d_model:512,dropout:0.053,fc_dropout:0.183,patch_size:16,stride:4,seq_len:192
-PatchTST CD-192 lr:0.0001,e_layers:3,d_ff:1.02e+03,d_model:128,dropout:0.705,fc_dropout:0.640,patch_size:16,stride:8,seq_len:512
+PatchTST CI-192 lr: 7.27e-06, e_layers: 4, d_ff: 1024, d_model: 512, dropout: 0.053, fc_dropout: 0.183, patch_size: 16, stride: 4, seq_len: 192
+PatchTST CD-192 lr: 0.0001, e_layers: 3, d_ff: 1024, d_model: 128, dropout: 0.705, fc_dropout: 0.640, patch_size: 16, stride: 8, seq_len: 512
 TSMixer CI-192 lr: 0.0005, num_blocks: 1, hidden_size: 64, dropout: 0.173, activation: relu, seq_len: 192
@@
 TimeMixer CD-192 lr: 0.0001, d_ff: 1024, d_model: 256, e_layers: 9, seq_len: 512
-PatchTST CI-336 lr’: 5.207e-05, ’e_layers’: 1, ’d_ff’: 512, ’d_model’: 256, ’dropout’: 0.572, ’fc_dropout’: 0.6308, ’patch_size’: 16, ’stride’: 4, ’seq_len’: 96
-PatchTST CD-336 lr’: 2.920e-06, ’e_layers’: 6, ’d_ff’: 1024, ’d_model’: 512, ’dropout’: 0.342, ’fc_dropout’: 0.404, ’patch_size’: 16, ’stride’: 4, ’seq_len’: 192
+PatchTST CI-336 lr: 5.21e-05, e_layers: 1, d_ff: 512, d_model: 256, dropout: 0.572, fc_dropout: 0.631, patch_size: 16, stride: 4, seq_len: 96
+PatchTST CD-336 lr: 2.92e-06, e_layers: 6, d_ff: 1024, d_model: 512, dropout: 0.342, fc_dropout: 0.404, patch_size: 16, stride: 4, seq_len: 192
 TSMixer CI-336 lr: 0.0002, num_blocks: 2, hidden_size: 64, dropout: 0.205, activation: relu, seq_len: 512
@@
 PatchTST CI-720 lr: 1.78e-06, e_layers: 2, d_ff: 512, d_model: 512, dropout: 0.438, fc_dropout: 0.405, patch_size: 16, stride: 4, seq_len: 336
-PatchTST CD-720 lr:0.0024,e_layers:2,d_ff:1.02e+03,d_model:512,dropout:0.450,fc_dropout:0.374,patch_size:16,stride:4,seq_len:512
+PatchTST CD-720 lr: 0.0024, e_layers: 2, d_ff: 1024, d_model: 512, dropout: 0.450, fc_dropout: 0.374, patch_size: 16, stride: 4, seq_len: 512
 TSMixer CI-720 lr: 0.0002, num_blocks: 2, hidden_size: 64, dropout: 0.205, activation: relu, seq_len: 512
@@
 TimeMixer CD-720 lr: 0.0005, d_ff: 512, d_model: 256, e_layers: 8, seq_len: 512
-Table 19. Hyperparameter settings for the ETTh1 dataset. The best variant for each model is shown.
-Model Setting Best Hyper. Params
-PatchTST CI-96 lr:3.50e-05,e_layers:1,d_ff:512,d_model:256,dropout:0.571,fc_dropout:0.486,patch_size:16,stride:4,seq_len:96
+Hyperparameter settings for the ETTh2 dataset across different forecasting horizons.
+Model Setting Best Hyperparameters
+[…]
+Model Setting Best Hyperparameters
+[…]
 PatchTST CD-96 lr: 0.0001, e_layers: 3, d_ff: 512, d_model: 512, dropout: 0.688, fc_dropout: 0.393, patch_size: 16, stride: 4, seq_len: 96
 TSMixer CI-96 lr: 0.0005, num_blocks: 1, hidden_size: 64, dropout: 0.166, activation: relu, seq_len: 192
-TSMixer CD-96 lr:0.0011,num_blocks:6,hidden_size:1.02e+03,dropout:0.746,activation:gelu,seq_len:192
+TSMixer CD-96 lr: 0.0011, num_blocks: 6, hidden_size: 1024, dropout: 0.746, activation: gelu, seq_len: 192
 CrossFormer CI-96 lr: 0.0003, e_layers: 5, d_ff: 512, d_model: 128, dropout: 0.707, seg_len: 3, baseline: 1, cross_factor: 10, seq_len: 96
@@
 TimeMixer CD-192 lr: 0.0004, d_ff: 256, d_model: 256, e_layers: 4, seq_len: 96
-PatchTST CI-336 lr’: 3.495e-05, ’e_layers’: 1, ’d_ff’: 512, ’d_model’: 256, ’dropout’: 0.5706, ’fc_dropout’: 0.486, ’patch_size’: 16, ’stride’: 4, ’seq_len’: 96
-PatchTST CD-336 lr’: 0.0002, ’e_layers’: 3, ’d_ff’: 512, ’d_model’: 512, ’dropout’: 0.6877, ’fc_dropout’: 0.5936, ’patch_size’: 16, ’stride’: 4, ’seq_len’: 96
+PatchTST CI-336 lr: 3.50e-05, e_layers: 1, d_ff: 512, d_model: 256, dropout: 0.571, fc_dropout: 0.486, patch_size: 16, stride: 4, seq_len: 96
+PatchTST CD-336 lr: 0.0002, e_layers: 3, d_ff: 512, d_model: 512, dropout: 0.688, fc_dropout: 0.594, patch_size: 16, stride: 4, seq_len: 96
 TSMixer CI-336 lr: 0.0097, num_blocks: 5, hidden_size: 32, dropout: 0.331, activation: relu, seq_len: 192
@@
 PatchTST CD-720 lr: 0.0001, e_layers: 3, d_ff: 512, d_model: 512, dropout: 0.518, fc_dropout: 0.802, patch_size: 16, stride: 4, seq_len: 96
-TSMixer CI-720 lr:0.0021,num_blocks:6,hidden_size:1.02e+03,dropout:0.141,activation:relu,seq_len:192
+TSMixer CI-720 lr: 0.0021, num_blocks: 6, hidden_size: 1024, dropout: 0.141, activation: relu, seq_len: 192
 TSMixer CD-720 lr: 0.0010, num_blocks: 1, hidden_size: 64, dropout: 0.033, activation: relu, seq_len: 192
@@
 TimeMixer CD-720 lr: 0.0033, d_ff: 512, d_model: 128, e_layers: 8, seq_len: 96
-Table 20. Hyperparameter settings for the ETTh2 dataset. The best variant for each model is shown.
-Model Setting Best Hyper. Params
-PatchTST CI-96 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
-PatchTST CD-96 lr:6.07e-06,e_layers:3,d_ff:256,d_model:1.02e+03,dropout:0.006,fc_dropout:0.642,patch_size:16,stride:8,seq_len:192
-TSMixer CI-96 lr:0.0097,num_blocks:2,hidden_size:1.02e+03,dropout:0.115,activation:relu,seq_len:192
+Hyperparameter settings for the ETTm1 dataset across different forecasting horizons.
+Model Setting Best Hyperparameters
+[…]
+Model Setting Best Hyperparameters
+[…]
+PatchTST CD-96 lr: 6.07e-06, e_layers: 3, d_ff: 256, d_model: 1024, dropout: 0.006, fc_dropout: 0.642, patch_size: 16, stride: 8, seq_len: 192
+TSMixer CI-96 lr: 0.0097, num_blocks: 2, hidden_size: 1024, dropout: 0.115, activation: relu, seq_len: 192
 TSMixer CD-96 lr: 0.0084, num_blocks: 2, hidden_size: 32, dropout: 0.264, activation: relu, seq_len: 512
@@
 TimeMixer CD-96 lr: 0.0009, d_ff: 512, d_model: 512, e_layers: 8, seq_len: 192
-PatchTST CI-192 lr:2.97e-06,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.081,fc_dropout:0.435,patch_size:16,stride:8,seq_len:192
-PatchTST CD-192 lr:2.97e-06,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.081,fc_dropout:0.435,patch_size:16,stride:8,seq_len:192
-TSMixer CI-192 lr:0.0097,num_blocks:2,hidden_size:1.02e+03,dropout:0.115,activation:relu,seq_len:192
+PatchTST CI-192 lr: 2.97e-06, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.081, fc_dropout: 0.435, patch_size: 16, stride: 8, seq_len: 192
+PatchTST CD-192 lr: 2.97e-06, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.081, fc_dropout: 0.435, patch_size: 16, stride: 8, seq_len: 192
+TSMixer CI-192 lr: 0.0097, num_blocks: 2, hidden_size: 1024, dropout: 0.115, activation: relu, seq_len: 192
 TSMixer CD-192 lr: 0.0005, num_blocks: 6, hidden_size: 64, dropout: 0.280, activation: relu, seq_len: 192
@@
 TimeMixer CD-192 lr: 0.0002, d_ff: 512, d_model: 1024, e_layers: 7, seq_len: 192
-PatchTST CI-336 lr’: 0.0003, ’e_layers’: 9, ’d_ff’: 256, ’d_model’: 128, ’dropout’: 0.228, ’fc_dropout’: 0.0716, ’patch_size’: 8, ’stride’: 8, ’seq_len’: 336
-PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.5389, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
-TSMixer CI-336 lr:0.0006,num_blocks:8,hidden_size:1.02e+03,dropout:0.166,activation:relu,seq_len:192
+PatchTST CI-336 lr: 0.0003, e_layers: 9, d_ff: 256, d_model: 128, dropout: 0.228, fc_dropout: 0.0716, patch_size: 8, stride: 8, seq_len: 336
+PatchTST CD-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.0137, fc_dropout: 0.5389, patch_size: 16, stride: 8, seq_len: 192
+TSMixer CI-336 lr: 0.0006, num_blocks: 8, hidden_size: 1024, dropout: 0.166, activation: relu, seq_len: 192
 TSMixer CD-336 lr: 0.0004, num_blocks: 8, hidden_size: 64, dropout: 0.215, activation: relu, seq_len: 192
@@
 TimeMixer CD-336 lr: 0.0021, d_ff: 256, d_model: 128, e_layers: 4, seq_len: 192
-PatchTST CI-720 lr:2.27e-06,e_layers:4,d_ff:1.02e+03,d_model:128,dropout:0.198,fc_dropout:0.356,patch_size:16,stride:8,seq_len:512
-PatchTST CD-720 lr:0.0019,e_layers:7,d_ff:1.02e+03,d_model:512,dropout:0.471,fc_dropout:0.375,patch_size:8,stride:8,seq_len:720
-TSMixer CI-720 lr:0.0097,num_blocks:2,hidden_size:1.02e+03,dropout:0.115,activation:relu,seq_len:192
-TSMixer CD-720 lr:0.0011,num_blocks:6,hidden_size:1.02e+03,dropout:0.339,activation:relu,seq_len:192
+PatchTST CI-720 lr: 2.27e-06, e_layers: 4, d_ff: 1024, d_model: 128, dropout: 0.198, fc_dropout: 0.356, patch_size: 16, stride: 8, seq_len: 512
+PatchTST CD-720 lr: 0.0019, e_layers: 7, d_ff: 1024, d_model: 512, dropout: 0.471, fc_dropout: 0.375, patch_size: 8, stride: 8, seq_len: 720
+TSMixer CI-720 lr: 0.0097, num_blocks: 2, hidden_size: 1024, dropout: 0.115, activation: relu, seq_len: 192
+TSMixer CD-720 lr: 0.0011, num_blocks: 6, hidden_size: 1024, dropout: 0.339, activation: relu, seq_len: 192
 CrossFormer CI-720 lr: 0.0001, e_layers: 6, d_ff: 128, d_model: 256, dropout: 0.272, seg_len: 7, baseline: 1, cross_factor: 20, seq_len: 336
@@
 TimeMixer CD-720 lr: 0.0001, d_ff: 1024, d_model: 1024, e_layers: 2, seq_len: 192
-Table 21. Hyperparameter settings for the ETTm1 dataset. The best variant for each model is shown.
-Model Setting Best Hyper. Params
-PatchTST CI-96 lr:0.0002,e_layers:7,d_ff:1.02e+03,d_model:1.02e+03,dropout:0.648,fc_dropout:0.175,patch_size:8,stride:4,seq_len:512
-PatchTST CD-96 lr:1.60e-06,e_layers:8,d_ff:256,d_model:1.02e+03,dropout:0.307,fc_dropout:0.623,patch_size:8,stride:4,seq_len:512
+Hyperparameter settings for the ETTm2 dataset across different forecasting horizons.
+Model Setting Best Hyperparameters
+[…]
+Model Setting Best Hyperparameters
+[…]
+PatchTST CD-96 lr: 1.60e-06, e_layers: 8, d_ff: 256, d_model: 1024, dropout: 0.307, fc_dropout: 0.623, patch_size: 8, stride: 4, seq_len: 512
 TSMixer CI-96 lr: 0.0097, num_blocks: 5, hidden_size: 32, dropout: 0.289, activation: relu, seq_len: 512
@@
 TimeMixer CD-96 lr: 5.00e-06, d_ff: 512, d_model: 512, e_layers: 3, seq_len: 720
-PatchTST CI-192 lr:7.02e-07,e_layers:2,d_ff:1.02e+03,d_model:512,dropout:0.292,fc_dropout:0.673,patch_size:8,stride:4,seq_len:336
-PatchTST CD-192 lr:1.82e-06,e_layers:4,d_ff:1.02e+03,d_model:1.02e+03,dropout:0.471,fc_dropout:0.405,patch_size:16,stride:8,seq_len:512
+PatchTST CI-192 lr: 7.02e-07, e_layers: 2, d_ff: 1024, d_model: 512, dropout: 0.292, fc_dropout: 0.673, patch_size: 8, stride: 4, seq_len: 336
+PatchTST CD-192 lr: 1.82e-06, e_layers: 4, d_ff: 1024, d_model: 1024, dropout: 0.471, fc_dropout: 0.405, patch_size: 16, stride: 8, seq_len: 512
 TSMixer CI-192 lr: 0.0010, num_blocks: 1, hidden_size: 64, dropout: 0.033, activation: relu, seq_len: 192
@@
 TimeMixer CD-192 lr: 0.0003, d_ff: 512, d_model: 128, e_layers: 4, seq_len: 720
-PatchTST CI-336 lr’: 0.0002, ’e_layers’: 7, ’d_ff’: 1024, ’d_model’: 1024, ’dropout’: 0.6482, ’fc_dropout’: 0.1748, ’patch_size’: 8, ’stride’: 4, ’seq_len’: 512
-PatchTST CD-336 lr’: 1.602e-06, ’e_layers’: 8, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.307, ’fc_dropout’: 0.6229, ’patch_size’: 8, ’stride’: 4, ’seq_len’: 512
+PatchTST CI-336 lr: 0.0002, e_layers: 7, d_ff: 1024, d_model: 1024, dropout: 0.6482, fc_dropout: 0.1748, patch_size: 8, stride: 4, seq_len: 512
+PatchTST CD-336 lr: 1.60e-06, e_layers: 8, d_ff: 256, d_model: 1024, dropout: 0.307, fc_dropout: 0.6229, patch_size: 8, stride: 4, seq_len: 512
 TSMixer CI-336 lr: 0.0097, num_blocks: 5, hidden_size: 32, dropout: 0.328, activation: relu, seq_len: 512
@@
 TimeMixer CD-336 lr: 1.00e-05, d_ff: 512, d_model: 128, e_layers: 1, seq_len: 720
-PatchTST CI-720 lr:2.27e-06,e_layers:4,d_ff:1.02e+03,d_model:128,dropout:0.198,fc_dropout:0.356,patch_size:16,stride:8,seq_len:512
-PatchTST CD-720 lr:0.0032,e_layers:3,d_ff:1.02e+03,d_model:128,dropout:0.318,fc_dropout:0.402,patch_size:16,stride:8,seq_len:512
+PatchTST CI-720 lr: 2.27e-06, e_layers: 4, d_ff: 1024, d_model: 128, dropout: 0.198, fc_dropout: 0.356, patch_size: 16, stride: 8, seq_len: 512
+PatchTST CD-720 lr: 0.0032, e_layers: 3, d_ff: 1024, d_model: 128, dropout: 0.318, fc_dropout: 0.402, patch_size: 16, stride: 8, seq_len: 512
 TSMixer CI-720 lr: 0.0097, num_blocks: 5, hidden_size: 32, dropout: 0.289, activation: relu, seq_len: 512
@@
 TimeMixer CD-720 lr: 1.00e-05, d_ff: 512, d_model: 128, e_layers: 1, seq_len: 720
-Table 22. Hyperparameter settings for the ETTm2 dataset. The best variant for each model is shown.
-Model Setting Best Hyper. Params
-PatchTST CI-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
+Hyperparameter settings for the Hopfield dataset across different forecasting horizons.
+Model Setting Best Hyperparameters
+[…]
+Model Setting Best Hyperparameters
+[…]
 PatchTST CD-96 lr: 0.0003, e_layers: 10, d_ff: 256, d_model: 128, dropout: 0.081, fc_dropout: 0.390, patch_size: 8, stride: 8, seq_len: 336
@@
 TimeMixer CD-192 lr: 0.0010, d_ff: 1024, d_model: 1024, e_layers: 10, seq_len: 192
-PatchTST CI-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.014, ’fc_dropout’: 0.539, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
-PatchTST CD-336 lr:0.0004,num_blocks:8,hidden_size:1.02e+03,dropout:0.159,activation:relu,seq_len:512
+PatchTST CI-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
+PatchTST CD-336 lr: 0.0004, num_blocks: 8, hidden_size: 1024, dropout: 0.159, activation: relu, seq_len: 512
 TSMixer CI-336 lr: 0.0011, num_blocks: 1, hidden_size: 256, dropout: 0.322, activation: relu, seq_len: 512
@@
 PatchTST CI-720 lr: 0.0003, e_layers: 10, d_ff: 256, d_model: 128, dropout: 0.081, fc_dropout: 0.390, patch_size: 8, stride: 8, seq_len: 336
-PatchTST CD-720 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
-TSMixer CI-720 lr:0.0002,num_blocks:8,hidden_size:1.02e+03,dropout:0.205,activation:relu,seq_len:512
+PatchTST CD-720 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
+TSMixer CI-720 lr: 0.0002, num_blocks: 8, hidden_size: 1024, dropout: 0.205, activation: relu, seq_len: 512
 TSMixer CD-720 lr: 0.0010, num_blocks: 1, hidden_size: 64, dropout: 0.033, activation: relu, seq_len: 192
@@
 TimeMixer CD-720 lr: 0.0010, d_ff: 1024, d_model: 1024, e_layers: 10, seq_len: 192
-Table 23. Hyperparameter settings for the Hopfield dataset. The best variant for each model is shown.
-Model Setting Best Hyper. Params
-PatchTST CI-96 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
+Hyperparameter settings for the Lorenz dataset across different forecasting horizons.
+Model Setting Best Hyperparameters
+[…]
+Model Setting Best Hyperparameters
+[…]
 PatchTST CD-96 lr: 0.0003, e_layers: 9, d_ff: 256, d_model: 128, dropout: 0.228, fc_dropout: 0.072, patch_size: 8, stride: 8, seq_len: 336
@@
 TimeMixer CD-96 lr: 0.0028, d_ff: 256, d_model: 128, e_layers: 4, seq_len: 192
-PatchTST CI-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
-PatchTST CD-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
+PatchTST CI-192 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
+PatchTST CD-192 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
 TSMixer CI-192 lr: 0.0041, num_blocks: 10, hidden_size: 256, dropout: 0.415, activation: gelu, seq_len: 96
@@
 TimeMixer CD-192 lr: 0.0016, d_ff: 256, d_model: 128, e_layers: 5, seq_len: 192
-PatchTST CI-336 lr’: 0.0003, ’e_layers’: 9, ’d_ff’: 256, ’d_model’: 128, ’dropout’: 0.2277, ’fc_dropout’: 0.0716, ’patch_size’: 8, ’stride’: 8, ’seq_len’: 336
-PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.5389, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
+PatchTST CI-336 lr: 0.0003, e_layers: 9, d_ff: 256, d_model: 128, dropout: 0.2277, fc_dropout: 0.0716, patch_size: 8, stride: 8, seq_len: 336
+PatchTST CD-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.0137, fc_dropout: 0.5389, patch_size: 16, stride: 8, seq_len: 192
 TSMixer CI-336 lr: 0.0041, num_blocks: 10, hidden_size: 256, dropout: 0.415, activation: gelu, seq_len: 96
@@
 TimeMixer CD-720 lr: 0.0041, d_ff: 256, d_model: 1024, e_layers: 6, seq_len: 96
-Table 24. Hyperparameter settings for the Lorenz dataset. The best variant for each model is shown.
-Model Setting Best Hyper. Params
-PatchTST CI-96 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
+Hyperparameter settings for the LorenzCoupled dataset.
+Model Setting Best Hyperparameters
+[…]
+Model Setting Best Hyperparameters
+[…]
 PatchTST CD-96 lr: 0.0003, e_layers: 9, d_ff: 256, d_model: 128, dropout: 0.228, fc_dropout: 0.072, patch_size: 8, stride: 8, seq_len: 336
@@
 TimeMixer CD-96 lr: 0.0006, d_ff: 1024, d_model: 128, e_layers: 5, seq_len: 192
-PatchTST CI-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
-PatchTST CD-192 lr:1.44e-05,e_layers:10,d_ff:256,d_model:1.02e+03,dropout:0.014,fc_dropout:0.539,patch_size:16,stride:8,seq_len:192
+PatchTST CI-192 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
+PatchTST CD-192 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.014, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
 TSMixer CI-192 lr: 0.0041, num_blocks: 10, hidden_size: 256, dropout: 0.415, activation: gelu, seq_len: 96
@@
 TimeMixer CD-192 lr: 0.0024, d_ff: 256, d_model: 512, e_layers: 6, seq_len: 96
-PatchTST CI-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.539, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
-PatchTST CD-336 lr’: 1.441e-05, ’e_layers’: 10, ’d_ff’: 256, ’d_model’: 1024, ’dropout’: 0.0137, ’fc_dropout’: 0.5389, ’patch_size’: 16, ’stride’: 8, ’seq_len’: 192
+PatchTST CI-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.0137, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
+PatchTST CD-336 lr: 1.44e-05, e_layers: 10, d_ff: 256, d_model: 1024, dropout: 0.0137, fc_dropout: 0.539, patch_size: 16, stride: 8, seq_len: 192
 TSMixer CI-336 lr: 0.0041, num_blocks: 10, hidden_size: 256, dropout: 0.415, activation: gelu, seq_len: 96
@@
 TimeMixer CD-336 lr: 0.0011, d_ff: 256, d_model: 256, e_layers: 7, seq_len: 96
-PatchTST CI-720 lr:0.0001,e_layers:2,d_ff:1.02e+03,d_model:1.02e+03,dropout:0.463,fc_dropout:0.708,patch_size:16,stride:4,seq_len:192
-PatchTST CD-720 lr:0.0018,e_layers:5,d_ff:1.02e+03,d_model:128,dropout:0.097,fc_dropout:0.623,patch_size:16,stride:8,seq_len:512
+PatchTST CI-720 lr: 0.0001, e_layers: 2, d_ff: 1024, d_model: 1024, dropout: 0.463, fc_dropout: 0.708, patch_size: 16, stride: 4, seq_len: 192
+PatchTST CD-720 lr: 0.0018, e_layers: 5, d_ff: 1024, d_model: 128, dropout: 0.097, fc_dropout: 0.623, patch_size: 16, stride: 8, seq_len: 512
 TSMixer CI-720 lr: 0.0005, num_blocks: 6, hidden_size: 64, dropout: 0.280, activation: relu, seq_len: 192
@@
 TimeMixer CD-720 lr: 0.0041, d_ff: 256, d_model: 1024, e_layers: 6, seq_len: 96
-Table 25. Hyperparameter settings for the LorenzCoupled dataset. The best variant for each model is shown.
-Model Setting Best Hyper. Params
-PatchTST CI-96
+Hyperparameter settings for the electricity dataset.
+Model Setting Best Hyperparameters
+[…]
+Model Setting Best Hyperparameters
+[…]
 PatchTST CD-96
-TSMixer CI-96 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
+TSMixer CI-96 lr: 0.0067, num_blocks: 9, hidden_size: 1024, dropout: 0.318, activation: relu, seq_len: 512
 TSMixer CD-96 lr: 0.0003, num_blocks: 1, hidden_size: 64, dropout: 0.166, activation: relu, seq_len: 512
@@
 PatchTST CD-192
-TSMixer CI-192 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
+TSMixer CI-192 lr: 0.0067, num_blocks: 9, hidden_size: 1024, dropout: 0.318, activation: relu, seq_len: 512
 TSMixer CD-192 lr: 0.0097, num_blocks: 6, hidden_size: 32, dropout: 0.315, activation: relu, seq_len: 512
@@
 iTransformers CD-192 lr: 0.0004, e_layers: 1, d_ff: 2048, d_model: 1024, dropout: 0.080, seq_len: 192
-TimeMixer CI-192 lr 0.00008, d_ff 512, d_model 512, e_layers 4, seq_len 192
+TimeMixer CI-192 lr: 0.00008, d_ff: 512, d_model: 512, e_layers: 4, seq_len: 192
 TimeMixer CD-192 lr: 0.0015, d_ff: 256, d_model: 256, e_layers: 7, seq_len: 512
-PatchTST CI-336 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
+PatchTST CI-336 lr: 0.0067, num_blocks: 9, hidden_size: 1024, dropout: 0.318, activation: relu, seq_len: 512
 PatchTST CD-336
@@
 CrossFormer CI-336 lr: 0.0080, e_layers: 1, d_ff: 128, d_model: 128, dropout: 0.189, seg_len: 8, baseline: 0, cross_factor: 9, seq_len: 192
-CrossFormer CD-336 lr’: 0.00017626081318364395, ’e_layers’: 7, ’d_ff’: 128, ’d_model’: 512, ’dropout’: 0.3870070203430017, ’seg_len’: 10, ’baseline’: 0, ’cross_factor’: 16, ’seq_len’: 192, ’affine’: 1
+CrossFormer CD-336 lr: 0.000176, e_layers: 7, d_ff: 128, d_model: 512, dropout: 0.387, seg_len: 10, baseline: 0, cross_factor: 16, seq_len: 192, affine: 1
 DLinear CI-336 lr: 0.0098, seq_len: 512
@@
 TimeMixer CD-720 lr: 0.0021, d_ff: 256, d_model: 128, e_layers: 4, seq_len: 192
-Table 26. Hyperparameter settings for the electricity dataset. The best variant for each model is shown.
-Model Setting Best Hyper. Params
-PatchTST CI-96 lr:0.0003,e_layers:9,d_ff:256,d_model:128,dropout:0.228,fc_dropout:0.072,patch_size:8,stride:8,seq_len:336
-PatchTST CD-96 lr:0.0001,e_layers:1,d_ff:1.02e+03,d_model:128,dropout:0.098,fc_dropout:0.465,patch_size:8,stride:8,seq_len:336
-TSMixer CI-96 lr:0.0002,num_blocks:5,hidden_size:1.02e+03,dropout:0.205,activation:relu,seq_len:512
+Hyperparameter settings for the weather dataset.
+Model Setting Best Hyperparameters
+[…]
+Model Setting Best Hyperparameters
+[…]
+PatchTST CD-96 lr: 0.0001, e_layers: 1, d_ff: 1024, d_model: 128, dropout: 0.098, fc_dropout: 0.465, patch_size: 8, stride: 8, seq_len: 336
+TSMixer CI-96 lr: 0.0002, num_blocks: 5, hidden_size: 1024, dropout: 0.205, activation: relu, seq_len: 512
 TSMixer CD-96 lr: 0.0010, num_blocks: 1, hidden_size: 64, dropout: 0.033, activation: relu, seq_len: 192
@@
 PatchTST CI-192 lr: 0.0001, e_layers: 4, d_ff: 256, d_model: 128, dropout: 0.469, fc_dropout: 0.405, patch_size: 8, stride: 8, seq_len: 512
-PatchTST CD-192 lr:0.0002,e_layers:1,d_ff:512,d_model:1.02e+03,dropout:0.010,fc_dropout:0.333,patch_size:16,stride:8,seq_len:192
-TSMixer CI-192 lr:0.0002,num_blocks:5,hidden_size:1.02e+03,dropout:0.247,activation:relu,seq_len:512
+PatchTST CD-192 lr: 0.0002, e_layers: 1, d_ff: 512, d_model: 1024, dropout: 0.010, fc_dropout: 0.333, patch_size: 16, stride: 8, seq_len: 192
+TSMixer CI-192 lr: 0.0002, num_blocks: 5, hidden_size: 1024, dropout: 0.247, activation: relu, seq_len: 512
 TSMixer CD-192 lr: 0.0010, num_blocks: 1, hidden_size: 64, dropout: 0.033, activation: relu, seq_len: 192
@@
 iTransformers CD-192 lr: 1.60e-05, e_layers: 9, d_ff: 2048, d_model: 1024, dropout: 0.012, seq_len: 192
-TimeMixer CI-192 lr 0.00008, d_ff 512, d_model 512, e_layers 4, seq_len 192
+TimeMixer CI-192 lr: 0.00008, d_ff: 512, d_model: 512, e_layers: 4, seq_len: 192
 TimeMixer CD-192 lr: 0.0005, d_ff: 512, d_model: 128, e_layers: 8, seq_len: 720
@@
 PatchTST CD-336 lr: 0.0007, e_layers: 8, d_ff: 1024, d_model: 512, dropout: 0.150, fc_dropout: 0.021, patch_size: 8, stride: 8, seq_len: 720
-TSMixer CI-336 lr:0.0002,num_blocks:5,hidden_size:1.02e+03,dropout:0.265,activation:relu,seq_len:512
+TSMixer CI-336 lr: 0.0002, num_blocks: 5, hidden_size: 1024, dropout: 0.265, activation: relu, seq_len: 512
 TSMixer CD-336 lr: 0.0001, num_blocks: 6, hidden_size: 256, dropout: 0.334, activation: gelu, seq_len: 512
@@
 PatchTST CI-720 lr: 0.0005, e_layers: 7, d_ff: 256, d_model: 256, dropout: 0.767, fc_dropout: 0.878, patch_size: 16, stride: 8, seq_len: 720
-PatchTST CD-720 lr:6.07e-06,e_layers:3,d_ff:256,d_model:1.02e+03,dropout:0.006,fc_dropout:0.642,patch_size:16,stride:8,seq_len:192
-TSMixer CI-720 lr:0.0003,num_blocks:2,hidden_size:1.02e+03,dropout:0.108,activation:relu,seq_len:512
+PatchTST CD-720 lr: 6.07e-06, e_layers: 3, d_ff: 256, d_model: 1024, dropout: 0.006, fc_dropout: 0.642, patch_size: 16, stride: 8, seq_len: 192
+TSMixer CI-720 lr: 0.0003, num_blocks: 2, hidden_size: 1024, dropout: 0.108, activation: relu, seq_len: 512
 TSMixer CD-720 lr: 0.0002, num_blocks: 8, hidden_size: 32, dropout: 0.281, activation: relu, seq_len: 512
@@
 TimeMixer CD-720 lr: 1.00e-05, d_ff: 512, d_model: 512, e_layers: 1, seq_len: 720
-Table 27. Hyperparameter settings for the weather dataset. The best variant for each model is shown.
-Model Setting Best Hyper. Params
-PatchTST CI-96
+Hyperparameter settings for the Traffic dataset.
+Model Setting Best Hyperparameters
+[…]
+Model Setting Best Hyperparameters
+[…]
 PatchTST CD-96
@@
 TSMixer CD-96 lr: 0.0097, num_blocks: 5, hidden_size: 32, dropout: 0.311, activation: relu, seq_len: 512
-CrossFormer CI-96 lr’: 0.000515701560401399, ’e_layers’: 7, ’d_ff’: 128, ’d_model’: 256, ’dropout’: 0.25730612136617825, ’seg_len’: 11, ’baseline’: 0, ’cross_factor’: 10, ’seq_len’: 96, ’affine’: 1
-CrossFormer CD-96 lr’: 5.2122436504604685e-05, ’e_layers’: 3, ’d_ff’: 256, ’d_model’: 512, ’dropout’: 0.12303593071645767, ’seg_len’: 8, ’baseline’: 0, ’cross_factor’: 7, ’seq_len’: 336, ’affine’: 0
+CrossFormer CI-96 lr: 0.000516, e_layers: 7, d_ff: 128, d_model: 256, dropout: 0.257, seg_len: 11, baseline: 0, cross_factor: 10, seq_len: 96, affine: 1
+CrossFormer CD-96 lr: 5.21e-05, e_layers: 3, d_ff: 256, d_model: 512, dropout: 0.123, seg_len: 8, baseline: 0, cross_factor: 7, seq_len: 336, affine: 0
 DLinear CI-96 lr: 0.0078, seq_len: 512
@@
 PatchTST CD-192
-TSMixer CI-192 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
+TSMixer CI-192 lr: 0.0067, num_blocks: 9, hidden_size: 1024, dropout: 0.318, activation: relu, seq_len: 512
 TSMixer CD-192 lr: 0.0097, num_blocks: 5, hidden_size: 32, dropout: 0.311, activation: relu, seq_len: 512
@@
 PatchTST CD-336
-TSMixer CI-336 lr:0.0067,num_blocks:9,hidden_size:1.02e+03,dropout:0.318,activation:relu,seq_len:512
+TSMixer CI-336 lr: 0.0067, num_blocks: 9, hidden_size: 1024, dropout: 0.318, activation: relu, seq_len: 512
 TSMixer CD-336 lr: 0.0097, num_blocks: 5, hidden_size: 32, dropout: 0.311, activation: relu, seq_len: 512
@@
 TimeMixer CD-720 From TimeMixer Paper
-Table 28. Hyperparameter settings for the Traffic dataset. The best variant for each model is shown.